### v1

# Constructing equivalences between fusion categories of quantum groups and of vertex operator algebras via quantum gauge groups

Abstract.
We summarize our results in [11] on a problem that Sergio Doplicher posed in the 90s in the setting of conformal nets, and on a related problem posed by Yi-Zhi Huang in the setting of vertex operator algebras, see e.g. [53]. The latter asks to find a direct construction of an equivalence due to Finkelberg between the quantum group modular fusion category at certain roots of unity and the fusion categories of affine vertex operator algebras at positive integer levels. The former asks for the construction of a quantum gauge group and field algebra for the category of localized endomorphisms of a conformal net, extending Doplicher-Roberts theory of the compact gauge group and field algebra in high dimensional algebraic QFT. In this summary we mostly focus on the affine VOA aspects (corresponding to the loop group conformal nets).
We solve Huang’s problem for all the Lie types for which it is known that the centralizer algebras of the non-negligible tensor powers of a given generating object $V$ for the quantum group fusion category (a list of $V$ for which fusion tensor powers are well defined has been given by Hans Wenzl in [87] for all the Lie types), is generated by the representation of the braid group associated to the $R$-matrix, and for Huang-Lepowsky ribbon braided tensor structure for the vertex operator algebras. In the classical type $A$, this property becomes the classical Schur-Weyl duality for the vector representation, and was used by Doplicher and Roberts to construct ${\rm SU}(d)$ as a special case of their unique compact gauge group and field algebra in algebraic QFT. Presently, known results in the literature on this generating property allow to construct our equivalence for the Lie types $A$, $B$, $C$, $D$, $G_{2}$. Our approach may be regarded in the setting of Connes noncommutative geometry.
Lack of knowledge of the braid group generation property (i.e. for the Lie types $E$ and $F$) allows us to show equivalence with Huang-Lepowsky ribbon braided tensor structure for all the ribbon morphisms, and part of those braiding morphisms, and associativity morphisms for which $V$ takes one or two components, respectively.
Our work also gives a unitary modular tensor category structure on the linear category of modules of the Zhu algebra associated to the minimum energy functor (or on the vertex operator algebra) for all the Lie types, equivalent to that from quantum groups. This algebra becomes a unitary weak quasi-Hopf algebra, solving a problem posed by Frenkel and Zhu for all the Lie types [34].
Our main tool is the construction of two semisimple quantum gauge groups and an equivalence of their representation categories: The quantum group fusion category provides a unitary coboundary weak Hopf algebra associated to Wenzl functor in [87]. We then apply a de-quantization continuous curve of Wenzl and a Drinfeld twist method to reach the Zhu algebra. The structures we find in both contexts extend Doplicher-Roberts symmetric functors for rigid symmetric tensor categories with simple unit which play a central role in their duality theory.
Dedicated to Sergio Doplicher, with gratitude
1. Introduction
Moore and Seiberg [63], [64] discovered a connection between their equations and a certain structure of a tensor category arising from rational conformal field theories. The notion of ribbon braided tensor category has been settled in mathematics, and we refer to [21], [66], [27], [26], and, for their relations to compact quantum groups, to [70], and more generally to weak (quasi-) Hopf algebras to [11]. Turaev gave a definition of modular tensor category, based on previous work with Reshetikhin on the construction of knot and 3-manifold invariants see [85]. Several authors contributed to construct modular fusion categories form quantum groups at roots of unity, and a more complete historical account on the quantum group side may be found in [80], [82]. see also the introduction of [11]. We denote by ${\mathcal{C}}({\mathfrak{g}},q)$ the resulting modular fusion category, with $q=e^{i\pi/d\ell}$, and $d\in\{1,2,3\}$ a suitable integer structure constant from Lie theory. Unitary tensor structures have been constructed by Wenzl and Xu [87], [92], and the former paper, as well as more recent papers by Wenzl and his coauthors play a major role in our work, and we try to explain this in this paper.
On the side of CFT, our main interest in this paper is the Wess-Zumino-Witten model. Modular, or more generally ribbon rigid braided tensor category constructions involved many authors among physicists and mathematicians.
Major work was carried out in a series of papers by Kazhdan and Lusztig in the mid 90s. Let ${\mathfrak{g}}$ be a complex simple Lie algebra. The authors construct rigid braided tensor structures for certain module categories of affine Lie algebras associated to ${\mathfrak{g}}$ at levels excluding the positive rationals. In their work they consider both generic and non generic values of the levels. In particular they consider the negative integer levels $-k$, $k\in{\mathbb{N}}$, and proved a braided tensor equivalence of their category ${\mathcal{O}}_{-k}$ with the non-semisimple category of certain modules of the Drinfeld-Jimbo quantum group $U_{q}({\mathfrak{g}})$ at the same type of roots of unity $q=e^{i\pi/d\ell}$, with the exception of some exceptional cases at low levels [57]. The order $\ell$ appearing in the root of unity and the positive integer $k$ are related by a well known equality, see e.g. Def. 19.1 and 20.2 in [11] and references therein.
Shortly after, Finkelberg considered a certain semisimple subquotient $\tilde{\mathcal{O}}_{-k}$ of the representation category constructed by Kazhdan and Lusztig, paralleling the subquotient construction in the setting of quantum groups, and described a passage $\tilde{\mathcal{O}}_{k}\to\tilde{\mathcal{O}}_{-k}$ claimed as a braided tensor equivalence from the category $\tilde{\mathcal{O}}_{k}$ of certain integral modules of the affine Lie algebra at the positive integer level $k$ to the subquotient rigid braided tensor category $\tilde{\mathcal{O}}_{-k}$. Combining with the equivalence by Kazhdan and Lusztig, Finkelberg derived a connection between $\tilde{\mathcal{O}}_{k}\to{\mathcal{C}}({\mathfrak{g}},q)$. See the introduction and Theorem 4.3 in [30].
The original work by Finkelberg refers to the cited work in physics by Moore and Seiberg and in algebraic geometry by Beilinson, Feigin, and Mazur [2] for the rigid braided tensor structure on $\tilde{\mathcal{O}}_{k}$. As I learned from discussions with Huang, rigidity was not proved in [2], and there was no proof of rigidity using this approach.
Remarkably, Huang and Lepowsky introduced and developed a rigorous direct construction of a modular (vertex) tensor category structure in a very general setting for vertex operator algebras, including as a particular case the affine vertex operator algebras at positive integer levels that we denote by $V_{{\mathfrak{g}}_{k}}$. The construction has been completed by Huang between 2005 and 2008 [47], [51] following a general construction of tensor product theory in the setting of vertex operator algebras by Huang and Lepowsky [42], [43], [44] and by Huang [45], [50]. In particular [45] gives the construction of associativity morphisms which is the first difficulty in the construction of the tensor structure. An historical account may be found in [52]. We also refer to the introductions of [46] and [53]. These works give in particular the first mathematical proof of modularity, that applies moreover to a large class of representation categories of vertex operator algebras.
We note that difficulties in the construction of associativity morphisms in the setting of Drinfeld category were earlier solved by Drinfeld in [23], [24], [25] by the introduction and construction of quasi-Hopf algebras, via differential Knizhnik-Zamolodchikov equation, leading to his remarkable Drinfeld-Kohno theorem connecting via a Drinfeld twist with the quantized universal enveloping algebra. In this case, rigidity of Drinfeld category is derived from the equivalence constructed by Drinfeld in Drinfeld-Kohno theorem.
In particular rigidity and modularity property in the setting of vertex operator algebras was first shown in [47]. To show modularity, and in particular rigidity, Huang first discovered the need of the Verlinde formula. The paper [50] gives a proof of the Verlinde formula in the general case. See also [48] for the proof of convergence and extension property assumed in [45] and [49] for the proof of modular invariance conjecture of Moore and Seiberg needed in the proof of the Verlinde formula.
By a private communication with Huang on the history of this problem, I learned that it finally turned out that the construction by Finkelberg described in [30] connecting $\tilde{\mathcal{O}}_{k}$ with Kazhdan and Lusztig structure inherited by $\tilde{\mathcal{O}}_{-k}$ could not suffice to derive rigidity of $\tilde{\mathcal{O}}_{k}$ from that of $\tilde{\mathcal{O}}_{-k}$. This implied that the proof of the equivalence $\tilde{\mathcal{O}}_{k}\to\tilde{\mathcal{O}}_{-k}$ was far from complete.
I have been informed by Huang that these doubts were motivated by two facts. On one hand, his earlier direct communications in the 90s had clarified that rigidity was not proved in [2]. On the other, the mentioned property that his proof of rigidity in the setting of vertex operator algebras involved the Verlinde formula indicated that more work was likely needed to complete the proof in [30]. These facts together with early claims of rigidity in [30], motivated Huang to a comparison of the results in [30] with their results on rigidity and this led to note a gap in a proposition of [30] with implications on the main result.
Huang’s study had led him to identify the gap in the fact that a suitable constant needed to be shown to be nonzero in a proposition useful to obtain the equivalence between the two local systems considered by Finkelberg and Kazhdan and Lusztig, and this proposition would imply derivation of rigidity in his case from previous work. On the other hand, Huang proof of rigidity in the setting of vertex operator algebras had similarly reduced to show nontriviality of a constant, and to prove this he had used the Verlinde formula as above alluded to, which was a main result in their work. This had led Huang to inform Finkelberg about utility of this formula to show rigidity in his case.
As later reported by Finkelberg in [31], it turned out that the completion of the proof of that proposition leading to rigidity of $\tilde{\mathcal{O}}_{k}$ and the equivalence $\tilde{\mathcal{O}}_{k}\to\tilde{\mathcal{O}}_{-k}$, involved the Verlinde formula as well (which became another main result in Finkelberg’s paper). This reflects also on Finkelberg composed equivalence $\tilde{\mathcal{O}}_{k}\to\tilde{\mathcal{O}}_{-k}\to{\mathcal{C}}({\mathfrak{g}},q)$, with Kazhdan-Lusztig equivalence that moreover still passes through the negative integer levels. This implies in particular that the theorem does not include a few exceptional cases, including ${\mathfrak{g}}=E_{8}$ and $k=2$, which correspond to cases not treated in Kazhdan-Lusztig theory.
Following the introductions in [46], [53], Finkelberg work [30], [31] may be reinterpreted as giving the construction of a braided tensor category structure on $\tilde{\mathcal{O}}_{k}$. One may also conclude that these papers give a second proof of rigidity and modularity of this category.
An approach to modularity based on algebraic geometry may be found in Theorem 7.0.1 [1] but the proof of rigidity is not given and, as already mentioned it was known to experts that there was no proof based on their approach at the time. The author thanks Y.-Z. Huang for this information. We point out the very recent paper by Etingof and Penneys where the authors show among other things that weak rigidity in the sense of [1] is equivalent to rigidity for certain semisimple braided tensor categories of moderate growth with simple unit, see [28] giving a third proof of rigidity and give applications to a simplification of Huang’s proof of rigidity. They also answer positively to a question posed in [1] that for every finite semisimple category over an algebraically closed field the data of a modular functor is essentially equivalent to the structure of a modular fusion category, completing the proof of rigidity in [1].
On the other hand, Finkelberg equivalence theorem is stated as Theorem 7.0.2 in [1] without proof. The braided tensor equivalence between $\tilde{\mathcal{O}}_{k}\to{\mathcal{C}}({\mathfrak{g}},q)$ obtained from the combined works by Finkelberg, Kazhdan and Lusztig is considered indirect by several authors. This leads to the desire of constructing the equivalence in a natural, direct, and explicit way, and this problem is discussed by Y.-Z. Huang in [53] in great detail, where we select only a small part.
Problem 4.4 in [53]. Find a direct construction of this equivalence without using the equivalence given by Kazhdan-Lusztig so that this equivalence covers all the cases, including the important ${\mathfrak{g}}=E_{8}$ and $k=2$ case.
We also mention Gannon’s observation in Sect. 6.2.3 of [35] regarding the desire of finding a direct connection between quantum groups and CFT.
A direct construction of the equivalence may be helpful to see how rigidity or modularity in the setting of CFT may be transported from corresponding results in the setting of quantum groups at roots of unity, which seem simpler, or vice versa, by a given explicit equivalence. Or to simplify or give new construction of other structures, e.g. unitary structures, with a quantum group viewpoint, or to see more connections with quantum groups from these fundamental models. The study of how unitarization from the quantum group fusion categories [87] was useful to derive unitarization in representation categories of vertex operator algebras via an explicit equivalence was important for us since the early stages of [11], and further motivated us to study the above problem.
Our approach to Huang problem is based on interpreting Finkelberg-Kazhdan-Lusztig equivalence as an extension to quantum groups of Doplicher-Roberts duality theory for the important class of WZW models that are described by affine Lie algebras at positive integer levels. We shall briefly come back to the previous history of these important models in physics. From the side of mathematical axiomatizations of CFT, we work with respect to Huang and Lepowsky braided tensor structure for the representation category of vertex operator algebras.
Our starting point to approach this problem arose from a problem generously posed by Sergio Doplicher in the 90s to the author in a different setting, that of algebraic quantum field theory. We try to explain this problem.
Algebraic quantum field theory had been initiated by Haag and Kastler in [40] in the 60s, and developed by founding breakthrough works by Doplicher, Haag, and Roberts [15], [16], [17], [18] in the 70s. We refer to the book by Haag [41], for an in-depth introduction to this theory, based on functional analysis, and more specifically on operator algebras, and also for more references.
In this context, I learned from private discussions that for high dimensional quantum field theories categorical structures in their study first emerged from the discovery by Doplicher of the notion of localized endomorphism of the local observable algebra, allowing composition of two of them which was understood as a strict tensor product structure. The mentioned works by Doplicher, Haag, and Roberts led to a deep construction of the structure of a rigid symmetric tensor $C^{*}$-category with simple unit, starting from first principles in physics.
Rigidity, called charge conjugation in this setting, is due to particle-antiparticle duality. In their papers rigidity was shown to imply deep and useful properties for the classification of the statistics parameter, and charges of para-Bose or para-Fermi type. Because of the symmetry, their results on the classification statistics parameter, which was shown to the the reciprocal of a positive integer $d$ up to a sign, and $C^{*}$-structure, in the 70s it arose the question of whether that was the category of representations of a compact group, because those data arising from a single localised endomorphism were also satisfied by the representation category of a compact Lie subgroup of ${\rm SU}(d)$, that would be a gauge group of global symmetries. They considered the abelian case in the 70s [17]. In that case, every object had dimension $1$, and in modern terminology these categories are referred to as pointed in e.g. [26].
In a series of important papers dated the end of the 80s, Doplicher and Roberts were able to solve the problem in full generality. Their construction gives a unique compact gauge group with representation category equivalent to that of a given symmetric rigid tensor $C^{*}$-category with simple unit, yielding the formulation of a new abstract duality theory for compact groups [19], [21],[20], and led to desired applications to high dimensional Algebraic QFT [22].
The term abstract duality refers to the property that the given categories do not have a tensor functor to the category of f.d. vector spaces from the beginning, raising the problem of constructing these structures. It is thus an abstract form of duality in the sense of Tannaka-Krein, where in the latter the tensor structure on the embedding functor is given from the start. In their theory, an important role is played by the notion of symmetric functor to the category of f.d. vector spaces, for which they show existence and uniqueness. This property allows to find their unique compact gauge group $G$.
In two or three dimensional quantum field theory, it was noted that the corresponding categories were braided rather than symmetric. This implies that Bose-Fermi alternative does not hold any longer. Models were first described by Streater and Wilde. Doplicher, Haag, Roberts theory of superselection sectors was generalized to braid groups statistics by Fredenhagen, Rheren, Schroer [32], [33]. The representation of the braid group in the tensor $C^{*}$-category is unitary.
The discovery of quantum groups first arose in physics literature in the work by Faddeev, introduced as noncommutative function algebras. In the mid 80s breakthrough papers by Drinfeld and Jimbo discovered them as quantization of universal enveloping algebras, and developed their properties. We refer to the introduction in [9] and more historical information at the end of each chapter in their book.
After the discovery of the quantum groups in the mid eighties, Doplicher asked the author to try to extend their theory and consequent applications on the construction of a quantum gauge group and a related field algebra. Of course, several studies were dedicated to extend Doplicher-Roberts duality theory to quantum groups, see Ch. IV.5 in [41] for more references on an early stage of this problem, see also the introduction in [11].
Our approach may be regarded in the setting of Connes noncommutative geometry [13], and follows naturally our previous work in the setting of compact quantum groups, see also the introduction in [12], [14]. In [73], [74], [78], [75], [76], [77], [79] we have developed at the level of tensor categories an analogue of Doplicher-Roberts construction of compact gauge group, although the analogy cannot be performed in a strict way due to new difficulties in the quantum case. We shall come back to this, see the mentioned papers and [11]. One important difference was that for rigid braided tensor categories we do not have a notion of braided tensor functor beyond the symmetric ones, because the category of vector spaces does not admit a braided symmetry. Giving up a compatibility property in the symmetric case implies lack of uniqueness of the functor, and thus of the gauge group. This was noticed early as an easy consequence from their construction.
To treat this difficulty in [10] and [11] we have taken into account Drinfeld notion of coboundary, which finally allows for a generalization. Moreover we have followed both the original choice by Mack and Schomerus [61] of a given dimension function corresponding to non-negligible tensor powers of the vector representation of quantum ${\mathfrak{s}l}_{2}$ at roots of unity and also the study of classification of compact quantum groups corresponding to a given representation category of deformation compact quantum groups by Neshveyev and Yamashita [71], [72]. They both considered a specific positive integer valued dimension function associated to the given category of the quantum group, which was the classical dimension function. This suggested to us to consider in [10] and [11] a specific dimension function on a functor related to classical data, and this was first Wenzl functor [87] for the quantum group fusion category, and then the minimum energy functor for the WZW model [93]. In was particularly important for us that in [87] Wenzl describes an explicit quantization curve associated to his functor that leads to an algebra that identifies with the Zhu algebra for that model in the setting of vertex operator algebras. Since in [10] we had constructed a natural quantum group structure on the former compatible with the tensor structure, this arose the question of constructing another natural quantum group structure on the latter at the time of [10], following the quantization curve on the unitary structure of tensor products with the aim of making it classical. This involved in [11] the construction of a Drinfeld twist, a notion that Drinfeld had introduced to study equivalence of what is known as the Drinfeld category with the quantum group category in a formal setting.
We refer to Sect. 23 in [11] for an extension of certain results by Neshveyev and Yamashita in the setting of quantum group fusion categories of type $A$ based on Kazhdan-Wenzl theory at roots of unity in our case, and on the weak Hopf algebras constructed in [10], in place of the compact quantum groups. This section is important for us as it suggested that in general the the associativity morphisms in a category arising from CFT may be of a simple nature from the viewpoint of associated quantum gauge groups that need to be constructed, and this finally proved to be the case, we shall come back to this with more precision.
On the other hand, in the setting of compact quantum groups, while extending Doplicher-Roberts theory starting from embedding part of the morphisms of the category into the Hilbert spaces [73], [75], [79], we had already started to fix a dimension function, and from that we constructed a tensor functor from that subcategory. To consider the full category, the analogue Doplicher-Roberts compact gauge group is described by an ergodic action of a larger quantum group on a quantum homogeneous space, rather than a quantum group itself, due to lack of commutativity of the $C^{*}$-algebra describing the homogeneous space constructed in [75], and this property is likely to hold also in other models arising from conformal field theory, unless one can show that certain known obstructions already met in [74] for compact quantum groups, vanish. We shall come back to this.
However, for the WZW models arising from affine Lie algebras at positive integer levels of the Lie types mentioned in the abstract, we are able to show in [11] that we obtain genuine quantum groups as quantum gauge groups, described as semisimple weak Hopf algebras with $R$-matrix, $C^{*}$-structure explicitly associated to Drinfeld coboundary matrix, and with unitary representations of the braid group, derived from the quantum group fusion category. This weak Hopf algebra has representation category equivalent to both the starting fusion category and also to that of the corresponding affine vertex operator algebra. Constructing the equivalence from the former to the latter as we do in [11] is our proof of Kazhdan-Lusztig-Finkelberg theorem, see Theorem 2.1.
Among our first references in the context of conformal field theory important for us to start to set up the connection with the work of Wenzl in quantum groups, is the work by Wassermann [86] in the setting of affine Lie algebras, where the key concept is played by the notion of primary field, and its initial term, which we use in a fundamental way to connect quantum group fusion category to representations of the Zhu algebra and their special tensor products with the fundamental representation, see Sect 33 in [11]. Most importantly, we describe a connection between Wenzl and Wassermann work via Wenzl quantisation curve in theorem 33.7, to which we refer the reader also for important historical references. The connection from affine Lie algebra setting to Huang-Lepowsky tensor product theory may be found in Sections 35 and 36.
Our abstract duality methods need new further tools as we shall see from the theory of quantum groups, mainly from the work of Drinfeld and Wenzl, that are not available in the theory of groups. However, we shall find that in both contexts, quantum groups at roots of unity and affine vertex operator algebras at positive integer levels, both representation categories are described by functors extending the symmetric tensor functors in Doplicher-Roberts abstract duality theory, see Sect. 27 in [11]. We call these structures hermitian coboundary weak quasi-Hopf algebras, as the hermitian structure is given by Drinfeld coboundary matrix, which is a modification of the $R$-matrix with a square root of the ribbon structure.
We notice on the other hand that a stronger property is needed to establish the connection, than the purely categorical characterisation of functors in Sect. 27. This property is expressed in terms of quantum groups, and the first example, albeit in a non-semisimple setting, is the hermitian structure of $U_{q}({\mathfrak{g}})$ at a root of unity by Theorem 29.9. This property says that the -involution enjoys a precise anticommutation relation (rather than commutation relation as is familiar in Hopf $C^{*}$-algebras) with the coproduct and was first noticed in [87]. We axiomatize this property in section 28, as a compatibility property of hermitian coboundary weak quasi-Hopf algebras. Further examples derived from this examples are our weak Hopf and quasi-Hopf algebras, see Sections 29, 30, 31.
We shall derive even more structure than rigidity e.g. unitarity and modularity on the vertex operator algebra representation category from the corresponding quantum group fusion category, via the equivalence, see Sects. 21, 22, 24 in [11]. Moreover our equivalence, which is given by an analogue of the Drinfeld twist in [23], [24], gives an understanding of the unitary structures of representation categories of affine vertex operator algebras as classical compact real forms, or untwisted. We refer to the outlook section for some more information on possible extension of our viewpoint to models beyond the affine vertex operator algebras at positive integer levels, see also [11].
Our main result is Theorem 2.1. In the subsequent sections we outline the plan of our proof. Our proof starts from the work of Wenzl [87] in the setting of quantum group fusion categories. A main idea that we follow is to consider the fundamental representation $V$ of ${\mathfrak{g}}$ for each Lie type as in [87] where the author explains its generating property and good fusion structure with respect to tensor products with an arbitrary irreducible representation allowing a well-defined truncation procedure to obtain non-negligible objects, with the aim of studying unitarizability of the fusion category. It should also be noted that a main application of that paper is construction of inclusion of subfactors of finite Jones index for all Lie types, extending his previous work.
For all the Lie types we obtain complete identification for the tensor product structure, all the ribbon structure, and those braiding and associativity morphisms where one or two variables are the fundamental representation $V$, respectively. These are parts (a), (b) of Theorem 2.1. The main method of our proof is the construction of semisimple weak Hopf or quasi-Hopf algebras in the two settings and an equivalence of their representation categories via the construction of a Drinfeld twist.
The term weak in the setting of quasi-Hopf algebras was first introduced by Mack and Schomerus motivated by the work of Drinfeld for quasi-Hopf algebras and Drinfeld-Kohno theorem [61]. Weakness refers to a non-unital coproduct, needed to exclude the non-physical representations with zero quantum dimension, and with the need of compatibility with the fusion rules of the category arising from a suitable model in CFT, described by a quantized universal enveloping algebra of ${\mathfrak{s}l}_{2}$ at a root of unity.
Mack and Schomerus did not consider the Hopf counterpart in their theory. Weak Hopf algebras were later introduced by Böhm and Szlachanyi motivated by the work of Mack and Schomerus, but with a strictly coassociative coproduct [5]. The definition of weak Hopf algebra has become standard and understood in the sense by Böhm and Szlachanyi.
In [11], and also in this paper, weak Hopf algebras are understood with a different meaning. While a different name would be helpful to avoid confusion, so far we have not been able to find a good name for our algebras to clarify the situation, and we use the term weak Hopf algebras in the sense defined in Sect 9 in [11].
In [11] we develop a general theory of weak Hopf algebras, review and expand the theory of weak quasi-Hopf algebras, mainly Tannaka-Krein aspects, both in the algebraic and in the $C^{*}$-algebraic setting, with non-trivial unitary structure, discuss a suitable Drinfeld-Kohno theorem, and its application to the construction of semisimple weak Hopf and quasi-Hopf algebras with structure naturally compatible with that of the involved ribbon braided tensor categories, for quantum groups at roots of unity and affine vertex operator algebras at positive integer levels, including unitary structure as heredited from the unitary structure of the category due to Wenzl [87]. We follow the strategy of Drinfeld for Drinfeld-Kohno theorem which includes the comparison of braiding morphisms of the two settings, but we do not include the comparison of the associativity morphisms in our analogue, Theorem 28.4 in [11] (more details are given later). We construct a Drinfeld twist for our weak Hopf algebras useful for the braided symmetry, and study associativity morphisms with a uniqueness theorem using the generating representation. We then use this twist to concretely construct unitary structures in the setting of the representation categories of affine vertex operator algebras at positive integer levels with the tensor structure due to Huang and Lepowsky, for all the mentioned Lie types, by transferring the unitary structure from quantum group fusion category constructed in [87].
Our weak Hopf algebras while having a non-unital coproduct, are not in the class of Bohm and Szlachanyi. For example their coproduct is not strictly coassociative. They are not extension of Bohm and Szlachanyi weak Hopf algebras, and we know from results in [11] that the intersection of the two classes are the genuine Hopf algebras.
On the other hand, in our opinion the weak Hopf algebras considered in [11] still deserve to be regarded as Hopf algebras in a weak sense, for various reasons. On one hand they seem to play the role of the quantized universal enveloping algebras in Drinfeld-Kohno theorem, following the original motivation of Mack and Schomerus. Moreover, their representation category is characterized by a tensor functor to the category of Hilbert spaces with non-invertible natural transformations (projections endowed with natural inverse inclusions, in our case) satisfying tensoriality diagrams, which reduces to the notion of tensor functor when the natural transformations are invertible. Their counit is antimultiplicative and the antipode satisfy the same relations as in the case of Hopf algebras.
We prove the complete identification of braiding and associativity morphisms of the two structures for all the classical Lie types $A$, $B$, $C$, $D$, and the case $G_{2}$ in part (c) of Theorem 2.1, by further refining methods of the previous part but we use more structure in Sect. $8$ of [11] that has not emerged in other parts.
For the type $A$ case, the extra structure that we need to use is known as quantum Schur-Weyl duality between the vector representation and the natural representation of the braid group between the tensor powers of the vector representations.
In the classical case, Schur-Weyl duality is a key property for the general linear groups starting with the work of Schur and Weyl. Schur-Weyl duality also takes an important role in Doplicher-Roberts abstract duality theory for compact groups and applications to construction of the field algebra. This classical duality emergences as the starting point for the construction of global gauge symmetry groups in tensor categories arising in [22].
Duality with the braid group representation is a concept at the basis of the origin of quantum groups following V. F. R. Jones [55].
More generally, for Lie types beyond $A$, we need a corresponding duality between the centralizer algebras of the truncated tensor powers of the fundamental representation $V$ of the Drinfeld-Jimbo quantum group $U_{q}({\mathfrak{g}})$ at the root of unity and the natural representation of the braid group in these algebras, and this leads to the restriction stated in (c). This allows us to apply our abstract uniqueness Theorem 8.19 in [11] for braided tensor structures in a pre-tensor category.
In our work, the need of quantum duality between the braid group representation and the fundamental representation $V$ emerged on the occasion of the proof of our earlier Lemma 7.10, which was not complete in the version 7 of [11], and we worked on completion with the same methods. It was for us a sudden realization of an even greater recognition than we had realized up to that point of the contribution of the authors on extensions of Schur-Weyl duality which has historical interest on his own in representation theory, and more work by Wenzl and his collaborators, and several other authors in the classical and quantum cases for the various Lie types. We shall come back to this in more detail.
In our setting, quantum extension of Schur-Weyl duality to the Lie types has a role to see uniqueness of the associativity morphisms in a tensor category associated to the given braided symmetry with a generating representation $V$, see Theorem 8.19 in [11]. This applies to both fusion categories associated to quantum groups at roots of unity and of affine VOAs.
Presently it is not known to the author whether the fundamental representation for the exceptional Lie types $E$ and $F$ satisfies braid group duality. The $E_{8}$ case is more delicate, because fusion of tensor powers of $U_{q}({\mathfrak{g}})$ at the root of unity is more complicated, as explained in [87], but more work by Wenzl is available for the study of the centralizer algebra [88]. We are grateful to M. Yamashita for pointing out this reference to us.
Among possible applications of our proof are the study of unitary structures of representation categories of vertex operator algebras, due to the clarity offered by a quantum group viewpoint of the same problem as explained in [87]; the study of equivalences between representation categories of conformal nets and vertex operator algebras via the construction of a common quantum gauge group; to extension of Doplicher-Roberts field algebra construction from the local observable algebra to the conformal field theory setting. Some more discussion may be found in the outlook section 13.
2. The main result
Theorem 2.1.
Let ${\mathfrak{g}}$ be a complex simple Lie algebra and let ${\rm Rep}(V_{{\mathfrak{g}}_{k}})$ be the module category of the corresponding affine vertex operator algebra $V_{{\mathfrak{g}}_{k}}$ at level $k\in{\mathbb{N}}$, endowed with the ribbon braided tensor structure introduced by Huang and Lepowsky and let $A(V_{{\mathfrak{g}}_{k}})$ be the Zhu algebra. Then
  • (a)
    The representation category ${\rm Rep}(A(V_{{\mathfrak{g}}_{k}}))$ of the Zhu algebra $A(V_{{\mathfrak{g}}_{k}})$ admits a natural structure of ribbon braided tensor category obtained by transporting Huang-Lepowsky structure to ${\rm Rep}(A(V_{{\mathfrak{g}}_{k}}))$. In this way, Zhu’s linear equivalence
    $Z:{\rm Rep}(V_{{\mathfrak{g}}_{k}})\to{\rm Rep}(A(V_{{\mathfrak{g}}_{k}}))$
    admits a natural structure of ribbon braided tensor equivalence.
  • (b)
    Let $A(V_{{\mathfrak{g}}_{k}})$ be endowed with the unitary compatible coboundary weak quasi-Hopf structure induced by ${\mathcal{C}}({\mathfrak{g}},q)$ via the construction of a weak tensor structure on Wenzl functor and of the Drinfeld twist following Theorem 31.1 in [11] (a), and let ${\rm Rep}(A(V_{{\mathfrak{g}}_{k}}))$ be endowed with the corresponding ribbon braided tensor structure.
    Then ${\rm Rep}(A(V_{{\mathfrak{g}}_{k}}))$ becomes a unitary coboundary modular tensor category in this way. Moreover the transport in (a) preserves the tensor products, the ribbon for all objects, the braid morphisms for pairs $(V_{\lambda},V)$ and $(V_{\lambda},V)$, and the associativity morphisms for triples $(V_{\lambda},V,V)$, $(V,V_{\lambda},V)$,$(V,V,V_{\lambda})$ with $V_{\lambda}$ an arbitrary irreducible of the fusion category.
  • (c)
    If ${\mathfrak{g}}$ is of one of the Lie types $A$, $B$, $C$, $D$, $G_{2}$ then the transport preserves the braided tensor structure for all objects. Thus $Z$ is a ribbon braided tensor equivalence. Specifically, the composition
[…]
    is a ribbon tensor equivalence by application of the indicated theorems.
The definition of unitary coboundary weak Hopf algebra, its restricted version named unitary coboundary weak Hopf algebra, are given in Sects. 29, 30 in [11], by abstracting features of the hermitian structure of $U_{q}({\mathfrak{g}})$ for $|q|=1$, discovered by Wenzl in [87] in his proof of unitarity of Kirillov inner product, for the associated fusion category as certain roots of unity, based on a complex Drinfeld coboundary matrix. The former more general notion admits a categorical characterization generalizing Doplicher-Roberts symmetric functors.
We would like to notice that the passages described in part (c) of Theorem 2.1 are explicit. The first and the last equivalences are due respectively, to the work by Zhu combined with that of Huang and Lepowsky. The last is the functor due to Wenzl enriched with our weak tensor structure, preserving Drinfeld coboundary structure derived from the ribbon and $R$-matrix, see the above cited sections in [11].
The middle equivalence in (2.1) is application of the explicit symmetry provided by our analogue of Drinfeld-Kohno theorem 28.4 in [11]. This symmetry is given by a Drinfeld twist explicitly defined by the action of the $R$-matrix, more precisely by the braided and ribbon structure in the two settings, that allow to see in the vertex operator algebra setting same structure as that in the quantum group setting.
3. Motivation and framework of our work. Constructing braided tensor equivalences via a quantum gauge group given by a discrete weak Hopf algebra and a Drinfeld twist.
Before going into the explanation of the statement and outline the main tools of the proof in more detail, we summarize the general framework that motivated and led to our work. In the following, where not indicated the cited sections and results refer to [11], and also for unexplained notation.
We were originally motivated by Doplicher-Roberts abstract duality theory for compact groups and their application to the construction of a global compact gauge group and reconstruction of the net of field algebras from the observable algebras in high dimensional quantum field theories [21], [22].
In the setting of low dimensional quantum field theories, starting with the fusion categories of quantum groups at all the roots of unity that are known to correspond to unitary structures by the work of Wenzl and Xu [87], [92], we in particular construct discrete quantum group-like structures with non-trivial twisted unitary structures for all the Lie types as in [87]. These discrete quantum groups are thought of as duals of compact quantum groups, but their preduals admit no analogy to the case of the compact quantum groups of Woronowicz, because they are non-associative algebras. However, Woronowicz work was strongly inspiring for us. Indeed our first examples in the type $A$ case were first constructed in our previous work [10] via an abstract Tannaka-Krein method in the search of a noncommutative function algebra, using the fusion rules of the corresponding fusion category.
On the other hand, their duals are quite well-behaved discrete $C^{*}$-algebras with almost associative coproducts, antimultiplicative antipode and multiplicative counit satisfying the same relation as for Hopf algebras. In this sense they resemble the duals of the compact quantum groups, the main differences being non-unitality of the coproduct and the unitary structure of the tensor product of their representations, which is twisted via a positive operator associated to the $R$-matrix (Drinfeld coboundary matrix), and makes the representation of the braid group unitary. The need of a twist for the tensor product representation arises from the fact that coproduct and -involution satisfy an anticommutativity condition, rather than commutativity as it usually is the case for Hopf algebras in the operator algebra setting. This anticommutativity is derived from the same property of the -involution of the non-semisimple $U_{q}({\mathfrak{g}})$, first pointed out by Wenzl [87].
The weak Hopf algebras constructed in [10] have been extended to all Lie types in Sections 29, 30 of [11] with uniform methods for all the Lie types.
Our algebras are examples of Mack-Schomerus weak quasi-Hopf algebras introduced in [61] as generalization of Drinfeld quasi-Hopf algebras with non-unital coproduct. For ${\mathfrak{g}}={\mathfrak{s}l}_{2}$ our examples reduce to Mack-Schomerus examples for the algebraic structure. We understand their lack of coassociativity as quite mild, because their representation category is endowed with a weak unitary tensor functor in the category of Hilbert spaces. Thus we regard them as the Hopf counterpart of weak quasi-Hopf algebras, in a position similar to that of Hopf algebras among quasi-Hopf algebras, that in our knowledge has been overlooked.
Here the term weak is understood in the sense first introduced by Mack and Schomerus, the coproduct is non-unital, a feature needed to exclude the non-physical representations with zero quantum dimension, and with the need of compatibility with the tensor structure of the category. The weakness of tensoriality of the embedding functor of their representation category to the category of Hilbert spaces means that the structure natural transformations are surjections onto subspaces of the tensor product spaces endowed with left inverses, rather than isomorphisms, and satisfy two tensoriality diagrams.
Much of Woronowicz theory can be extended to our case. From this via a quantization isomorphism described by Wenzl in [87] and the construction of a Drinfeld twist we construct and describe in some more detail later a quantum group-like structure on the Zhu algebra associated to the affine models described as in the statement of the main result.
These constructions are in some quantum analogy with Doplicher-Roberts construction in the setting of high dimensional algebraic quantum field theory, in that the categories that we study are not given as Tannakian categories. In our case, differently from their case, we have naturally defined linear functors emerging as the minimum energy functor due to Zhu in the setting of vertex operator algebras. Also differently form their case, where the linear functor needed to be constructed but the tensor structure was trivial, in our case we need to construct tensor structures associated to these naturally occurring linear functors to the category of vector spaces.
Our starting point is what we shall refer to as Wenzl functor associated to the quantum group fusion category described in [87]. This paper plays a central role, both on the side of quantum groups and on the side of affine vertex operator algebras, as we next try to explain.
As said, on the the side of quantum groups, our approach has been influenced by results in quantum group theory by Woronowicz but also by the dual viewpoint by Majid for quasi-Hopf algebras, extended to weak quasi-Hopf algebras by Haring-Oldenburg. We start reviewing these Tannakian reconstruction theorems from the viewpoint of Majid and Haring-Oldenburg in Sect. 5 in [11], where the reader may find indication of the original references. Moreover Kazhdan and Wenzl theory is also important to us, and the reader may consult Sect. 23 in [11] and references therein. It is by no means an exaggeration to assert that our paper [11] would have not existed without the paper by Neshveyev, Tuset on Kazhdan-Lusztig theorem for compact quantum groups [68]. This paper was strongly inspiring for our strategy, although methods are different. For similar reasons, we were also guided by the presentation of Drinfeld-Kohno theorem in the book by Bakalov and Kirillov [1]. These last two references first inspired our work in the search of a unified treatment following a Tannakian approach in the idea of Majid, in our case. An important aspect in our case as far as Majid Tannakian duality for quasi-Hopf algebras, is the search of natural tensor structures for the involved embedding functors into the category of Hilbert spaces (Wenzl functor and Zhu functor in the two settings).
Primary fields first emerged in the physics literature, [3], using previous work by Feigin and Fuks [29] and Kac [56] on the representation theory of the Virasoro algebra. In 1984 Witten introduced what are now known the WZW models, due to Wess, Zumino, Novikov and Witten [90]. In the same year, Knizhnik and Zamolodchikov showed that their correlation functions satisfy a system of partial differential equations [59].
Tsuchiya and Kanie started a rigorous study in [84]. Goddard, Kent and Olive realized the chiral unitary Virasoro minimal models via a coset construction [37].
Moore and Seiberg conjectured some analytic properties of the quantum fields and described the construction of a modular category associated to a general conformal field theory.
The mathematical contribution by Kohno, Tsuchiya and Kanie [84], and the subsequent works by A. Wassermann [86] and Toledano-Laredo [83] on affine Lie algebras, compared with the structure that had previously emphasized by Wenzl in the study ofquantum group fusion categories and their unitary structures [87] that led to our approach. The reader may find reference to more of the several original papers in mathematics and physics in Sections 31–36 in [11].
4. The general plan of our proof.
We briefly explain here the plan of our proof of the main theorem for parts (a), (b), and (c). In the following sections we describe the main ideas to prove these parts in greater detail.
Part (a) is an application of a general construction of a natural structure of a (ribbon, braided) tensor equivalence on a given equivalence ${\mathcal{C}}\to{\rm Rep}(A)$ of semisimple linear categories endowed with an inverse linear equivalence ${\rm Rep}(A)\to{\mathcal{C}}$, with $A$ a discrete algebra. We outline the main ideas of this tensor structure construction in the next section. The reader may refer to Section 15 in [11] for the detailed construction.
For part (b), we follow a suitable Drinfeld twist method originally used by Drinfeld for Drinfeld-Kohno theorem [23], [24], [25], and in our case for different categories, defined in the statement of the main result. This is the main part of the result. The first main step is the construction of a class of finite dimensional weak Hopf $C^{*}$-algebras $A_{W}({\mathfrak{g}},q)$ for all Lie types and positive integer levels, extending the former examples of [10] in the type $A$ case. They have unitary coboundary structures associated to the fusion category of quantum groups at roots of unity, associated to a specific linear fiber functor $W:{\mathcal{C}}({\mathfrak{g}},q)\to{\rm Vec}$, that emerges from the work of Wenzl [87]. Complete definitions of unitary coboundary structures on weak quasi-Hopf algebras are given in [11], and are motivated by the work of several authors, among them Drinfeld, Wenzl, Mack and Schomerus. See Sections 3, 5, 9, 10, 11, 26, 28 in [11].
In Sect. 8 of [11] we prove two main abstract uniqueness results for associativity morphisms and braided symmetries in a semisimple pre-tensor category with a generating object. These are Theorem 8.13 and its extension Theorem 8.19. We explain here briefly how Theorem 8.19 applies to imply part (c) of Theorem 2.1.
The proof of parts (b) and (c) is based on the general theory of semisimple weak quasi-Hopf algebras, weak Hopf algebras, and the theory of Drinfeld twist treated in Sections 3–6, and 8–16 in [11]. This abstract theory splits in two parts. Theory on which part (b) of Theorem 2.1 is based is developed in Sections 3–6, and 9–16. The applicative part of the proof of part (b) is developed in Sections 19–36. The abstract aspects that lead to the proof of part (c) of Theorem 2.1 are the main subject of Section 8.
For part (b), we develop an abstract analogue of Drinfeld-Kohno theorem, Theorem 28.4, and apply it to the setting of fusion categories of quantum groups at suitable roots of unity and of module categories of affine vertex operator algebras at positive integer levels. We construct weak Hopf algebras $A_{W}({\mathfrak{g}},q)$ associated to the quantum group fusion category for all the Lie types in Section 30, by constructing a weak tensor structure on a linear functor from this category to the category of vector spaces, that emerges in [87] for the special tensor products described in (b). The weak tensor structure is constituted by fusion intertwiners with respect to the action of $U_{q}({\mathfrak{g}})$.
By application of our analogue of Drinfeld-Kohno theorem we obtain the structure of a weak quasi-Hopf algebra on the Zhu algebra $A(V_{{\mathfrak{g}}_{k}})$ associated to a corresponding affine vertex operator algebra $V_{{\mathfrak{g}}_{k}}$. Thus a natural weak quasitensor structure defined by fusion intertwiners with respect to the action of ${\mathfrak{g}}$ compatible with the fusion rules of modules of the vertex operator algebra. This induces the structure of a ribbon braided tensor category on the module category of $V_{{\mathfrak{g}}_{k}}$, via transport from the the Zhu algebra $A(V_{{\mathfrak{g}}_{k}})$.
By construction, this structure is ribbon braided tensor equivalent to the structure arising from the quantum group fusion category. The equivalence is explicitly described by a Drinfeld twist analogous to the Drinfeld twist appearing in Drinfeld’s first proof of Drinfeld-Kohno theorem [24]. In our case, the unitary structure of the fusion category of quantum groups at roots of unity in [87] is used in the construction of the Drinfeld twist of $A_{W}({\mathfrak{g}},q)$.
On the module category of $V_{{\mathfrak{g}}_{k}}$ we also have the ribbon braided tensor structure constructed by Huang and Lepowsky, and we need to compare the two structures.
5. Outline of the construction of a weak quasi-tensor structure on the Zhu algebra $A(V_{{\mathfrak{g}}_{k}})$ via Wenzl quantisation isomorphism and Drinfeld twist method.
Let $({\mathcal{C}},\otimes,\iota,\alpha)$ be a semisimple complex tensor category (possibly braided, rigid, ribbon) and let
${\mathcal{F}}:{\mathcal{C}}\to{\rm Vec}$
be a faithful linear functor. There are two natural ways to make the representation category ${\rm Rep}(A)$ of $A={\rm Nat}_{0}({\mathcal{F}})$ into a tensor category (or a category with the same type of structure as ${\mathcal{C}}$) equivalent to ${\mathcal{C}}$ with all the structure.
The first way is traditionally referred to as Tannaka-Krein duality. One needs a (weak, quasi-) tensor structure $(F,G)$ on ${\mathcal{F}}$, and then apply Theorem 6.6 to obtain a (weak, quasi)-bialgebra structure on $A={\rm Nat}_{0}({\mathcal{F}})$ (with extra structure) with the desired equivalence property. The (weak quasi-) tensor structure
$({\mathcal{F}},F,G):{\mathcal{C}}\to{\rm Vec}$
on a given functor, may be easily identified in some cases, e.g. representation categories of groups, simple complex Lie algebras, or quantum groups or more generally whenever one has a tensor category that naturally admits a tensor functor to ${\rm Vec}$, as is the example of the compact quantum groups of Woronowicz. In these cases, Tannakian constructions are ubiquitous.
But in other cases the structure $(F,G)$ needs to be constructed, and in this case the situation is perhaps similar to the problem met by Doplicher and Roberts and Deligne in their abstract duality theories. In Doplicher-Roberts duality for compact groups, the embedding functor itself ${\mathcal{F}}$ needed to be constructed from an abstract symmetric strict tensor category satisfying suitable conditions, and the tensor structure of ${\mathcal{F}}$ was trivial.
The Tannakian construction endows $A$ with an associator which depends on both the associativity morphisms $\alpha$ of ${\mathcal{C}}$ and the structure $(F,G)$, following Majid version for Drinfeld quasi-Hopf algebras. Thus Tannaka-Krein duality is particularly useful in cases where the associativity morphisms $\alpha$ are not too complicated, in that in this case it puts emphasis on $F$ and $G$. This may allow to understand most of the associativity property of the category by understanding $F$ and $G$.
Fusion categories that arise from quantum groups at roots of unity, by the work of many authors, among them Drinfeld, Jimbo, Lusztig, Andersen, Wenzl, are realized in a natural way as strict fusion categories. However, in this case tensor functors to ${\rm Vec}$ do not exist. A natural linear functor has been naturally constructed in [87] for the roots of unity of interest for the aims of [11], i.e. those that are related to the affine Lie algebras at positive integer levels.
Tensor structures on Wenzl functor ${\mathcal{F}}$ do not exist in general, and one has to construct more general structures, from the inner structure of the category, in order to reduce to the weak quasi-bialgebra method.
A suitable notion of weak Hopf algebra property is a natural concept to look for, by the analogy with Drinfeld-Kohno theorem that first motivated the work of physicists Mack and Schomerus in [61]. The adjective weak refers to the need of considering truncated fusion rules that respect fusion rules of physical models, and rule out representations of $U_{q}({\mathfrak{g}})$ at roots of unity of vanishing quantum dimension. They accordingly introduced the notion of weak analogues of Drinfeld quasi-Hopf algebras, and called them weak quasi-Hopf algebras, a key concept that we rely on. They constructed a semisimple weak quasi-Hopf algebra associated to a suitable model in conformal field theory for ${\mathfrak{s}l}_{2}$ and truncated fusion rules. However, the weak Hopf property was overlooked by them.
We introduce a notion of weak Hopf algebra in [11]. This property was first noticed in [10] in the type $A$.
In Sect. 30 we shall construct a weak tensor structure on Wenzl functor, and derive a weak Hopf finite dimensional $C^{*}$-algebra $A_{W}({\mathfrak{g}},q)$ associated to fusion categories of quantum groups at roots of unity for all Lie types, with extra structure, that is the braiding, unitary structures given by a Drinfeld coboundary matrix, also rigid and ribbon, extending the work of [10] for ${\mathfrak{s}l}_{N}$ in turn extending and expanding the work by Mack and Schomerus for ${\mathfrak{s}l}_{2}$ [61].
We use the weak Hopf algebra property of $A_{W}({\mathfrak{g}},q)$ as a replacement of the Hopf algebra property of $U_{h}({\mathfrak{g}})$ in Drinfeld-Kohno theorem. This property indeed plays a crucial role in our work, but in a different way as compared to Drinfeld’s proof of Drinfeld-Kohno theorem. In our case, we use the weak Hopf property jointly with fusion and duality properties of a fundamental generating representation $V$ in the application, to study the same aim as Drinfeld, comparison of the braided symmetries and associativity morphisms in the two settings of fusion categories of quantum groups at roots of unity and of module categories of vertex operator algebras.
The unitary aspects of our weak Hopf algebras $A_{W}({\mathfrak{g}},q)$ are based on the results of [87], on the use of Drinfeld coboundary matrix.
The weak tensor property as compared to the weak quasi-tensor property may prove useful for further constructions of the quantum fields from the observables, with the aim of extending the work of Doplicher and Roberts in high dimensional algebraic quantum field theory, as compared to previous attempts in the literature on low dimensional algebraic quantum field theory, that suffered the difficulties due to a possibly more complicated associator.
6. Main tools to prove (a): How to make ${\rm Rep}(A({{V_{\mathfrak{g}}}_{k}}))$ into a braided tensor category with Huang-Lepowsky structure.
A second way to construct a braided tensor category structure on ${\rm Rep}(A)$ from a braided tensor category ${\mathcal{C}}$ is useful when ${\mathcal{C}}$ has a natural faithful linear functor
${\mathcal{F}}:{\mathcal{C}}\to{\rm Vec}$
given from the start and a braided tensor category structure $({\mathcal{C}},\otimes,\iota,\alpha,d)$ is constructed as a second step. Then in this case it may be more natural to first consider the discrete algebra $A={\rm Nat}_{0}({\mathcal{F}})$ and then transfer of all the structure of ${\mathcal{C}}$ to the representation category of $A$.
We discuss this transfer in Sect. 15 without the need of a weak quasi-tensor structure on ${\mathcal{F}}$, in abstract terms. One only needs an inverse linear equivalence ${\mathcal{S}}:{\rm Rep}(A)\to{\mathcal{C}}$ of the Tannakian linear equivalence ${\mathcal{E}}:{\mathcal{C}}\to{\rm Rep}(A)$. In this situation we may define a braided tensor category structure in ${\rm Rep}(A)$ from the braided tensor category structure in ${\mathcal{C}}$ with the property that ${\rm Rep}(A)$ is equivalent to ${\mathcal{C}}$ as a braided tensor category. The tensor product, associativity morphisms, braiding morphisms of ${\rm Rep}(A)$ are explicitly defined via the equivalence ${\mathcal{E}}$ and its inverse ${\mathcal{S}}$.
Far-reaching historic examples of such ${\mathcal{F}}$ may be found in conformal field theories (e.g. in the setting of models of conformal nets and in that of vertex operator algebras) as the minimum energy functor, before the braided tensor category structure was constructed. In these cases we refer to the braided tensor tensor category structure of modules of vertex operator algebras by Huang and Lepowsky or to the unitary braided tensor category structure of localized endomorphisms of conformal nets.
In the setting of vertex operator algebras, the minimum energy functor is also known as Zhu’s functor [93]. Notably, Zhu constructed a semisimple algebra $A(V)$ directly from modules of the vertex operator algebra $V$, naturally isomorphic to ${\rm Nat}_{0}({\mathcal{F}})$. This algebra is well known, and is referred to as the Zhu algebra. He also constructed a linear equivalence ${\mathcal{E}}:{\rm Rep}(V)\to{\rm Rep}(A(V))$ and also a natural linear inverse equivalence ${\mathcal{S}}:{\rm Rep}(A(V))\to{\rm Rep}(V)$ in the same paper if the vertex operator algebra $V$ satisfies suitable semisimplicity conditions.
Section 15 abstracts Zhu linear equivalence and its inverse, by considering a pair of the Tannakian linear equivalence ${\mathcal{E}}:{\mathcal{C}}\to{\rm Rep}(A)$ associated to ${\mathcal{F}}:{\mathcal{C}}\to{\rm Vec}$ and a linear inverse equivalence ${\mathcal{S}}:{\rm Rep}(A)\to{\mathcal{C}}$ and describes how to make ${\rm Rep}(A)$ into a braided tensor category.
Thus, in the case of vertex operator algebras satisfying semisimplicity conditions, one may reduce the analysis of the category of infinite dimensional representations of the vertex operator algebra $V$ to a category of finite dimensional representations of the Zhu algebra $A(V)$ in a natural way. This transfer is the passage described in part (a) of the statement of Theorem 2.1 for the affine vertex operator algebras at a positive integer level $k$.
We show in [11] that a weak quasi-tensor structure $(F_{0},G_{0})$ can be associated to the minimum energy functor ${\mathcal{F}}$ in the case of affine vertex operator algebras at positive integer level. In other words at the end we shall construct natural weak quasi-tensor structures $({\mathcal{F}},F_{0},G_{0})$, associated to the weak tensor structure of Wenzl functor of a corresponding fusion category of a quantum group at a root of unity. Since the latter is a rigid and modular tensor category, we in particular directly obtain rigidity and modularity of ${\rm Rep}(V_{{\mathfrak{g}}_{k}})$ from the corresponding quantum group fusion category. In this way we rediscover a result by Huang for the special case of affine vertex operator algebras at a positive integer level [51].
In the final sections of [11] we discuss in greater detail how Huang and Lepowsky theory together with previous work by Frenkel and Zhu provides ${\mathcal{F}}$ naturally with the weak quasi-tensor structure $(F_{0},G_{0})$, such that $F_{0}$ is the 0-grade part of Huang-Lepowsky tensor structure map $F$.
We shall first identify a quantum group fusion category at a fixed suitable root of unity with that of a corresponding affine VOA fusion category or conformal net fusion category as pre-tensor categories, then we identify them with all the structure on special variables which do not generate all the variables linearly. By analysis of a cohomological kind on any braided tensor category with a generating representation, and further braid group duality properties well studied in the literature for quantum groups, we shall extend this identification to all variables. We shall give a more detailed overview of our proof of Theorem 2.1 in the following sections.
7. A central method in the quantum group and VOA setting: the fundamental generating representation $V$ for all Lie types, following H.Wenzl.
A central method of proof for both parts (b) and (c) is the use of the fundamental generating representation $V$ and its properties, following primarily the work of Wenzl. The paper [87] plays a main role especially for the proof of (b). For part (c) we shall later describe results and related works in the literature on which our proof is based.
Part (b) of Theorem 2.1 gives coincidence of the two braided tensor structures for all Lie types (two pairs of associativity morphisms together with their braiding morphisms satisfying the two hexagonal diagrams) on a specific collection of their variables, in the two settings of quantum groups and vertex operator algebras. These collections do not generate additively all the variables of the braiding morphisms or associativity morphisms, because they do not contain all the irreducible objects.
Coincidence of the associativity morphisms and braiding morphisms on the special objects is verified in a direct way, relying on similar structural properties of fusion of two corresponding generating representations in the two cases. In the quantum group case, fusion arises in the setting of non-semisimple representations of $U_{q}({\mathfrak{g}})$ considered in [87], while in the vertex operator algebra case fusion arises in a semisimple setting. To compare the braiding and associativity morphisms on the special variables, we use a Drinfeld twist that we construct trivialising, in a sense, the non-trivial unitary structure in the setting of representations of quantum groups at roots of unity to that of the affine vertex operator algebra. This idea emerges in the construction and analysis of the unitary structure of the fusion category of quantum groups at roots of unity of [87].
The twist takes one unitary structure of a special tensor product arising in the quantum group setting from [87] to the classical unitary structure corresponding to the compact real form of ${\mathfrak{g}}$ (and this in turn remarkably leads to the construction of a unitary tensor structure on the module category of the vertex operator algebra $V_{{\mathfrak{g}}_{k}}$ in a uniform way). This implies that the twist leaves the corresponding structural properties invariant on these special variables. To see this, we work on the side of the vertex operator algebras and we compare with the work by Frenkel and Zhu [34]. An abstract concept describing these common structural properties of the associativity morphisms is introduced in Sect. 8 with the name of a ${\mathcal{V}}$-pre-associator of CFT type. The name is motivated by our understanding of the structure of Huang-Lepowsky associativity morphisms.
Direct verification of coincidence of the two braided tensor structures on all the irreducible objects is more complicated due to multiplicity of fusion and non-semisimplicity of $U_{q}({\mathfrak{g}})$. To this aim rather than pursuing a closer analogy with the original Drinfeld-Kohno theorem for all objects, we develop two uniqueness theorems, Theorem 8.13 and its generalization Theorem 8.19. This generalization is made necessary by the actual conclusions reached in our work in the applicative sections on the comparison between special cases of the two associativity morphisms and their corresponding braided symmetries. Thus the uniqueness theorem 8.19 is formulated with suitable assumptions, where again in the application the abstract notion of ${\mathcal{V}}$-pre-associator of CFT-type for the fundamental representation $V$ is used in a central way both on the side of quantum groups and that of vertex operator algebras to verify coincidence of these special braid and associativity morphisms, after application of our Drinfeld-Kohno theorem. But this brief description of Theorems 8.13 and Theorem 8.19 put together with coincidence of the two braided tensor structures on special collections of objects proved in part (b), is still not enough to describe our method of proof to complete the proof of part (c) and with it the proof of Theorem 2.1.
We need more that we describe in the following sections. The first next step is a somewhat complicated cohomological reduction work done in Theorems 8.13 and 8.19, where we use cohomology of the pentagon equation and its interplay with the hexagonal equations defining the braided symmetry for a general semisimple braided tensor category endowed with a generating object.
The cohomological reduction is useful to reduce the problem to the last step, which uncovers a central duality property between the fundamental representation of the quantum group and corresponding representations of the braid group in the intertwining spaces of the fundamental representation, in the various Lie types. The first example of this is Schur-Weyl duality in the type $A$ case, and works in the literature that we shall recall greatly extend this first classical example to to other Lie types and to the quantum cases. We shall come back to this, but let us anticipate similarities and differences with the original Drinfeld-Kohno theorem, useful for the proof of (b), that will also useful to clarify how we were led to uncover the key role of duality of the fundamental representation $V$, for the proof of the last step (c).
8. Main tools to prove (b): Analogies and differences between Drinfeld-Kohno Theorem and Theorem 28.4 in [11]
We next outline general ideas that make our strategy possible, useful to prove part (b) of Theorem 2.1.
As already explained, the starting point of our work is the paper [87]. In that paper, a fundamental generating representation $V$ is described for each simple Lie algebra ${\mathfrak{g}}$, and it is used for the construction of the unitary structure in the associated quantum group fusion category. In our approach we follow the work of Wenzl on the use of the generating representation $V$.
Secondly, in our approach the notion of coboundary matrix first introduced by Drinfeld in his work on Drinfeld-Kohno theorem for quasi-Hopf algebras in a formal setting, plays a major two-fold role. In our case we adopt the idea of Drinfeld coboundary that we describe for complex weak quasi-Hopf algebras in an analytic setting. This general theory is considered in Sections 26, 27, 28.
On one hand the coboundary matrix in an analytic setting induces a unitary structure which shows striking differences with unitary structures of usual Hopf $C^{*}$-algebras, such as for example the compact quantum groups, where the -involution commutes with the coproduct. Drinfeld original coboundary is used in [87] in the setting of representations of the quantized universal enveloping algebra $U_{q}({\mathfrak{g}})$ with $q$ a suitable type of root of unity, to construct the unitary braided tensor structure of the associated fusion category. This already gives a striking connection between unitary structures and ribbon braided tensor equivalences established by Drinfeld in Drinfeld-Kohno theorem, for different categories.
Similarities with the original Drinfeld-Kohno theorem regard the treatment of special braiding morphisms, that is based as in Drinfeld proof on a twist given by a square root of the Drinfeld coboundary matrix. This approach is particularly useful for braiding on tensor products of the form $V_{\lambda}\otimes V$ or $V\otimes V_{\lambda}$, where multiplicity-free and structural fusion properties of $V$ allow for a direct verification of the equivalence between the two braided symmetries on these objects. Moreover in our case on special triples (or $n$-tuples) where two (or consecutive $n-1$-tuples) coordinates are fixed to $V$, verification of the equivalence of the two associativity morphisms similarly put in a nutshell reduces to the same properties of $V$. We do not use the KZ equation to this aim, or in any part of our work, other than the fact that the KZ equation was used by Drinfeld, Kazhdan and Lusztig, and by Huang and Lepowsky in the setting of affine vertex operator algebras, to construct the structure of a braided tensor category.
9. Differences between the original Drinfeld-Kohno theorem and our analogue Theorem 28.4 in the treatment of the associativity morphisms, to prove (c).
Differences regard mainly the treatment of the associativity morphisms, that in our case as before explained resorts to determining the braiding morphisms via an analogue of Drinfeld square root construction of the coboundary matrix, and to braid group duality for $V$ and to cohomology of a braided tensor structure of a generating representation. Thus in our case the method of proof for coincidence of the associativity morphisms depends on the braid group action primarily and categorical cohomological methods of a kind perhaps different than those used by Drinfeld in the original Drinfeld-Kohno theorem. It would be interesting to study whether the two methods are related.
Differently from the original Drinfeld-Kohno theorem, we do not study the problem of equivalence associativity morphisms in our version version of Drinfeld-Kohno theorem. We do this in Section 8 in [11] with the mentioned different ideas that appeal to duality properties of the fundamental representation $V$ with respect to the representation of the braid group in the tensor powers of $V$.
More specifically, we first construct semisimple weak Hopf algebras $A_{W}({\mathfrak{g}},q)$ associated to fusion categories of quantum groups at roots of unity for all Lie types. We show that $A_{W}({\mathfrak{g}},q)$ satisfy a triviality property of the associator, which makes the definining weak quasi-tensor functor (via a suitable Tannakian duality construction) into a weak tensor functor.
These $A_{W}({\mathfrak{g}},q)$ play the role of the quantum group over a formal variable in Drinfeld-Kohno theorem. This construction is based on the analysis in [87] and the triviality property of the associator is based on properties observed by Gelfand and Kazhdan for morphisms between non negligible and negligible objects [36], following the same ideas as in [10] in the type $A$ case.
Then we use a quantization continuous curve that is described in [87] to connect irreducible representations of $A_{W}({\mathfrak{g}},q)$ with irreducible representations of the Zhu algebra $A({V_{\mathfrak{g}}}_{k})$ associated to a corresponding affine Lie algebra, or vertex operator algebra at a positive integer level.
We apply the natural isomorphism given by extremes of Wenzl continuos curve defining quantization of a suitable finite class of representations of the classical Lie algebra (the Weyl alcove defined by the roots of unity) to representations of $U_{q}({\mathfrak{g}})$. Then we apply our Drinfeld-Kohno theorem to endow the Zhu algebra with the structure of a weak quasi-Hopf algebra with coboundary matrix.
These two passages, in analogy with the original Drinfeld-Kohno theorem, allow to construct the structure of a weak quasi-Hopf algebra in our case on the Zhu algebra $A({V_{\mathfrak{g}}}_{k})$. At this point we can use Zhu’s linear equivalence and transport the ribbon braided tensor structure to the module category of the associated affine vertex operator algebra. As already explained we verify that braided symmetry and associativity morphisms correspond to the correct braided symmetry and associativity morphisms in the setting of vertex operator algebras on special pairs or triples of representations of the weak quasi-Hopf algebra.
The unitary structure associated to a coboundary of a semisimple weak quasi-Hopf algebra allows directly unitary representations of the braid group, and, similarly to the case of $U_{q}({\mathfrak{g}})$, anticommutes with the comultiplication. This property resembles the antimultiplicativity property of the -involution in a noncommutative topological space. The unitary structure enters in the construction of the tensor equivalence in our case via an analytic construction of a Drinfeld twist first performed by Drinfeld for the Drinfeld category.
Drinfeld coboundary is closely related to the ribbon structure, and thus it is relevant for the ribbon preserving property of the KLF equivalence, similarly to the case of Drinfeld-Kohno theorem.
10. Main tools to prove (c): cohomological reduction of a braided tensor structure of a category with a generating object $V$ to braid group duality of $V$
Theorem 8.13 and Theorem 8.19 show uniqueness of the associativity morphisms (and also braiding morphisms in the latter case) under some assumptions. These assumptions all aim to determine the associativity morphisms uniquely via data from the braiding, but lead to analysis of a different nature in the application.
The first assumption requires knowledge that in our application, via the methods previously alluded to (i.e. Wenzl isomorphism given by a continuous curve describing quantization of all the irreducible representations of the fusion category [87] and recalled also in [11], combined with application of our Drinfeld twist) one gets a braided symmetry coinciding, for all pairs of objects, with the braided symmetry associated to Huang-Lepowsky braided tensor category structure.
Moreover in Theorem 8.13 we require that the associativity morphisms coincide on the special triples. These assumptions allow a cohomological analysis on the braided symmetry of a tensor category with a generating object. This analysis does not suffice to complete the proof Theorem 8.13. Before describing the further key property of the fundamental representation that we need, we add some more comments on our methods regarding this part.
In the application on the other hand, while as discussed in the previous parts we reach coincidence of associativity morphisms on special triples thanks to the structure of Huang-Lepowsky associativity morphisms and Drinfeld twist method, regarding the braiding morphisms we limit ourselves to verify the coincidence of the two braided symmetries on the special pairs of objects where one of the variables is fixed to $V$. These special braiding morphisms generate all the other braiding morphisms, but the generating property of these special morphisms depends also on the associativity morphisms. This poses the need of a more detailed analysis to obtain a uniqueness theorem for both braiding and associativity morphisms based on knowledge of their coincidence on special variables that in our approach we can verify directly. This is the aim of Theorem 8.19, with respect to Theorem 8.13, which on the other hand is useful to present the main ideas of how a full braided symmetry determines the associativity morphisms in case of a generating representation.
The assumptions on coincidence of the associativity morphisms and braided symmetry as required in Theorem 8.19 are verified in the proof of Theorem 2.1 on the side of vertex operator algebras. For the associativity morphisms, methods are similar to those used to verify the assumption of Theorem 8.13. We apply Wenzl isomorphism and Drinfeld twist construction and then we use the structural properties of the Huang-Leposwky associativity morphisms in Theorem 31.9. For the special braiding morphisms we take into account the formulas given by Toledano-Laredo in Theorem 31.12. These two comparisons first motivated our method of proof for both braided symmetry and associativity morphisms.
The proof of (c) of Theorem 2.1 is reduced to the proof of (b) and application of Theorem 8.19. More in detail, for part (c) we use the fact that the special variables of the associativity morphisms and braided symmetry generate all the triples multiplicatively, in the sense that all the components except one in the variables of the associativity and braiding morphisms are taken by the generating representation in each Lie type. In the proof of Theorem 8.19, and the antecedent Theorems 8.13, we explain the generating property of these special values of the associativity and braiding morphisms with the pentagon equation of associativity morphisms and the hexagonal equations of a braided symmetry in a semisimple tensor category with a generating object but, as anticipated, another important property concerning the braid group representations in the category is needed, to determine the associativity morphisms uniquely.
11. The role of braid group duality for the generating object in determining the full braided tensor structure for the proof of (c).
The uniqueness Theorems 8.13 and 8.19 both apply under a central duality property of the fundamental representation $V$ with respect to the braid group representation in the quantum group tensor category, which is at the core of quantum group representation theory. The first instance of this property is Schur-Weyl duality for the general linear group. This duality says that the algebra of intertwining operators, from a tensor power of the vector representation $V$ of the general linear group to itself (also called centralizer algebras for the tensor powers of the vector representation) is generated by permutation operators [91].
The generalization of Schur-Weyl duality to Lie types beyond $A$ in classical Lie groups and quantum groups is an enormous subject developed by many authors, extending from the classical more familiar cases of type $A$ and vector representation to other Lie types beyond $A$, with respect to a specified generating representation $V$, and to the quantum group case.
We need the property that the centralizer algebras $(V^{n},V^{n})$ with respect to the action of the Drinfeld-Jimbo quantum group $U_{q}({\mathfrak{g}})$ are generated by the natural representation of the braid group induced by the $R$-matrix. In our case $q$ is a root of unity as usually assumed corresponding to the unitary cases of the associated fusion category, specified in [87], [82], see also Definition 20.2 in [11]. The tensor powers $V^{n}$ appearing in our setting are canonical choices of truncations of the full tensor powers of the quantum group $U_{q}({\mathfrak{g}})$ obtained by induction over $n$.
The construction of the iterated truncations $V^{n}$ of $U_{q}({\mathfrak{g}})$ at the root of unity is explained in [87] and is the starting point to construct our weak Hopf algebras in Sections 29 and 30. The duality property between the truncated tensor powers $V^{n}$ of the fundamental representation and the representation of the braid group is needed as a further property to complete the proof of FKL theorem on coincidence between associativity morphisms and braiding morphisms on all pairs and triples and is used in our cohomological theorems 8.13 and 8.19. These theorems complete with cohomogical arguments what was proved up to that point in our work, coincidence of these structure maps on pairs and triples where one or two coordinated are fixed to the fundamental representation $V$. To do so, the way how our weak Hopf algebras are described in the application to the quantum group fusion category is also used.
For all the Lie types, the truncated tensor powers $V^{n}$ are inductively defined over $n$ by a canonical decomposition, which is into irreducible components for ${\mathfrak{g}}\neq E_{8}$, of a full tensor product of representations $V_{\lambda}\otimes V$ of $U_{q}({\mathfrak{g}})$, with $V_{\lambda}$ an arbitrary irreducible of the semisimplified fusion category. Then one selects irreducible summands with nonzero quantum dimension. The fusion of the fundamental representation of $E_{8}$ is described in [87] as well, and is more delicate.
Returning to the question of whether the braid group representation generates the centralizer algebras $(V^{n},V^{n})$ for the truncated tensor powers of $U_{q}({\mathfrak{g}})$ at the root of unity, this question can be studied in the same way as for the case of a corresponding summand with $q$ not a root of unity, as in the truncated summand the root of unity property is not visible. This is explained among other things in Prop. 2.4 in [87], see also Prop 29.2 for the same result with a slightly more detailed proof, where an important role is played by the continuous curve described in that proposition, that passes through generic values of $q$. See also Section 3.5 in [87] for the special tensor products $V_{\lambda}\otimes V$ of representations of $U_{q}({\mathfrak{g}})$. Thus it suffices to have the duality property between the full algebras $(V^{\otimes n},V^{\otimes n})$ of $U_{q}({\mathfrak{g}})$ when $q$ is not a root of unity.
We recall that the fundamental representation $V$ are listed in Sect. 3.5 in [87] for all the Lie types, and in particular are given by
    The vector representation for Lie types $A$ and $C$,
    The spinor representation for Lie types $B$ and $D$. In particular $V$ is the direct sum of the two spinor representations for type $D_{n}$ with $n$ even.
    The 7-dimensional representation for Lie types $G_{2}$.
The Lie types $A$ and $C$ for the vector representation are well known. If $q$ is not a root of unity, the centralizer algebras $(V^{\otimes n},V^{\otimes n})$ in the type $A$ case are described by Hecke algebras, a quotient of the complex braid group algebra by well known generators and relations. The first result in this direction is due to Jimbo [54]. For the Lie types $B$, $C$, $D$, when $V$ is the vector representation, Murakami and Birman and Wenzl introduced and studied what are known as the BMW algebras [67], [4]. A nice exposition of the Hecke and BMW algebras may be found in the book by Goodman, de la Harpe and Jones [38]. Duality for the vector representation is of quantum groups was studied by Kirillov and Reshetikhin [58]. A summary of these results and references to the original papers for the vector representation for all the classical Lie types may be found in Theorem 10.2.5 of [9].
The centralizer algebra of the fundamental representation for types $B$ and $D$ is studied in [81], [89].
Type $G_{2}$ was shown by Lehrer and Zhang [60], Morrison [65], and Martirosyan and Wenzl [62].
In the application of Theorems 8.13 and 8.19, we work on the side of the quantum group fusion category and by the results above, our assumptions on braid group duality for the fundamental representation of our weak Hopf algebras $A_{W}({\mathfrak{g}},q)$ constructed in Section 30 are verified for the defining structure arising from Tannakian construction on the quantum group fusion category ${\mathcal{C}}({\mathfrak{g}},q))$ and Wenzl functor.
12. More remarks from the viewpoint of Doplicher-Roberts reconstruction in AQFT on the relevance of duality for the fundamental representation in the construction of quantum gauge group
We also remark that Schur-Weyl duality plays an important role in the abstract duality theory for compact groups by Doplicher and Roberts. It is thanks to the fact that the usual permutation symmetry of the category of Hilbert spaces represents all the the intertwining operators between tensor powers of the defining representation of the unitary groups ${\rm SU}(d)$, that their construction of the compact gauge group associated to the category of localized endomorphisms of the observable algebras in algebraic quantum field theory takes place, as a closed subgroup $G\subseteq{\rm SU}(d)$.
If one regards affine vertex operator algebras at positive integer levels as models corresponding to the case of a gauge group given by the simple compact Lie group ${\rm SU}(d)$ in Doplicher-Roberts theory of quantum gauge fields (i.e.versus the case of gauge group $G$ given by any compact Lie group (a proper closed subgroup of ${\rm SU}(d)$), then from this viewpoint one may expect that duality properties for the fundamental representation $V$ of a quantum group should play a central role in our approach to FKL theorem. The analogue of permutation symmetry in our case is its generalization to the braid group symmetry, and therefore it is natural to consider all the Lie types for which braid group duality is known to hold for the generating representation $V$, as a first step towards a Doplicher-Roberts reconstruction of a field algebra from the local observable algebras in low dimensional quantum field theory.
It may therefore be somewhat surprising that the need to use such an important property as generalizations of Schur-Weyl duality to Lie types beyond $A$ and to the quantum case, has not been noted before as a key property to understand the proof of Theorem 2.1, in the literature but also in the previous versions of [11], given that our approach is originally inspired by Doplicher-Roberts duality theory for compact groups with similar applicative intents to the construction of a quantum gauge group that we construct from the emerging categories in low dimensional quantum field theory as an analogue of a gauge group, aimed to the construction of the quantum field algebra from the local observable algebras.
Perhaps similarly surprising is the fact that in our approach we do not invoke modularity. On the other hand by a private discussion with M. Yamashita, we learned that modularity is used in their recent work on classification of fusion categories with fusion rules of type BCD [39].
Thus the fact that braid group duality for the fundamental representation $V$ was the relevant missing observation helpful to complete the proof of Theorem 2.1 in the previous version of [11], which led us to Theorems 8.13 and 8.19 as a completion of Lemma 7.10 in version 7 of [11], was for us a further important recognition of the relevance of the role of quantum groups. We refer the reader to the beautiful historical paper by V. F. R. Jones [55] for a general first introduction to the origin of quantum groups as solutions to the Yang-Baxter equation.
We find it remarkable the deep application that braid group duality has to quantum field theory. This application led us to evaluate even more the relevance of the quantum group viewpoint than as noted up to that point in our study.
We also would like to point out that our results of the categorical characterization of the class of unitary coboundary weak Hopf algebras that we obtain in Sect. 26, 27, 29 in [11] extends the notion of symmetric functors that plays a key role in Doplicher-Roberts theory [21], [22].
13. Outlook
In Doplicher-Roberts theory, a unique gauge compact subgroup of ${\rm SU}(d)$ is in particular constructed in the case of a rigid symmetric tensor $C^{*}$-category with simple unit and a single generating object, to describe the category when the generating property of the permutation group is not satisfied. Their theory has applications to the category of localized endomorphisms arising from the local observable algebras in high dimensional AQFT, and leads to the construction of the field algebra [21], [22].
For the class of the compact quantum groups of Woronowicz there is a construction of their quantum homogeneous spaces and Hilbert bimodule representations developed for non-symmetric rigid tensor categories with simple unit [75], [79]. The construction of the homogeneous space and bimodule representations of the compact quantum group extends to compact quantum groups the idea of closed subgroup construction in Doplicher-Roberts theory. In that quantum case the homogeneous space is an ergodic action of a compact quantum group of a unital $C^{*}$-algebra, and may not arise from a quantum subgroup.
In the setting of compact quantum groups, unitary representations of the braid group are not known to arise in a natural way. On the other hand, they arise from fusion categories at roots of unity in [87], and this paper was our first motivation to study these quantum groups and construct weak Hopf $C^{*}$-algebras in the sense of [11], with precedents the examples by Mack and Schomerus for ${\mathfrak{s}l}_{2}$, described as weak quasi-Hopf algebras in those early papers where the study of the unitary structure was not addressed yet [61], and its extension to ${\mathfrak{s}l}_{N}$ studied in [10] where the study of the unitary structure was included following the results of [87].
Our weak Hopf $C^{*}$-algebras are suitable extensions of the compact quantum groups, and in our experience the failure of strict co-associativity has played a mild role, and has allowed the development of a theory extending that of Hopf algebras, or of compact quantum groups if one allows twisted unitary structure, called $\Omega$-involutions in [11] for which classes of examples arising from ribbon tensor $C^{*}$-categories allow unitary representations of the braid group in a natural way, called unitary coboundary weak Hopf $C^{*}$-algebras in the same paper.
It is then desirable to search for further generalizations of quantum homogeneous spaces and bimodule representations to these weak Hopf algebras to describe ribbon braided tensor $C^{*}$-categories with unitary representations of the braid group, such as the representation categories of models of CFT: conformal nets where the unitary structure is given from the start, or related vertex operator algebras where is constructed. In the affine case that we have studied in our papers, the unitary, rigid, and modular structure in the latter can be transported from that of the quantum group fusion category via the equivalence. This may be useful for those models for which, in analogy to the theory of Doplicher-Roberts for rigid symmetric tensor categories, the generating property of the braid group is not satisfied for the algebras of intertwining morphisms of the category. Or also for models for which weak Hopf algebras can not arise from failure of submultiplicativity of the integer valued dimension function of representations of the Zhu algebra. Natural classes of models to study in this direction are referred to as coset models.
Such an extension may prove useful to construct a field algebra from a braided tensor $C^{*}$-category, for which conformal nets provides examples, from the algebra of local observables, taking into consideration the quantum group fusion categories at roots of unity and the associated weak Hopf $C^{*}$-algebras constructed in [11], as developments of early works in that direction of the 90s.
Other interesting future developments may regard applications in noncommutative geometry in the sense of Connes [13] of our weak Hopf algebras $A_{W}({\mathfrak{g}},q)$ associated to the quantum group fusion category ${\mathcal{C}}({\mathfrak{g}},q)$ or their twisted weak quasi-Hopf algebra structures on the Zhu algebra of an affine vertex operator algebra at a positive integer level $V_{{\mathfrak{g}}_{k}}$ to the study of Dirac operators, by expanding an analog in conformal field theory of the groundbreaking work by Neshveyev and Tuset on Dirac operators for compact quantum groups in the setting of Kazhdan-Lusztig theorem for real values of the quantization parameter $q$ [69].
 [^1] The claims, arguments, and reasoning of the recent papers [6], [7], [8] are essentially dependent on the results and in-depth discussions between S.C. and C.P. that took place from 2018 to 2022 on the contents of an entire project including [11] and the topics of the above papers, during a period in which S.C. was a co-author, before breaking the collaboration in early 2023, with a minor contribution in a draft of a single section. In addition, C.P. department digital devices were subject to a serious hacking attack for several months in 2021 during the previous collaboration, which was recognized by authorities, that severely damaged her intellectual property and health.
Acknowledgements The author is grateful to Yi-Zhi Huang for several comments on the proof of modularity for fusion categories of affine vertex operator algebras, and the history of Finkelberg-Kazhdan-Lusztig theorem. She is also grateful to Marco V. Giannone for pointing out to her incompleteness of the proof of Lemma 7.10 in version 7 of [11] and to Hans Wenzl for kind email correspondence and for references on duality for tensor powers of the fundamental representation of Drinfeld-Jimbo quantum groups for the Lie types $B$, $D$, $G_{2}$, leading to the present paper. Finally she thanks Pavel Etingof for informing her of [28].

### v2

@@
 Turaev gave a definition of modular tensor category, based on previous work with Reshetikhin on the construction of knot and 3-manifold invariants see [89].
-Several authors contributed to construct modular fusion categories form quantum groups at roots of unity, and a more complete historical account on the quantum group side may be found in [80], [82].
-see also the introduction of [11].
+Several authors contributed to construct modular fusion categories form quantum groups at roots of unity, and a more complete historical account on the quantum group side may be found in [83], [85], see also the introduction of [11].
 We denote by ${\mathcal{C}}({\mathfrak{g}},q)$ the resulting modular fusion category, with $q=e^{i\pi/d\ell}$, and $d\in\{1,2,3\}$ a suitable integer structure constant from Lie theory.
@@
 Modular, or more generally ribbon rigid braided tensor category constructions involved many authors among physicists and mathematicians.
-Major work was carried out in a series of papers by Kazhdan and Lusztig in the mid 90s.
+A main difficulty in constructing tensor categories from models of conformal field theories is the construction of the associativity morphisms compatible with naturally assigned braiding morphisms via the two hexagonal equations.
 Let ${\mathfrak{g}}$ be a complex simple Lie algebra.
+This problem was first solved by Drinfeld in [23], [25] [26] for the Drinfeld category with the introduction of the notion and construction of non-trivial examples of quasi-Hopf algebras over $U({\mathfrak{g}})[[h]]$, via the Knizhnik-Zamolodchikov differential equation which is satisfied by the correlation functions of the WZW model.
+Drinfeld used this equation to construct the associativity morphisms compatible with naturally assigned braiding morphisms.
+Furthermore Drinfeld established his remarkable Drinfeld-Kohno theorem expanding an earlier theorem by Kohno [62] stating that the representation of the braid group arising from monodromy of the KZ equation is equivalent to that arising from the action of the $R$-matrix of $U_{h}({\mathfrak{g}})$ with respect to the vector representation, for the classical Lie types.
+Drinfeld-Kohno theorem connects via a Drinfeld twist, the quasi-Hopf algebra $U({\mathfrak{g}})[[h]]$ with the quantized universal enveloping algebra $U_{h}({\mathfrak{g}})$ which is a Hopf algebra.
+In this case, rigidity of Drinfeld category is derived from the equivalence constructed in Drinfeld-Kohno theorem and rigidity of the representation category of $U_{h}({\mathfrak{g}})$.
+Major work was carried out in a series of papers by Kazhdan and Lusztig in the mid 90s in the setting of affine Lie algebras.
 The authors construct rigid braided tensor structures for certain module categories of affine Lie algebras associated to ${\mathfrak{g}}$ at levels excluding the positive rationals.
@@
 As I learned from discussions with Huang, rigidity was not proved in [2], and there was no proof of rigidity using this approach.
-Remarkably, Huang and Lepowsky introduced and developed a rigorous direct construction of a modular (vertex) tensor category structure in a very general setting for vertex operator algebras, including as a particular case the affine vertex operator algebras at positive integer levels that we denote by $V_{{\mathfrak{g}}_{k}}$.
+Remarkably, Huang and Lepowsky introduced and developed a rigorous direct construction of a modular (vertex) tensor category structure in a very general setting for vertex operator algebras, including as a particular case the affine vertex operator algebras at positive integer levels $k$ that we denote by $V_{{\mathfrak{g}}_{k}}$.
 The construction has been completed by Huang between 2005 and 2008 [49], [53] following a general construction of tensor product theory in the setting of vertex operator algebras by Huang and Lepowsky [44], [45], [46] and by Huang [47], [52].
@@
 These works give in particular the first mathematical proof of modularity, that applies moreover to a large class of representation categories of vertex operator algebras.
-We note that difficulties in the construction of associativity morphisms in the setting of Drinfeld category were earlier solved by Drinfeld in [23], [24], [25] by the introduction and construction of quasi-Hopf algebras, via differential Knizhnik-Zamolodchikov equation, leading to his remarkable Drinfeld-Kohno theorem connecting via a Drinfeld twist with the quantized universal enveloping algebra.
-In this case, rigidity of Drinfeld category is derived from the equivalence constructed by Drinfeld in Drinfeld-Kohno theorem.
 In particular rigidity and modularity property in the setting of vertex operator algebras was first shown in [49].
@@
 We also mention Gannon’s observation in Sect. 6.2.3 of [36] regarding the desire of finding a direct connection between quantum groups and CFT.
+From the side of mathematical axiomatizations of CFT, we work with respect to Huang and Lepowsky braided tensor structure for the representation category of vertex operator algebras.
 A direct construction of the equivalence may be helpful to see how rigidity or modularity in the setting of CFT may be transported from corresponding results in the setting of quantum groups at roots of unity, which seem simpler, or vice versa, by a given explicit equivalence.
 Or to simplify or give new construction of other structures, e.g.
-unitary structures, with a quantum group viewpoint, or to see more connections with quantum groups from these fundamental models.
-The study of how unitarization from the quantum group fusion categories [87] was useful to derive unitarization in representation categories of vertex operator algebras via an explicit equivalence was important for us since the early stages of [11], and further motivated us to study the above problem.
+unitary structures, with a quantum group viewpoint, or to see more connections with quantum groups of models derived from or generalizing these fundamental models.
+The study of how unitarization from the quantum group fusion categories [91] was useful to derive unitarization in representation categories of vertex operator algebras via an explicit equivalence that passes through weak quasi-Hopf algebras was important for us since the early stages of applications of the theory of weak quasi-Hopf algebras [11], and further motivated us to study the above problem.
+Indeed one of our first results on unitarization was that weak quasi-Hopf algebras associated to fusion categories known to be equivalent allow to transport the unitary structure of one category to unitarize the other.
+In the process, knowledge of the explicit equivalence gives an understanding of the transported unitary structure.
+We have results on the structure of the unitary structure of the affine vertex operator algebra at positive integer levels transported by our equivalence, which asserts that it must correspond to the classical compact real form, and necessarily equals the unitary structure previously directly constructed by Gui, see [11] for more details and references.
 Our approach to Huang problem is based on interpreting Finkelberg-Kazhdan-Lusztig equivalence as an extension to quantum groups of Doplicher-Roberts duality theory for the important class of WZW models that are described by affine Lie algebras at positive integer levels.
 We shall briefly come back to the previous history of these important models in physics.
-From the side of mathematical axiomatizations of CFT, we work with respect to Huang and Lepowsky braided tensor structure for the representation category of vertex operator algebras.
-Our starting point to approach this problem arose from a problem generously posed by Sergio Doplicher in the 90s to the author in a different setting, that of algebraic quantum field theory.
+Our starting point to approach this problem arose from a related problem generously posed by Sergio Doplicher in the 90s to the author in a different setting, that of algebraic quantum field theory.
 We try to explain this problem.
@@
 In this context, I learned from private discussions that for high dimensional quantum field theories categorical structures in their study first emerged from the discovery by Doplicher of the notion of localized endomorphism of the local observable algebra, allowing composition of two of them which was understood as a strict tensor product structure.
-The mentioned works by Doplicher, Haag, and Roberts led to a deep construction of the structure of a rigid symmetric tensor $C^{*}$-category with simple unit, starting from first principles in physics.
-Rigidity, called charge conjugation in this setting, is due to particle-antiparticle duality.
-In their papers rigidity was shown to imply deep and useful properties for the classification of the statistics parameter, and charges of para-Bose or para-Fermi type.
-Because of the symmetry, their results on the classification statistics parameter, which was shown to the the reciprocal of a positive integer $d$ up to a sign, and $C^{*}$-structure, in the 70s it arose the question of whether that was the category of representations of a compact group, because those data arising from a single localised endomorphism were also satisfied by the representation category of a compact Lie subgroup of ${\rm SU}(d)$, that would be a gauge group of global symmetries.
+The mentioned works by Doplicher, Haag, and Roberts led to a deep construction of the structure of a rigid symmetric strict tensor $C^{*}$-category with simple unit, starting from first principles in physics.
+Rigidity, called charge conjugation in this setting, describes particle-antiparticle duality.
+In their papers rigidity is derived as a weak limit of inner automorphisms induced by transfer operators moving the support of the given localized endomorphism $\rho$ to infinity.
+This gives a left inverse $\Phi_{\rho}$ of $\rho$, which is a completely positive map.
+From this the conjugate localized endomorphism $\overline{\rho}$ is derived.
+The left inverse may be described in the category as a map induced by the adjoint action of a solution of the conjugate equations on a intertwiner translated to the right by the tensor product.
+It was shown that charge conjugation implies deep and useful properties for the classification of the statistics parameter (which is the scalar $\Phi_{\rho}(\varepsilon)$, with $\varepsilon$ the action of the generator of the permutation group in the algebra of intertwining operators of the tensor powers of $\rho$), and of charges into charges of para-Bose or para-Fermi type.
+Because of the representation of the permutation group in the category, together with their results on the classification of the statistics parameter, which was shown to the the reciprocal of a positive integer $d$ (the order of the parastatistics) up to a sign (Bose or Fermi cases), and the $C^{*}$-structure, in the 70s it arose the question of whether that was the category of representations of a compact group, because those data arising from a single localised endomorphism were also satisfied by the representation category of a compact Lie subgroup of ${\rm SU}(d)$, that would be a gauge group of global symmetries.
 They considered the abelian case in the 70s [17].
@@
 In a series of important papers dated the end of the 80s, Doplicher and Roberts were able to solve the problem in full generality.
-Their construction gives a unique compact gauge group with representation category equivalent to that of a given symmetric rigid tensor $C^{*}$-category with simple unit, yielding the formulation of a new abstract duality theory for compact groups [19], [21],[20], and led to desired applications to high dimensional Algebraic QFT [22].
+Their construction gives a unique compact gauge group with representation category equivalent to that of a given symmetric rigid strict tensor $C^{*}$-category with simple unit, yielding the formulation of a new abstract duality theory for compact groups [19], [21],[20], and led to desired applications to high dimensional Algebraic QFT [22].
 The term abstract duality refers to the property that the given categories do not have a tensor functor to the category of f.d.
@@
 We refer to the introduction in [9] and more historical information at the end of each chapter in their book.
+Woronowicz developed a theory of compact quantum groups as noncommutative function algebras in the setting of $C^{*}$-algebras and constructed the first important examples as deformation of ${\rm SU}(d)$ by a real parameter, see [73] and reference to the original papers, via a concrete Tannaka-Krein duality for $C^{*}$-algebras extending the classical case.
 After the discovery of the quantum groups in the mid eighties, Doplicher asked the author to try to extend their theory and consequent applications on the construction of a quantum gauge group and a related field algebra.
@@
 Our approach may be regarded in the setting of Connes noncommutative geometry [13], and follows naturally our previous work in the setting of compact quantum groups, see also the introduction in [12], [14].
-In [73], [74], [78], [75], [76], [77], [79] we have developed at the level of tensor categories an analogue of Doplicher-Roberts construction of compact gauge group, although the analogy cannot be performed in a strict way due to new difficulties in the quantum case.
-We shall come back to this, see the mentioned papers and [11].
+In [76], [77], [81], [78], [79], [80], [82] we have developed at the level of rigid tensor $C^{*}$-categories with simple unit an analogue of Doplicher-Roberts construction of a compact gauge group and field algebra, or more precisely the analogue of [21] in the setting of compact quantum groups.
+The analogy cannot be performed in a strict way due to new difficulties in the quantum case, see the mentioned papers and [11].
 One important difference was that for rigid braided tensor categories we do not have a notion of braided tensor functor beyond the symmetric ones, because the category of vector spaces does not admit a braided symmetry.
@@
 This was noticed early as an easy consequence from their construction.
-To treat this difficulty in [10] and [11] we have taken into account Drinfeld notion of coboundary, which finally allows for a generalization.
+To describe our approach, we recall that Drinfeld original motivation for introducing quasi-Hopf algebras was to interpret them as symmetries to describe braided tensor categories from important models of rational conformal field theories motivated by the work by Moore and Seiberg.
+Although as noted in [24], they could not describe precisely simple models such as the Ising, Drinfeld evaluated quasi-Hopf algebras a good description of quantum symmetries for the WZW model in an approximate way, i.e.
+in the limit to infinity of the central charge.
+From this perspective, we regard the equivalence described by Drinfeld-Kohno theorem as a Tannakian counterpart at the level of quasi-Hopf algebras of Kazhdan-Lusztig-Finkelberg equivalence $\tilde{\mathcal{O}}_{k}\to{\mathcal{C}}({\mathfrak{g}},q)$.
+This point of view is the author understanding of the simplified proof given by Neshveyev and Tuset of Drinfeld-Kohno theorem following the proof by Kazhdan and Lusztig of the same theorem over the complex numbers, and for real values of the deformation parameter $q$, in the setting of compact quantum groups.
+Their approach inspired us to a similar strategy in our case of an abstract duality theory for both the fusion categories of quantum groups at roots of unity and of affine vertex operator algebras at positive integer levels [71].
+There have been studies that construct analogues of Doplicher-Roberts field algebra using generalizations of quasi-Hopf algebras (the weak quasi-Hopf algebras) as quantum gauge groups, and they showed complications reflecting non-triviality of the associator, see [86].
+It would be desirable to be able to use quantum symmetry groups that may be regarded in the family of Hopf algebras.
+To treat the difficulty of finding a replacement of Doplicher-Roberts symmetric functors, in [10] and [11] we have taken into account Drinfeld notion of coboundary symmetry $\overline{R}$, which plays a role in the construction of Drinfeld twist in Drinfeld-Kohno theorem [25].
+This is a symmetry of the category obtained modifying the braided symmetry defined by the $R$-matrix by a square root of the $2$-coboundary defined by the ribbon structure, and by construction the coboundary symmetry is an involutive symmetry, similarly to the braided symmetry.
+This symmetry finally allows for a generalization of the notion of symmetric functor.
 Moreover we have followed both the original choice by Mack and Schomerus [64] of a given dimension function corresponding to non-negligible tensor powers of the vector representation of quantum ${\mathfrak{s}l}_{2}$ at roots of unity and also the study of classification of compact quantum groups corresponding to a given representation category of deformation compact quantum groups by Neshveyev and Yamashita [74], [75].
 They both considered a specific positive integer valued dimension function associated to the given category of the quantum group, which was the classical dimension function.
-This suggested to us to consider in [10] and [11] a specific dimension function on a functor related to classical data, and this was first Wenzl functor [87] for the quantum group fusion category, and then the minimum energy functor for the WZW model [93].
-In was particularly important for us that in [87] Wenzl describes an explicit quantization curve associated to his functor that leads to an algebra that identifies with the Zhu algebra for that model in the setting of vertex operator algebras.
-Since in [10] we had constructed a natural quantum group structure on the former compatible with the tensor structure, this arose the question of constructing another natural quantum group structure on the latter at the time of [10], following the quantization curve on the unitary structure of tensor products with the aim of making it classical.
-This involved in [11] the construction of a Drinfeld twist, a notion that Drinfeld had introduced to study equivalence of what is known as the Drinfeld category with the quantum group category in a formal setting.
+This suggested to us to consider in [10] and [11] a specific dimension function on a functor related to classical data, and this was first Wenzl functor [91] for the quantum group fusion category now at roots of unity, and then the minimum energy functor for the WZW model [98] in the setting of vertex operator algebras.
+It was particularly important for us that in [91] Wenzl describes an explicit quantization curve associated to his functor that leads to an algebra that identifies with the Zhu algebra for that model in the setting of vertex operator algebras.
+Since in [10] for the type $A$ case we had constructed a natural quantum group structure on a semisimple associative algebra from the quantum group fusion category endowed with Wenzl functor, compatible with the tensor structure, this arose the question of constructing another natural quantum group structure on the Zhu algebra at the time of [10], following the quantization curve on the unitary structure of tensor products with the aim of making it classical.
+This involved in [11] the construction of a Drinfeld twist in our setting, the above mentioned notion that Drinfeld had introduced to study equivalence of the Drinfeld category with the quantum group category in a formal setting.
+We have done this for all the Lie types via our version of Drinfeld-Kohno Theorem 28.4 and applications to affine vertex operator algebras at positive integer level, via Zhu linear equivalence, in [11].
 We refer to Sect.
 23 in [11] for an extension of certain results by Neshveyev and Yamashita in the setting of quantum group fusion categories of type $A$ based on Kazhdan-Wenzl theory at roots of unity in our case, and on the weak Hopf algebras constructed in [10], in place of the compact quantum groups.
-This section is important for us as it suggested that in general the the associativity morphisms in a category arising from CFT may be of a simple nature from the viewpoint of associated quantum gauge groups that need to be constructed, and this finally proved to be the case, we shall come back to this with more precision.
-On the other hand, in the setting of compact quantum groups, while extending Doplicher-Roberts theory starting from embedding part of the morphisms of the category into the Hilbert spaces [73], [75], [79], we had already started to fix a dimension function, and from that we constructed a tensor functor from that subcategory.
-To consider the full category, the analogue Doplicher-Roberts compact gauge group is described by an ergodic action of a larger quantum group on a quantum homogeneous space, rather than a quantum group itself, due to lack of commutativity of the $C^{*}$-algebra describing the homogeneous space constructed in [75], and this property is likely to hold also in other models arising from conformal field theory, unless one can show that certain known obstructions already met in [74] for compact quantum groups, vanish.
-We shall come back to this.
-However, for the WZW models arising from affine Lie algebras at positive integer levels of the Lie types mentioned in the abstract, we are able to show in [11] that we obtain genuine quantum groups as quantum gauge groups, described as semisimple weak Hopf algebras with $R$-matrix, $C^{*}$-structure explicitly associated to Drinfeld coboundary matrix, and with unitary representations of the braid group, derived from the quantum group fusion category.
-This weak Hopf algebra has representation category equivalent to both the starting fusion category and also to that of the corresponding affine vertex operator algebra.
-Constructing the equivalence from the former to the latter as we do in [11] is our proof of Kazhdan-Lusztig-Finkelberg theorem, see Theorem 2.1.
+This section is important for us as it suggested that in general the associativity morphisms in a category arising from CFT may be of a simple nature from the viewpoint of the associated quantum gauge groups that need to be constructed, and this finally proved to be the case, we shall come back to this with more precision.
+On the other hand, in the process of extending Doplicher-Roberts theory in the setting of compact quantum groups, we had started by embedding in a tensorial way part of the morphisms of the category into the Hilbert spaces [76], [78], [82], and in doing this we had already made a choice of a dimension function, and from that choice we had constructed a tensor functor from that subcategory.
+To consider the full category, the analogue Doplicher-Roberts compact gauge group is described by an ergodic action of a larger quantum group on a quantum homogeneous space described by a $C^{*}$-algebra, rather than a quantum group itself, due to lack of commutativity of this $C^{*}$-algebra [78].
+This property is likely to hold also in other models arising from conformal field theory starting from the minimum energy functor, unless one can show that certain known obstructions already met in [77] for compact quantum groups, vanish, we shall come back to this in the outlook section 13.
+However, for the WZW models arising from affine Lie algebras at positive integer levels in [10] for the type $A$ case, and in [11] for all the Lie types, we obtain genuine quantum groups as quantum gauge groups, described as semisimple weak Hopf algebras with $R$-matrix, $C^{*}$-structure explicitly associated to Drinfeld coboundary matrix, and with unitary representations of the braid group, derived from the quantum group fusion category.
+We prove that these weak Hopf algebras have representation category equivalent to both the starting fusion category and also to that of the corresponding affine vertex operator algebra for the classical Lie types and $G_{2}$.
+Constructing the equivalence from the former to the latter in the setting of semisimple weak quasi-Hopf algebras as we do in [11] is our proof of Kazhdan-Lusztig-Finkelberg theorem, see Theorem 2.1.
 Among our first references in the context of conformal field theory important for us to start to set up the connection with the work of Wenzl in quantum groups, is the work by Wassermann [90] in the setting of affine Lie algebras, where the key concept is played by the notion of primary field, and its initial term, which we use in a fundamental way to connect quantum group fusion category to representations of the Zhu algebra and their special tensor products with the fundamental representation, see Sect 33 in [11].
@@
 27.
-This property is expressed in terms of quantum groups, and the first example, albeit in a non-semisimple setting, is the hermitian structure of $U_{q}({\mathfrak{g}})$ at a root of unity by Theorem 29.9.
+This property is expressed in terms of quantum groups, and the first example, in a non-semisimple setting, is the hermitian structure of $U_{q}({\mathfrak{g}})$ at a root of unity by Theorem 29.9.
 This property says that the -involution enjoys a precise anticommutation relation (rather than commutation relation as is familiar in Hopf $C^{*}$-algebras) with the coproduct and was first noticed in [91].
@@
 Further examples derived from this examples are our weak Hopf and quasi-Hopf algebras, see Sections 29, 30, 31.
-We shall derive even more structure than rigidity e.g.
-unitarity and modularity on the vertex operator algebra representation category from the corresponding quantum group fusion category, via the equivalence, see Sects.
+This is not a direct consequence of the corresponding property of $U_{q}({\mathfrak{g}})$, as its coproduct is not compatible with the fusion tensor product.
+As above motivated, we shall derive more structure than rigidity and modularity, such as unitarization of the vertex operator algebra representation category from the corresponding quantum group fusion category, via the explicit equivalence, see Sects.
 21, 22, 24 in [11].
-Moreover our equivalence, which is given by an analogue of the Drinfeld twist in [23], [24], gives an understanding of the unitary structures of representation categories of affine vertex operator algebras as classical compact real forms, or untwisted.
+To this aim it is important that our equivalence passes through the weak quasi-Hopf algebras, and the fact that these are naturally associated to the categories.
+Indeed it was known that a mere tensor equivalence between the categories would not suffice to transport the unitary structures see [43].
+Sect.
+15 in [11] treats this for general tensor categories that admit submultiplicative dimension functions (called weak dimension functions).
+Moreover our equivalence, which is given by an analogue of the Drinfeld twist in [23], [25], at the level of the associated weak Hopf or quasi-Hopf algebras gives an understanding of the unitary structures of representation categories of affine vertex operator algebras as classical compact real forms, or untwisted.
 We refer to the outlook section for some more information on possible extension of our viewpoint to models beyond the affine vertex operator algebras at positive integer levels, see also [11].
@@
 In the subsequent sections we outline the plan of our proof.
-Our proof starts from the work of Wenzl [87] in the setting of quantum group fusion categories.
 A main idea that we follow is to consider the fundamental representation $V$ of ${\mathfrak{g}}$ for each Lie type as in [91] where the author explains its generating property and good fusion structure with respect to tensor products with an arbitrary irreducible representation allowing a well-defined truncation procedure to obtain non-negligible objects, with the aim of studying unitarizability of the fusion category.
 It should also be noted that a main application of that paper is construction of inclusion of subfactors of finite Jones index for all Lie types, extending his previous work.
-For all the Lie types we obtain complete identification for the tensor product structure, all the ribbon structure, and those braiding and associativity morphisms where one or two variables are the fundamental representation $V$, respectively.
+For all the Lie types we obtain complete identification for the tensor product structure, all the ribbon structure, and those braiding and associativity morphisms where one or two variables are the fundamental representation $V$, respectively, with Huang-Lepowsky ribbon braided tensor structure.
 These are parts (a), (b) of Theorem 2.1.
-The main method of our proof is the construction of semisimple weak Hopf or quasi-Hopf algebras in the two settings and an equivalence of their representation categories via the construction of a Drinfeld twist.
 The term weak in the setting of quasi-Hopf algebras was first introduced by Mack and Schomerus motivated by the work of Drinfeld for quasi-Hopf algebras and Drinfeld-Kohno theorem [64].
@@
 The definition of weak Hopf algebra has become standard and understood in the sense by Böhm and Szlachanyi.
-In [11], and also in this paper, weak Hopf algebras are understood with a different meaning.
+In [10], [11], and also in this paper, weak Hopf algebras are understood with a different meaning.
 While a different name would be helpful to avoid confusion, so far we have not been able to find a good name for our algebras to clarify the situation, and we use the term weak Hopf algebras in the sense defined in Sect 9 in [11].
@@
 We follow the strategy of Drinfeld for Drinfeld-Kohno theorem which includes the comparison of braiding morphisms of the two settings, but we do not include the comparison of the associativity morphisms in our analogue, Theorem 28.4 in [11] (more details are given later).
-We construct a Drinfeld twist for our weak Hopf algebras useful for the braided symmetry, and study associativity morphisms with a uniqueness theorem using the generating representation.
-We then use this twist to concretely construct unitary structures in the setting of the representation categories of affine vertex operator algebras at positive integer levels with the tensor structure due to Huang and Lepowsky, for all the mentioned Lie types, by transferring the unitary structure from quantum group fusion category constructed in [87].
+We construct a Drinfeld twist for our weak Hopf algebras useful for the braided symmetry in a way similar to Drinfeld method, by taking a square root of the coboundary matrix, in our case this is slightly more delicate because of the truncation procedure and analytic aspects.
+Similarly to Drinfeld-Kohno theorem, we establish equality of the associativity morphisms arising from quantum group fusion category after the twist applied to our weak Hopf algebras with Huang-Lepowsky associativity morphisms via a uniqueness theorem of a cohomological nature, but in our case this theorem seems different, in that we use cohomology of the pentagon equation on the generating representation $V$ and extensions to quantum case and Lie types of Schur-Weyl duality.
+We shall come back to this in the last part of the introduction and in the following sections of this paper.
+We apply our concrete twist to transport unitary structures to representation categories of affine vertex operator algebras at positive integer levels with the tensor structure due to Huang and Lepowsky, for all the mentioned Lie types, by transferring the unitary structure from quantum group fusion category as constructed in [91].
 Our weak Hopf algebras while having a non-unital coproduct, are not in the class of Bohm and Szlachanyi.
@@
 They are not extension of Bohm and Szlachanyi weak Hopf algebras, and we know from results in [11] that the intersection of the two classes are the genuine Hopf algebras.
-On the other hand, in our opinion the weak Hopf algebras considered in [11] still deserve to be regarded as Hopf algebras in a weak sense, for various reasons.
+On the other hand, in our opinion the weak Hopf algebras considered in [10] and [11] still deserve to be regarded as Hopf algebras in a weak sense, for various reasons.
 On one hand they seem to play the role of the quantized universal enveloping algebras in Drinfeld-Kohno theorem, following the original motivation of Mack and Schomerus.
@@
 $8$ of [11] that has not emerged in other parts.
-For the type $A$ case, the extra structure that we need to use is known as quantum Schur-Weyl duality between the vector representation and the natural representation of the braid group between the tensor powers of the vector representations.
+As above mentioned, in the type $A$ case, the extra structure that we need to use is known as quantum Schur-Weyl duality between the vector representation and the natural representation of the braid group between the tensor powers of the vector representations.
 In the classical case, Schur-Weyl duality is a key property for the general linear groups starting with the work of Schur and Weyl.
-Schur-Weyl duality also takes an important role in Doplicher-Roberts abstract duality theory for compact groups and applications to construction of the field algebra.
+We note that Schur-Weyl duality also takes an important role in Doplicher-Roberts abstract duality theory for compact groups and applications to construction of the field algebra.
 This classical duality emergences as the starting point for the construction of global gauge symmetry groups in tensor categories arising in [22].
@@
 This allows us to apply our abstract uniqueness Theorem 8.19 in [11] for braided tensor structures in a pre-tensor category.
-In our work, the need of quantum duality between the braid group representation and the fundamental representation $V$ emerged on the occasion of the proof of our earlier Lemma 7.10, which was not complete in the version 7 of [11], and we worked on completion with the same methods.
+In our work, the need of quantum duality between the braid group representation and the fundamental representation $V$ emerged on the occasion of the proof of our earlier Lemma 7.10, which was not complete in version 7 of [11], and we worked on completion with the same methods.
 It was for us a sudden realization of an even greater recognition than we had realized up to that point of the contribution of the authors on extensions of Schur-Weyl duality which has historical interest on his own in representation theory, and more work by Wenzl and his collaborators, and several other authors in the classical and quantum cases for the various Lie types.
@@
 2. The main result
+Let ${\mathfrak{g}}$ be a complex simple Lie algebra and let ${\rm Rep}(V_{{\mathfrak{g}}_{k}})$ be the module category of the corresponding affine vertex operator algebra $V_{{\mathfrak{g}}_{k}}$ at level $k\in{\mathbb{N}}$, endowed with the ribbon braided tensor structure introduced by Huang and Lepowsky.
+Let $A(V_{{\mathfrak{g}}_{k}})$ be the Zhu algebra.
+Let $V$ denote the fundamental representation of ${\mathfrak{g}}$ for each Lie type, defined in [91], (see also Sect.
+11 for the classical types and $G_{2}$).
+We also denote by $V$ the corresponding quantized representation of $U_{q}({\mathfrak{g}})$, $q=e^{i\pi/\ell d}$, the associated object of the quantum group fusion category ${\mathcal{C}}({\mathfrak{g}},q)$, with $\ell=k+\check{h}$, $\check{h}$ the dual Coxeter number of ${\mathfrak{g}}$.
+We use the same notation for the corresponding representation of the vertex operator algebra $V_{{\mathfrak{g}}_{k}}$ and of the Zhu algebra $A(V_{{\mathfrak{g}}_{k}})$.
+Similarly for the irreducible objects $V_{\lambda}$ of the involved categories.
+Let $A_{W}({\mathfrak{g}},q)$ denote the unitary compatible coboundary weak Hopf algebra associated to ${\mathcal{C}}({\mathfrak{g}},q)$ via the construction of a weak tensor structure on Wenzl functor $W:{\mathcal{C}}({\mathfrak{g}},q)\to{\rm Vec}$ following Theorems 30.17, in [11].
+We shall give a brief exposition of $A_{W}({\mathfrak{g}},q)$ in Sect.
+4.
 Theorem 2.1.
-Let ${\mathfrak{g}}$ be a complex simple Lie algebra and let ${\rm Rep}(V_{{\mathfrak{g}}_{k}})$ be the module category of the corresponding affine vertex operator algebra $V_{{\mathfrak{g}}_{k}}$ at level $k\in{\mathbb{N}}$, endowed with the ribbon braided tensor structure introduced by Huang and Lepowsky and let $A(V_{{\mathfrak{g}}_{k}})$ be the Zhu algebra.
-Then
+  • We have that:
   • (a)
-    The representation category ${\rm Rep}(A(V_{{\mathfrak{g}}_{k}}))$ of the Zhu algebra $A(V_{{\mathfrak{g}}_{k}})$ admits a natural structure of ribbon braided tensor category obtained by transporting Huang-Lepowsky structure to ${\rm Rep}(A(V_{{\mathfrak{g}}_{k}}))$.
+    The representation category ${\rm Rep}(A(V_{{\mathfrak{g}}_{k}}))$ admits a natural structure of ribbon braided tensor category obtained by transporting Huang-Lepowsky structure to ${\rm Rep}(A(V_{{\mathfrak{g}}_{k}}))$.
 In this way, Zhu’s linear equivalence
@@
   • (b)
-    Let $A(V_{{\mathfrak{g}}_{k}})$ be endowed with the unitary compatible coboundary weak quasi-Hopf structure induced by ${\mathcal{C}}({\mathfrak{g}},q)$ via the construction of a weak tensor structure on Wenzl functor and of the Drinfeld twist following Theorem 31.1 in [11] (a), and let ${\rm Rep}(A(V_{{\mathfrak{g}}_{k}}))$ be endowed with the corresponding ribbon braided tensor structure.
-    Then ${\rm Rep}(A(V_{{\mathfrak{g}}_{k}}))$ becomes a unitary coboundary modular tensor category in this way.
+    Let $A(V_{{\mathfrak{g}}_{k}})$ be endowed with the twisted unitary compatible coboundary weak quasi-Hopf structure induced by the weak Hopf algebra $A_{W}({\mathfrak{g}},q)$ via Wenzl path and the construction of the Drinfeld twist, following Theorem 31.1 in [11] (a), and let ${\rm Rep}(A(V_{{\mathfrak{g}}_{k}}))$ be endowed with the corresponding ribbon braided tensor structure.
+    Then ${\rm Rep}(A(V_{{\mathfrak{g}}_{k}}))$ becomes a unitary coboundary modular tensor category in this way with the structure derived from ${\mathcal{C}}({\mathfrak{g}},q)$ and the equivalence of Theorem 31.1 in [11],
+    ${\rm Rep}(A(V_{{\mathfrak{g}}_{k}}))\xrightarrow{}{\rm Rep}(A_{W}({\mathfrak{g}},q)).$
     Moreover the transport in (a) preserves the tensor products, the ribbon for all objects, the braid morphisms for pairs $(V_{\lambda},V)$ and $(V_{\lambda},V)$, and the associativity morphisms for triples $(V_{\lambda},V,V)$, $(V,V_{\lambda},V)$,$(V,V,V_{\lambda})$ with $V_{\lambda}$ an arbitrary irreducible of the fusion category.
@@
     is a ribbon tensor equivalence by application of the indicated theorems.
-The definition of unitary coboundary weak Hopf algebra, its restricted version named unitary coboundary weak Hopf algebra, are given in Sects.
+The definition of unitary coboundary weak Hopf algebra, its restricted version named unitary compatible coboundary weak Hopf algebra, are given in Sects.
 29, 30 in [11], by abstracting features of the hermitian structure of $U_{q}({\mathfrak{g}})$ for $|q|=1$, discovered by Wenzl in [91] in his proof of unitarity of Kirillov inner product, for the associated fusion category as certain roots of unity, based on a complex Drinfeld coboundary matrix.
@@
 The weak Hopf algebras constructed in [10] have been extended to all Lie types in Sections 29, 30 of [11] with uniform methods for all the Lie types.
+The construction of associativity morphisms satisfying the two hexagonal equations with respect to natural braiding morphisms associated to representations of the universal enveloping algebra $U({\mathfrak{g}})[[h]]$ of a simple Lie algebra (in a formal setting) with classical coproduct, were first solved by Drinfeld with the introduction and construction of quasi-Hopf algebras, via the Knizhnik-Zamolodchikov differential equation.
+This category is known as Drinfeld category, motivated by previous work by Kohno on monodromy of a connection associated to the vector representation for all ${\mathfrak{g}}$ of classical type [62].
+Drinfeld developed the remarkable and beautiful Drinfeld-Kohno theorem connecting $U({\mathfrak{g}})[[h]]$ with its natural $R$-matrix and non-trivial co-associator with the quantized universal enveloping algebra $U_{h}({\mathfrak{g}})$ via a Drinfeld twist, describing a ribbon braided tensor equivalence.
+In this case, rigidity of Drinfeld category is derived from the equivalence constructed by Drinfeld in Drinfeld-Kohno theorem [23], [25], [26].
 Our algebras are examples of Mack-Schomerus weak quasi-Hopf algebras introduced in [64] as generalization of Drinfeld quasi-Hopf algebras with non-unital coproduct.
@@
 Moore and Seiberg conjectured some analytic properties of the quantum fields and described the construction of a modular category associated to a general conformal field theory.
-The mathematical contribution by Kohno, Tsuchiya and Kanie [84], and the subsequent works by A.
+In the same paper [62] Kohno had also considered monodromy of $n$-point functions of a conformal field theory on ${\mathbb{P}}^{1}$ associated to the affine Lie algebra $\hat{\mathfrak{s}l}_{2}$.
+The mathematical contribution by Tsuchiya and Kanie [88], and the subsequent works by A.
 Wassermann [90] and Toledano-Laredo [87] on affine Lie algebras, compared with the structure that had previously emphasized by Wenzl in the study ofquantum group fusion categories and their unitary structures [91] that led to our approach.
@@
 We do not use the KZ equation to this aim, or in any part of our work, other than the fact that the KZ equation was used by Drinfeld, Kazhdan and Lusztig, and by Huang and Lepowsky in the setting of affine vertex operator algebras, to construct the structure of a braided tensor category.
+Similarities with the original Drinfeld-Kohno theorem also concern the identification method between the two associativity morphisms, in our case the twisted structure of that arising from quantum groups and that coming from Huang Lepowsky tensor structure both on the Zhu algebra.
+In Proposition 3.16 of [25], Drinfeld identifies the most difficult parts of the corresponding strictures for Drinfeld category using his uniqueness Theorem 3.15 based on cohomology of the same paper.
+In our case, we similarly develop a uniqueness theorem but for the associativity morphisms which we verify of our weak Hopf algebras $A_{W}({\mathfrak{g},q})$ and with very different methods, based on the generating representation $V$ for which we develop cohomological reduction Theorem 8.19 in [11] and extension to quantum case and different Lie types of Schur-Weyl duality.
+We discuss this more in the next two sections.
 9. Differences between the original Drinfeld-Kohno theorem and our analogue Theorem 28.4 in the treatment of the associativity morphisms, to prove (c).
@@
 A summary of these results and references to the original papers for the vector representation for all the classical Lie types may be found in Theorem 10.2.5 of [9].
-The centralizer algebra of the fundamental representation for types $B$ and $D$ is studied in [81], [89].
+The centralizer algebra of the fundamental representation for types $B$ and $D$ involve difficult work by Wenzl and his coauthors.
+Central theorems are Theorem 5.2 in [94] for the even case, and the previous Theorem 3.3 in [93] for the odd case.
+A special case has been studied by Rowell and Wenzl in [84].
 Type $G_{2}$ was shown by Lehrer and Zhang [63], Morrison [68], and Martirosyan and Wenzl [65].
@@
 Other interesting future developments may regard applications in noncommutative geometry in the sense of Connes [13] of our weak Hopf algebras $A_{W}({\mathfrak{g}},q)$ associated to the quantum group fusion category ${\mathcal{C}}({\mathfrak{g}},q)$ or their twisted weak quasi-Hopf algebra structures on the Zhu algebra of an affine vertex operator algebra at a positive integer level $V_{{\mathfrak{g}}_{k}}$ to the study of Dirac operators, by expanding an analog in conformal field theory of the groundbreaking work by Neshveyev and Tuset on Dirac operators for compact quantum groups in the setting of Kazhdan-Lusztig theorem for real values of the quantization parameter $q$ [72].
- [^1] The claims, arguments, and reasoning of the recent papers [6], [7], [8] are essentially dependent on the results and in-depth discussions between S.C.
+ [^1] Claims, arguments, and reasoning of the recent papers [6], [7], [8] are essentially dependent on the results and in-depth discussions between S.C.
 and C.P.
 that took place from 2018 to 2022 on the contents of an entire project including [11] and the topics of the above papers, during a period in which S.C.
-was a co-author, before breaking the collaboration in early 2023, with a minor contribution in a draft of a single section.
+was a co-author, before breaking the collaboration in early 2023 by his decision, with a minor contribution in a draft of a single section.
 In addition, C.P.
-department digital devices were subject to a serious hacking attack for several months in 2021 during the previous collaboration, which was recognized by authorities, that severely damaged her intellectual property and health.
+department digital devices were subject to a serious hacking attack for several months in 2021 during the previous collaboration, which was recognized by authorities, that severely damaged her intellectual property and health, and the hacked material has not been restituted.
 Acknowledgements The author is grateful to Yi-Zhi Huang for several comments on the proof of modularity for fusion categories of affine vertex operator algebras, and the history of Finkelberg-Kazhdan-Lusztig theorem.
 She is also grateful to Marco V.
-Giannone for pointing out to her incompleteness of the proof of Lemma 7.10 in version 7 of [11] and to Hans Wenzl for kind email correspondence and for references on duality for tensor powers of the fundamental representation of Drinfeld-Jimbo quantum groups for the Lie types $B$, $D$, $G_{2}$, leading to the present paper.
+Giannone for pointing out to her incompleteness of the proof of Lemma 7.10 in version 7 of [11] and to Hans Wenzl for kind email correspondence on the centraliser algebras for the spinor representations, and for references on duality for tensor powers of the fundamental representation of Drinfeld-Jimbo quantum groups for the Lie types $B$, $D$, $G_{2}$, leading to the present paper, and to Makoto Yamashita for interesting discussions on [40] and [92] during his visit at the math department of Rome, Sapienza, and interesting comments on our work.
 Finally she thanks Pavel Etingof for informing her of [29].

### v4

@@
 Abstract.
-We summarize our results in [11] on a problem that Sergio Doplicher posed in the 90s in the setting of conformal nets, and on a related problem posed by Yi-Zhi Huang in the setting of vertex operator algebras, see e.g.
-[55].
+We explain the structure of proof of our results on a problem that Sergio Doplicher posed in the 90s in the setting of conformal nets, and on a related problem posed by Yi-Zhi Huang in the setting of vertex operator algebras.
 The latter asks to find a direct construction of an equivalence due to Finkelberg between the quantum group modular fusion category at certain roots of unity and the fusion categories of affine vertex operator algebras at positive integer levels.
 The former asks for the construction of a quantum gauge group and field algebra for the category of localized endomorphisms of a conformal net, extending Doplicher-Roberts theory of the compact gauge group and field algebra in high dimensional algebraic QFT.
-In this summary we mostly focus on the affine VOA aspects (corresponding to the loop group conformal nets).
-We solve Huang’s problem for all the Lie types for which it is known that the centralizer algebras of the non-negligible tensor powers of a given generating object $V$ for the quantum group fusion category (a list of $V$ for which fusion tensor powers are well defined has been given by Hans Wenzl in [91] for all the Lie types), is generated by the representation of the braid group associated to the $R$-matrix, and for Huang-Lepowsky ribbon braided tensor structure for the vertex operator algebras.
+In this summary we mostly focus on the affine VOA.
+We solve Huang’s problem for all the Lie types for which it is known that the centralizer algebras of the non-negligible tensor powers of a given generating object $V$ for the quantum group fusion category (a list of $V$ for which fusion tensor powers are well defined has been given by Hans Wenzl for all the Lie types), is generated by the representation of the braid group associated to the $R$-matrix, and for Huang-Lepowsky ribbon braided tensor structure for the vertex operator algebras.
 In the classical type $A$, this property becomes the classical Schur-Weyl duality for the vector representation, and was used by Doplicher and Roberts to construct ${\rm SU}(d)$ as a special case of their unique compact gauge group and field algebra in algebraic QFT.
@@
 Our approach may be regarded in the setting of Connes noncommutative geometry.
-Lack of knowledge of the braid group generation property (i.e.
-for the Lie types $E$ and $F$) allows us to show equivalence with Huang-Lepowsky ribbon braided tensor structure for all the ribbon morphisms, and part of those braiding morphisms, and associativity morphisms for which $V$ takes one or two components, respectively.
-Our work also gives a unitary modular tensor category structure on the linear category of modules of the Zhu algebra associated to the minimum energy functor (or on the vertex operator algebra) for all the Lie types, equivalent to that from quantum groups.
-This algebra becomes a unitary weak quasi-Hopf algebra, solving a problem posed by Frenkel and Zhu for all the Lie types [35].
-Our main tool is the construction of two semisimple quantum gauge groups and an equivalence of their representation categories: The quantum group fusion category provides a unitary coboundary weak Hopf algebra associated to Wenzl functor in [91].
-We then apply a de-quantization continuous curve of Wenzl and a Drinfeld twist method to reach the Zhu algebra.
-The structures we find in both contexts extend Doplicher-Roberts symmetric functors for rigid symmetric tensor categories with simple unit which play a central role in their duality theory.
+We also solve a problem by Frenkel and Zhu on the construction of a (unitary) weak quasi-Hopf algebra structure on the Zhu algebra, associated to the minimum energy functor for all the Lie types, equivalent to that from quantum groups.
+Our main tool is the construction of a unitary coboundary weak Hopf $C^{*}$-algebra in a sense extending Drinfeld coboundary quasi-Hopf algebras on a formal variable, associated to Wenzl functor for the quantum group fusion category, and of a Drinfeld twist.
+We apply a de-quantization continuous curve of Wenzl and a Drinfeld twist method to reach the Zhu algebra.
+The two Drinfeld coboundary symmetries we find in both contexts extend Doplicher-Roberts symmetric functors for rigid symmetric tensor categories with simple unit which play a central role in their duality theory.
 Dedicated to Sergio Doplicher, with gratitude
@@
 A main difficulty in constructing tensor categories from models of conformal field theories is the construction of the associativity morphisms compatible with naturally assigned braiding morphisms via the two hexagonal equations.
+Another main difficulty is to show rigidity of the category.
 Let ${\mathfrak{g}}$ be a complex simple Lie algebra.
-This problem was first solved by Drinfeld in [23], [25] [26] for the Drinfeld category with the introduction of the notion and construction of non-trivial examples of quasi-Hopf algebras over $U({\mathfrak{g}})[[h]]$, via the Knizhnik-Zamolodchikov differential equation which is satisfied by the correlation functions of the WZW model.
+These problems were first solved by Drinfeld in [25], [27] [28] for the Drinfeld category with the introduction of the notion and construction of non-trivial examples of quasi-Hopf algebras over $U({\mathfrak{g}})[[h]]$, via the Knizhnik-Zamolodchikov differential equation which is satisfied by the correlation functions of the WZW model.
 Drinfeld used this equation to construct the associativity morphisms compatible with naturally assigned braiding morphisms.
-Furthermore Drinfeld established his remarkable Drinfeld-Kohno theorem expanding an earlier theorem by Kohno [62] stating that the representation of the braid group arising from monodromy of the KZ equation is equivalent to that arising from the action of the $R$-matrix of $U_{h}({\mathfrak{g}})$ with respect to the vector representation, for the classical Lie types.
+Furthermore he established rigidity of this category by showing a remarkable equivalence with the module category arising from the quantized universal enveloping algebra $U_{h}({\mathfrak{g}})$, which is a Hopf algebra.
+To construct the equivalence, he used Drinfeld coboundary matrix ${\overline{R}}$, which is a modification of the $R$-matrix by a square root of the coboundary defined by the ribbon structure.
+This is his remarkable Drinfeld-Kohno theorem, expanding an earlier theorem by Kohno [69] stating that the representation of the braid group arising from monodromy of the KZ equation is equivalent to that arising from the action of the $R$-matrix of $U_{h}({\mathfrak{g}})$ with respect to the vector representation, for the classical Lie types.
 Drinfeld-Kohno theorem connects via a Drinfeld twist, the quasi-Hopf algebra $U({\mathfrak{g}})[[h]]$ with the quantized universal enveloping algebra $U_{h}({\mathfrak{g}})$ which is a Hopf algebra.
-In this case, rigidity of Drinfeld category is derived from the equivalence constructed in Drinfeld-Kohno theorem and rigidity of the representation category of $U_{h}({\mathfrak{g}})$.
+As above mentioned, rigidity of Drinfeld category is derived from the equivalence constructed in Drinfeld-Kohno theorem and rigidity of the representation category of $U_{h}({\mathfrak{g}})$.
 Major work was carried out in a series of papers by Kazhdan and Lusztig in the mid 90s in the setting of affine Lie algebras.
@@
 Def.
-19.1 and 20.2 in [11] and references therein.
+20.1 and 21.2 in [13] and references therein.
 Shortly after, Finkelberg considered a certain semisimple subquotient $\tilde{\mathcal{O}}_{-k}$ of the representation category constructed by Kazhdan and Lusztig, paralleling the subquotient construction in the setting of quantum groups, and described a passage $\tilde{\mathcal{O}}_{k}\to\tilde{\mathcal{O}}_{-k}$ claimed as a braided tensor equivalence from the category $\tilde{\mathcal{O}}_{k}$ of certain integral modules of the affine Lie algebra at the positive integer level $k$ to the subquotient rigid braided tensor category $\tilde{\mathcal{O}}_{-k}$.
@@
 In the process, knowledge of the explicit equivalence gives an understanding of the transported unitary structure.
-We have results on the structure of the unitary structure of the affine vertex operator algebra at positive integer levels transported by our equivalence, which asserts that it must correspond to the classical compact real form, and necessarily equals the unitary structure previously directly constructed by Gui, see [11] for more details and references.
+We have results on the structure of the unitary structure of the affine vertex operator algebra at positive integer levels transported by our equivalence, which asserts that it must correspond to the classical compact real form, and necessarily equals the unitary structure previously directly constructed by Gui, see [13] for more details.
+Our starting point to approach this problem arose from a related problem generously posed by Sergio Doplicher in the 90s to the author in a different setting, that of algebraic quantum field theory.
+We try to explain this problem.
 Our approach to Huang problem is based on interpreting Finkelberg-Kazhdan-Lusztig equivalence as an extension to quantum groups of Doplicher-Roberts duality theory for the important class of WZW models that are described by affine Lie algebras at positive integer levels.
 We shall briefly come back to the previous history of these important models in physics.
-Our starting point to approach this problem arose from a related problem generously posed by Sergio Doplicher in the 90s to the author in a different setting, that of algebraic quantum field theory.
-We try to explain this problem.
 Algebraic quantum field theory had been initiated by Haag and Kastler in [47] in the 60s, and developed by founding breakthrough works by Doplicher, Haag, and Roberts [17], [18], [19], [20] in the 70s.
@@
 Rigidity, called charge conjugation in this setting, describes particle-antiparticle duality.
-In their papers rigidity is derived as a weak limit of inner automorphisms induced by transfer operators moving the support of the given localized endomorphism $\rho$ to infinity.
+In their papers rigidity is derived as a weak limit of inner automorphisms induced by transfer operators moving the support of the given localized endomorphism $\rho$ to (space-like) infinity.
 This gives a left inverse $\Phi_{\rho}$ of $\rho$, which is a completely positive map.
@@
 The left inverse may be described in the category as a map induced by the adjoint action of a solution of the conjugate equations on a intertwiner translated to the right by the tensor product.
-It was shown that charge conjugation implies deep and useful properties for the classification of the statistics parameter (which is the scalar $\Phi_{\rho}(\varepsilon)$, with $\varepsilon$ the action of the generator of the permutation group in the algebra of intertwining operators of the tensor powers of $\rho$), and of charges into charges of para-Bose or para-Fermi type.
-Because of the representation of the permutation group in the category, together with their results on the classification of the statistics parameter, which was shown to the the reciprocal of a positive integer $d$ (the order of the parastatistics) up to a sign (Bose or Fermi cases), and the $C^{*}$-structure, in the 70s it arose the question of whether that was the category of representations of a compact group, because those data arising from a single localised endomorphism were also satisfied by the representation category of a compact Lie subgroup of ${\rm SU}(d)$, that would be a gauge group of global symmetries.
+It was shown that the values of the statistics parameter (which is the constant $\Phi_{\rho}(\varepsilon)$, with $\varepsilon$ the action of the generator of the permutation group in the algebra of intertwining operators of the tensor powers of $\rho$), can be completely classified, and it can only take the values given by the reciprocal of a positive integer $d$ (the order of the parastatistics) up to a sign which classifies charges of para-Bose or para-Fermi type.
+This analysis also needed to prove that the statistics parameter can not take the value 00 and was excluded by physically motivated conditions [20], [48].
+Because of the representation of the permutation group in the category, together with their results on the classification of the statistics parameter, and the $C^{*}$-structure, in the 70s it arose the question of whether that was the category of representations of a compact group, because those data arising from a single localised endomorphism were also satisfied by the representation category of a compact Lie subgroup of ${\rm SU}(d)$, that would be a gauge group of global symmetries.
 They considered the abelian case in the 70s [19].
@@
 The analogy cannot be performed in a strict way due to new difficulties in the quantum case, see the mentioned papers and [13].
-One important difference was that for rigid braided tensor categories we do not have a notion of braided tensor functor beyond the symmetric ones, because the category of vector spaces does not admit a braided symmetry.
+One important difference was that for rigid braided tensor categories we do not have braided tensor functors to the categories of vector spaces, beyond the symmetric ones, because the category of vector spaces does not admit a braided symmetry which is not a permutation symmetry.
 Giving up a compatibility property in the symmetric case implies lack of uniqueness of the functor, and thus of the gauge group.
@@
 To treat the difficulty of finding a replacement of Doplicher-Roberts symmetric functors, in [12] and [13] we have taken into account Drinfeld notion of coboundary symmetry $\overline{R}$, which plays a role in the construction of Drinfeld twist in Drinfeld-Kohno theorem [27].
-This is a symmetry of the category obtained modifying the braided symmetry defined by the $R$-matrix by a square root of the $2$-coboundary defined by the ribbon structure, and by construction the coboundary symmetry is an involutive symmetry, similarly to the braided symmetry.
+This is a symmetry of the category obtained modifying the braided symmetry defined by the $R$-matrix by a square root of the $2$-coboundary defined by the ribbon structure, and by construction the coboundary symmetry is an involutive symmetry, similarly to the permutation symmetry.
 This symmetry finally allows for a generalization of the notion of symmetric functor.
@@
 This involved in [13] the construction of a Drinfeld twist in our setting, the above mentioned notion that Drinfeld had introduced to study equivalence of the Drinfeld category with the quantum group category in a formal setting.
-We have done this for all the Lie types via our version of Drinfeld-Kohno Theorem 28.4 and applications to affine vertex operator algebras at positive integer level, via Zhu linear equivalence, in [11].
+We have done this for all the Lie types via our version of Drinfeld-Kohno Theorem 30.4 in [13] and applications to affine vertex operator algebras at positive integer level, via Zhu linear equivalence, in [13].
 We refer to Sect.
@@
 This section is important for us as it suggested that in general the associativity morphisms in a category arising from CFT may be of a simple nature from the viewpoint of the associated quantum gauge groups that need to be constructed, and this finally proved to be the case, we shall come back to this with more precision.
-On the other hand, in the process of extending Doplicher-Roberts theory in the setting of compact quantum groups, we had started by embedding in a tensorial way part of the morphisms of the category into the Hilbert spaces [76], [78], [82], and in doing this we had already made a choice of a dimension function, and from that choice we had constructed a tensor functor from that subcategory.
-To consider the full category, the analogue Doplicher-Roberts compact gauge group is described by an ergodic action of a larger quantum group on a quantum homogeneous space described by a $C^{*}$-algebra, rather than a quantum group itself, due to lack of commutativity of this $C^{*}$-algebra [78].
+On the other hand, in the process of extending Doplicher-Roberts theory in the setting of compact quantum groups, we had started by embedding in a tensorial way part of the morphisms of the category into the Hilbert spaces.
+This was the tensor subcategory generated by solutions of the conjugate equations [83], [85], [89].
+Via Tannaka-Krein-Woronowicz duality this subcategory corresponds to Wang and Van Daele orthogonal or universal compact quantum groups $A_{o}(F)$, or $A_{u}(F)$ [97], their representation category has been studied by Banica [2], [3].
+The choice of the compact quantum group arising in this way depends on whether a generating object $\rho$ is selfconjugate as in the case of $A_{o}(F)$ or not as for $A_{u}(F)$.
+In doing this embedding we had already made a choice of a dimension function, the rank of the associated invertible matrix $F$, and from that choice we had constructed a tensor functor from that subcategory.
+To consider the full category, the analogue Doplicher-Roberts compact gauge group is described by an ergodic action of the corresponding quantum group on a quantum homogeneous space rather than a quantum group itself, due to lack of commutativity of this $C^{*}$-algebra [85].
+The homogeneous space is described by a $C^{*}$-algebra, naturally associated to the category and to the embedding of the subcategory.
 This property is likely to hold also in other models arising from conformal field theory starting from the minimum energy functor, unless one can show that certain known obstructions already met in [84] for compact quantum groups, vanish, we shall come back to this in the outlook section 13.
 However, for the WZW models arising from affine Lie algebras at positive integer levels in [12] for the type $A$ case, and in [13] for all the Lie types, we obtain genuine quantum groups as quantum gauge groups, described as semisimple weak Hopf algebras with $R$-matrix, $C^{*}$-structure explicitly associated to Drinfeld coboundary matrix, and with unitary representations of the braid group, derived from the quantum group fusion category.
-We prove that these weak Hopf algebras have representation category equivalent to both the starting fusion category and also to that of the corresponding affine vertex operator algebra for the classical Lie types and $G_{2}$.
-Constructing the equivalence from the former to the latter in the setting of semisimple weak quasi-Hopf algebras as we do in [11] is our proof of Kazhdan-Lusztig-Finkelberg theorem, see Theorem 2.1.
-Among our first references in the context of conformal field theory important for us to start to set up the connection with the work of Wenzl in quantum groups, is the work by Wassermann [90] in the setting of affine Lie algebras, where the key concept is played by the notion of primary field, and its initial term, which we use in a fundamental way to connect quantum group fusion category to representations of the Zhu algebra and their special tensor products with the fundamental representation, see Sect 33 in [11].
-Most importantly, we describe a connection between Wenzl and Wassermann work via Wenzl quantisation curve in theorem 33.7, to which we refer the reader also for important historical references.
-The connection from affine Lie algebra setting to Huang-Lepowsky tensor product theory may be found in Sections 35 and 36.
+We prove that these weak Hopf algebras have representation category equivalent to both the starting fusion category for all the Lie types and also to that of the corresponding affine vertex operator algebra with ribbon braided Huang-Lepowsky tensor structure, for the classical Lie types and $G_{2}$.
+Constructing the equivalence from the former to the latter in the setting of semisimple weak quasi-Hopf algebras as we do in [13] for these Lie types is our proof of Kazhdan-Lusztig-Finkelberg theorem, see Theorem 2.1.
+Among our first references in the context of conformal field theory important for us to start to set up the connection with the work of Wenzl in quantum groups, is the work by Wassermann [98] in the setting of affine Lie algebras, where the key concept is played by the notion of primary field, and its initial term, which we use in a fundamental way to connect quantum group fusion category to representations of the Zhu algebra and their special tensor products with the fundamental representation, see Sect.
+35 in [13].
+Most importantly, we describe a connection between Wenzl and Wassermann work via Wenzl quantisation curve, and building on their work in [13] we obtain Theorems 3.1, 3.2, Cor.
+3.3, and Theorem 3.4 which compares with the tensor product structure by Huang and Lepowsky, to which we refer the reader also for important historical references.
+The connection from affine Lie algebra setting to Huang-Lepowsky tensor product theory may be found in Sections 34–39 in [13].
 Our abstract duality methods need new further tools as we shall see from the theory of quantum groups, mainly from the work of Drinfeld and Wenzl, that are not available in the theory of groups.
@@
 We notice on the other hand that a stronger property is needed to establish the connection, than the purely categorical characterisation of functors in Sect.
-27.
-This property is expressed in terms of quantum groups, and the first example, in a non-semisimple setting, is the hermitian structure of $U_{q}({\mathfrak{g}})$ at a root of unity by Theorem 29.9.
+29 in [13].
+This property is expressed in terms of quantum groups, and the first example, in a non-semisimple setting, is the hermitian structure of $U_{q}({\mathfrak{g}})$ at a root of unity by Theorem 31.9 in [13].
 This property says that the -involution enjoys a precise anticommutation relation (rather than commutation relation as is familiar in Hopf $C^{*}$-algebras) with the coproduct and was first noticed in [99].
-We axiomatize this property in section 28, as a compatibility property of hermitian coboundary weak quasi-Hopf algebras.
-Further examples derived from this examples are our weak Hopf and quasi-Hopf algebras, see Sections 29, 30, 31.
-This is not a direct consequence of the corresponding property of $U_{q}({\mathfrak{g}})$, as its coproduct is not compatible with the fusion tensor product.
-As above motivated, we shall derive more structure than rigidity and modularity, such as unitarization of the vertex operator algebra representation category from the corresponding quantum group fusion category, via the explicit equivalence, see Sects.
-21, 22, 24 in [11].
+We axiomatize this property in Sect.
+30 in [13], as a compatibility property of hermitian coboundary weak quasi-Hopf algebras.
+Further examples are our weak bi-algebras $A_{W}({\mathfrak{g}},q)$ associated to Wenzl functor $W:{\mathcal{C}}({\mathfrak{g}},q)\to{\rm Vec}$ and to the quasi-bialgebras structures on the Zhu algebra $A({V_{{\mathfrak{g}}_{k}}})$ associated to Zhu functor ${\rm Rep}({V_{{\mathfrak{g}}_{k}}})\to{\rm Vec}$ via our Drinfeld twist construction, see Sections 32, 34 in [13].
+The anticommutation property of the -involution and coproduct of $A_{W}({\mathfrak{g}},q)$ is not a direct consequence of the corresponding property of $U_{q}({\mathfrak{g}})$, as its coproduct is not compatible with the fusion tensor product.
+We need some delicate analysis of the fusion tensor product described in [99].
+For $A({V_{{\mathfrak{g}}_{k}}})$, we show that the twist makes the coproduct of $A({V_{{\mathfrak{g}}_{k}}})$ symmetric, and this is a result analogous to the symmetry property of the usual coproduct of Drinfeld quasi-Hopf algebra $U({\mathfrak{g}})[[h]]$.
+For details we refer to the cited sections.
+On the other hand, as an algebra $A_{W}({\mathfrak{g}},q)$ is a natural quotient of $U_{q}({\mathfrak{g}})$, and the antipode of $A_{W}({\mathfrak{g}},q)$ is directly derived from that of $U_{q}({\mathfrak{g}})$.
+In this way, $A_{W}({\mathfrak{g}},q)$ becomes a weak Hopf algebra.
+It follows that ${\rm Rep}(A_{W}({\mathfrak{g}},q))$ becomes a rigid ribbon braided tensor category naturally equivalent to ${\rm Rep}(A({V_{{\mathfrak{g}}_{k}}}))$ via our twist, and therefore so does ${\rm Rep}({V_{{\mathfrak{g}}_{k}}})$ via Zhu equivalence.
+As above motivated, we shall derive more structure beyond rigidity and modularity on ${\rm Rep}({V_{{\mathfrak{g}}_{k}}})$ from the quantum group fusion category, a further example is unitarization of the vertex operator algebra representation category from the corresponding quantum group fusion category, via the explicit equivalence, see Sects.
+22, 23, 24, 26 in [13].
 To this aim it is important that our equivalence passes through the weak quasi-Hopf algebras, and the fact that these are naturally associated to the categories.
@@
 16 in [13] treats this for general tensor categories that admit submultiplicative dimension functions (called weak dimension functions).
-Moreover our equivalence, which is given by an analogue of the Drinfeld twist in [23], [25], at the level of the associated weak Hopf or quasi-Hopf algebras gives an understanding of the unitary structures of representation categories of affine vertex operator algebras as classical compact real forms, or untwisted.
+Our equivalence, since it is given by a Drinfeld twist analogous to [25], [27], and is also motivated by the requirement of making the unitary structure classical, at the level of the associated weak Hopf or quasi-Hopf algebras gives an understanding of the unitary structures of representation categories of affine vertex operator algebras as classical compact real forms, or untwisted.
 We refer to the outlook section for some more information on possible extension of our viewpoint to models beyond the affine vertex operator algebras at positive integer levels, see also [13].
@@
 A main idea that we follow is to consider the fundamental representation $V$ of ${\mathfrak{g}}$ for each Lie type as in [99] where the author explains its generating property and good fusion structure with respect to tensor products with an arbitrary irreducible representation allowing a well-defined truncation procedure to obtain non-negligible objects, with the aim of studying unitarizability of the fusion category.
-It should also be noted that a main application of that paper is construction of inclusion of subfactors of finite Jones index for all Lie types, extending his previous work.
+It should also be noted that a main application of that paper is construction of inclusion of subfactors of finite Jones index [63] for all Lie types, extending his previous work.
 For all the Lie types we obtain complete identification for the tensor product structure, all the ribbon structure, and those braiding and associativity morphisms where one or two variables are the fundamental representation $V$, respectively, with Huang-Lepowsky ribbon braided tensor structure.
-These are parts (a), (b) of Theorem 2.1.
+These are parts (a), (b), (c) of Theorem 2.1.
 The term weak in the setting of quasi-Hopf algebras was first introduced by Mack and Schomerus motivated by the work of Drinfeld for quasi-Hopf algebras and Drinfeld-Kohno theorem [71].
@@
 In [12], [13], and also in this paper, weak Hopf algebras are understood with a different meaning.
-While a different name would be helpful to avoid confusion, so far we have not been able to find a good name for our algebras to clarify the situation, and we use the term weak Hopf algebras in the sense defined in Sect 9 in [11].
+While a different name would be helpful to avoid confusion, so far we have not been able to find a good name for our algebras to clarify the situation, and we use the term weak Hopf algebras in the sense defined in Sect.
+10 in [13].
 In [13] we develop a general theory of weak Hopf algebras, review and expand the theory of weak quasi-Hopf algebras, mainly Tannaka-Krein aspects, both in the algebraic and in the $C^{*}$-algebraic setting, with non-trivial unitary structure, discuss a suitable Drinfeld-Kohno theorem, and its application to the construction of semisimple weak Hopf and quasi-Hopf algebras with structure naturally compatible with that of the involved ribbon braided tensor categories, for quantum groups at roots of unity and affine vertex operator algebras at positive integer levels, including unitary structure as heredited from the unitary structure of the category due to Wenzl [99].
@@
 We apply our concrete twist to transport unitary structures to representation categories of affine vertex operator algebras at positive integer levels with the tensor structure due to Huang and Lepowsky, for all the mentioned Lie types, by transferring the unitary structure from quantum group fusion category as constructed in [99].
+On the other hand, the remaining Lie types may be either derived from Finkelberg equivalence, making now the ribbon braided fusion category $\tilde{\mathcal{O}}_{k}$ unitary, or treated with our partial results in the setting of Huang-Lepowsky structure for representation categories of vertex operator algebras.
+Direct constructions of unitary structures in this setting have been obtained by Gui as already mentioned [43], [44], [45], [46].
 Our weak Hopf algebras while having a non-unital coproduct, are not in the class of Bohm and Szlachanyi.
@@
 Their counit is antimultiplicative and the antipode satisfy the same relations as in the case of Hopf algebras.
-We prove the complete identification of braiding and associativity morphisms of the two structures for all the classical Lie types $A$, $B$, $C$, $D$, and the case $G_{2}$ in part (c) of Theorem 2.1, by further refining methods of the previous part but we use more structure in Sect.
-$8$ of [11] that has not emerged in other parts.
+As already mentioned, we prove the complete identification of braiding and associativity morphisms of the two structures for all the classical Lie types $A$, $B$, $C$, $D$, and the case $G_{2}$ in part (d) of Theorem 2.1, by further refining methods of the previous part but we use more structure in Sect.
+9 of [13] that has not emerged in other parts.
 As above mentioned, in the type $A$ case, the extra structure that we need to use is known as quantum Schur-Weyl duality between the vector representation and the natural representation of the braid group between the tensor powers of the vector representations.
@@
 Duality with the braid group representation is a concept at the basis of the origin of quantum groups following V. F. R. Jones [64].
-More generally, for Lie types beyond $A$, we need a corresponding duality between the centralizer algebras of the truncated tensor powers of the fundamental representation $V$ of the Drinfeld-Jimbo quantum group $U_{q}({\mathfrak{g}})$ at the root of unity and the natural representation of the braid group in these algebras, and this leads to the restriction stated in (c).
+More generally, for Lie types beyond $A$, we need a corresponding duality between the centralizer algebras of the truncated tensor powers of the fundamental representation $V$ of the Drinfeld-Jimbo quantum group $U_{q}({\mathfrak{g}})$ at the root of unity and the natural representation of the braid group in these algebras, and this leads to the restriction stated in (d).
 This allows us to apply our abstract uniqueness Theorem 9.19 in [13] for braided tensor structures in a pre-tensor category.
@@
   • (a)
-    The representation category ${\rm Rep}(A(V_{{\mathfrak{g}}_{k}}))$ admits a natural structure of ribbon braided tensor category obtained by transporting Huang-Lepowsky structure to ${\rm Rep}(A(V_{{\mathfrak{g}}_{k}}))$.
+    The linear representation category ${\rm Rep}(A(V_{{\mathfrak{g}}_{k}}))$ admits a natural structure of ribbon braided tensor category ${\rm Rep}_{\rm HL}(A(V_{{\mathfrak{g}}_{k}}))$ obtained as a transport from Huang-Lepowsky structure on ${\rm Rep}_{\rm HL}(V_{{\mathfrak{g}}_{k}})$ to ${\rm Rep}(A(V_{{\mathfrak{g}}_{k}}))$.
 In this way, Zhu’s linear equivalence
-    $Z:{\rm Rep}(V_{{\mathfrak{g}}_{k}})\to{\rm Rep}(A(V_{{\mathfrak{g}}_{k}}))$
+    $Z:{\rm Rep}_{\rm HL}(V_{{\mathfrak{g}}_{k}})\to{\rm Rep}_{\rm HL}(A(V_{{\mathfrak{g}}_{k}}))$
     admits a natural structure of ribbon braided tensor equivalence.
   • (b)
-    Let $A(V_{{\mathfrak{g}}_{k}})$ be endowed with the twisted unitary compatible coboundary weak quasi-Hopf structure induced by the weak Hopf algebra $A_{W}({\mathfrak{g}},q)$ via Wenzl path and the construction of the Drinfeld twist, following Theorem 31.1 in [11] (a), and let ${\rm Rep}(A(V_{{\mathfrak{g}}_{k}}))$ be endowed with the corresponding ribbon braided tensor structure.
-    Then ${\rm Rep}(A(V_{{\mathfrak{g}}_{k}}))$ becomes a unitary coboundary modular tensor category in this way with the structure derived from ${\mathcal{C}}({\mathfrak{g}},q)$ and the equivalence of Theorem 31.1 in [11],
-    ${\rm Rep}(A(V_{{\mathfrak{g}}_{k}}))\xrightarrow{}{\rm Rep}(A_{W}({\mathfrak{g}},q)).$
-    Moreover the transport in (a) preserves the tensor products, the ribbon for all objects, the braid morphisms for pairs $(V_{\lambda},V)$ and $(V_{\lambda},V)$, and the associativity morphisms for triples $(V_{\lambda},V,V)$, $(V,V_{\lambda},V)$,$(V,V,V_{\lambda})$ with $V_{\lambda}$ an arbitrary irreducible of the fusion category.
+    The Zhu algebra $A(V_{{\mathfrak{g}}_{k}})$ becomes a twisted unitary compatible coboundary weak quasi-Hopf algebra with structure induced by the weak Hopf algebra $A_{W}({\mathfrak{g}},q)$ via Wenzl path and the construction of the Drinfeld twist.
+Let ${\rm Rep}_{\rm QG}(A(V_{{\mathfrak{g}}_{k}}))$ be endowed with the corresponding ribbon braided tensor structure.
+    At the level of representation categories, ${\rm Rep}_{\rm QG}(A(V_{{\mathfrak{g}}_{k}}))$ becomes a unitary coboundary modular tensor category in this way with the structure derived from ${\mathcal{C}}({\mathfrak{g}},q)$ and the equivalence
+    ${\rm Rep}_{\rm QG}(A(V_{{\mathfrak{g}}_{k}}))\xrightarrow{}{\rm Rep}(A_{W}({\mathfrak{g}},q)),$
+    obtained in Theorem 3.2 in [13]. Moreover ${\rm Rep}_{\rm QG}(A(V_{{\mathfrak{g}}_{k}}))$ and ${\rm Rep}_{\rm HL}(A(V_{{\mathfrak{g}}_{k}}))$ have the same tensor product bifunctor.
   • (c)
-    If ${\mathfrak{g}}$ is of one of the Lie types $A$, $B$, $C$, $D$, $G_{2}$ then the transport preserves the braided tensor structure for all objects.
-Thus $Z$ is a ribbon braided tensor equivalence.
-Specifically, the composition
+    The linear equivalence
+    $Z:{\rm Rep}_{\rm HL}(V_{{\mathfrak{g}}_{k}})\to{\rm Rep}_{\rm QG}(A(V_{{\mathfrak{g}}_{k}}))$
+    endowed with the same structure maps as in (a) but w.r.t.
+the new ribbon braided tensor structure of ${\rm Rep}_{\rm QG}(A(V_{{\mathfrak{g}}_{k}}))$ described in (b), satisfies the ribbon braided tensor equivalence equations for the ribbon structure for all objects, the braid morphisms for pairs $(V_{\lambda},V)$ and $(V_{\lambda},V)$, and the associativity morphisms for triples
+    $(V_{\lambda},V,V),\quad(V,V_{\lambda},V),\quad(V,V,V_{\lambda})$
+    with $V_{\lambda}$ an arbitrary irreducible object of the fusion category.
+  • (d)
+    If ${\mathfrak{g}}$ is of one of the Lie types $A$, $B$, $C$, $D$, $G_{2}$ then Zhu’s equivalence in (c) is a ribbon braided tensor equivalence. Specifically, the composition
 […]
@@
 The definition of unitary coboundary weak Hopf algebra, its restricted version named unitary compatible coboundary weak Hopf algebra, are given in Sects.
-29, 30 in [11], by abstracting features of the hermitian structure of $U_{q}({\mathfrak{g}})$ for $|q|=1$, discovered by Wenzl in [91] in his proof of unitarity of Kirillov inner product, for the associated fusion category as certain roots of unity, based on a complex Drinfeld coboundary matrix.
-The former more general notion admits a categorical characterization generalizing Doplicher-Roberts symmetric functors.
-We would like to notice that the passages described in part (c) of Theorem 2.1 are explicit.
+31, 32 in [13], by abstracting features of the hermitian structure of $U_{q}({\mathfrak{g}})$ for $|q|=1$, discovered by Wenzl in [99] in his proof of unitarity of Kirillov inner product, for the associated fusion category as certain roots of unity, based on a complex Drinfeld coboundary matrix.
+The more general notion admits a categorical characterization generalizing Doplicher-Roberts symmetric functors.
+For part (b), we refer the reader to Theorem 3.2 (a) in [13] and further indications therein where the statement is explained.
+We would like to notice that the passages described in (2.1) are explicit.
 The first and the last equivalences are due respectively, to the work by Zhu combined with that of Huang and Lepowsky.
@@
 This anticommutativity is derived from the same property of the -involution of the non-semisimple $U_{q}({\mathfrak{g}})$, first pointed out by Wenzl [99].
-The weak Hopf algebras constructed in [10] have been extended to all Lie types in Sections 29, 30 of [11] with uniform methods for all the Lie types.
+The weak Hopf algebras constructed in [12] have been extended to all Lie types in Sections 31, 32 of [13] with uniform methods for all the Lie types.
 The construction of associativity morphisms satisfying the two hexagonal equations with respect to natural braiding morphisms associated to representations of the universal enveloping algebra $U({\mathfrak{g}})[[h]]$ of a simple Lie algebra (in a formal setting) with classical coproduct, were first solved by Drinfeld with the introduction and construction of quasi-Hopf algebras, via the Knizhnik-Zamolodchikov differential equation.
@@
 We start reviewing these Tannakian reconstruction theorems from the viewpoint of Majid and Haring-Oldenburg in Sect.
-5 in [11], where the reader may find indication of the original references.
+6, 7, in [13], where the reader may find indication of the original references.
 Moreover Kazhdan and Wenzl theory is also important to us, and the reader may consult Sect.
@@
 Wassermann [98] and Toledano-Laredo [94] on affine Lie algebras, compared with the structure that had previously emphasized by Wenzl in the study ofquantum group fusion categories and their unitary structures [99] that led to our approach.
-The reader may find reference to more of the several original papers in mathematics and physics in Sections 31–36 in [11].
+The reader may find reference to more of the several original papers in mathematics and physics in Sections 34–39 in [13].
 4. The general plan of our proof.
-We briefly explain here the plan of our proof of the main theorem for parts (a), (b), and (c). In the following sections we describe the main ideas to prove these parts in greater detail.
+We briefly explain here the plan of our proof of the main theorem for parts (a), (b), (c), and (d). In the following sections we describe the main ideas to prove these parts in greater detail.
 Part (a) is an application of a general construction of a natural structure of a (ribbon, braided) tensor equivalence on a given equivalence ${\mathcal{C}}\to{\rm Rep}(A)$ of semisimple linear categories endowed with an inverse linear equivalence ${\rm Rep}(A)\to{\mathcal{C}}$, with $A$ a discrete algebra.
@@
 Complete definitions of unitary coboundary structures on weak quasi-Hopf algebras are given in [13], and are motivated by the work of several authors, among them Drinfeld, Wenzl, Mack and Schomerus.
-See Sections 3, 5, 9, 10, 11, 26, 28 in [11].
 In Sect.
@@
 These are Theorem 9.13 and its extension Theorem 9.19.
-We explain here briefly how Theorem 8.19 applies to imply part (c) of Theorem 2.1.
-The proof of parts (b) and (c) is based on the general theory of semisimple weak quasi-Hopf algebras, weak Hopf algebras, and the theory of Drinfeld twist treated in Sections 3–6, and 8–16 in [11].
+We explain here briefly how Theorem 9.19 applies to imply part (d) of Theorem 2.1.
+The proof of parts (b), (c), and (d) is based on the general theory of semisimple weak quasi-Hopf algebras, weak Hopf algebras, and the theory of Drinfeld twist treated in Sections 5–8, and 10–18 in [13].
 This abstract theory splits in two parts.
-Theory on which part (b) of Theorem 2.1 is based is developed in Sections 3–6, and 9–16.
-The applicative part of the proof of part (b) is developed in Sections 19–36.
-The abstract aspects that lead to the proof of part (c) of Theorem 2.1 are the main subject of Section 8.
-For part (b), we develop an abstract analogue of Drinfeld-Kohno theorem, Theorem 28.4, and apply it to the setting of fusion categories of quantum groups at suitable roots of unity and of module categories of affine vertex operator algebras at positive integer levels.
-We construct weak Hopf algebras $A_{W}({\mathfrak{g}},q)$ associated to the quantum group fusion category for all the Lie types in Section 30, by constructing a weak tensor structure on a linear functor from this category to the category of vector spaces, that emerges in [91] for the special tensor products described in (b).
+Theory on which part (b), (c) of Theorem 2.1 is based is developed in Sections 5–8, and 10–17.
+The applicative part of the proof of part (b), (c) is developed in Sections 20–39.
+The abstract aspects that lead to the proof of part (d) of Theorem 2.1 are the main subject of Section 9.
+For parts (b), (c) we develop an abstract analogue of Drinfeld-Kohno theorem 30.4 in [13], and apply it to the setting of fusion categories of quantum groups at suitable roots of unity and of module categories of affine vertex operator algebras at positive integer levels.
+We construct weak Hopf algebras $A_{W}({\mathfrak{g}},q)$ associated to the quantum group fusion category for all the Lie types in Section 32 in [13], by constructing a weak tensor structure on a linear functor from this category to the category of vector spaces, that emerges in [99] for the special tensor products described in (b), (c).
 The weak tensor structure is constituted by fusion intertwiners with respect to the action of $U_{q}({\mathfrak{g}})$.
@@
 The first way is traditionally referred to as Tannaka-Krein duality.
-One needs a (weak, quasi-) tensor structure $(F,G)$ on ${\mathcal{F}}$, and then apply Theorem 6.6 to obtain a (weak, quasi)-bialgebra structure on $A={\rm Nat}_{0}({\mathcal{F}})$ (with extra structure) with the desired equivalence property.
+One needs a (weak, quasi-) tensor structure $(F,G)$ on ${\mathcal{F}}$, and then apply Theorem 8.6 in [13] to obtain a (weak, quasi)-bialgebra structure on $A={\rm Nat}_{0}({\mathcal{F}})$ (with extra structure) with the desired equivalence property.
 The (weak quasi-) tensor structure
@@
 We discuss this transfer in Sect.
-15 without the need of a weak quasi-tensor structure on ${\mathcal{F}}$, in abstract terms.
+16 in [13] without the need of a weak quasi-tensor structure on ${\mathcal{F}}$, in abstract terms.
 One only needs an inverse linear equivalence ${\mathcal{S}}:{\rm Rep}(A)\to{\mathcal{C}}$ of the Tannakian linear equivalence ${\mathcal{E}}:{\mathcal{C}}\to{\rm Rep}(A)$.
@@
 He also constructed a linear equivalence ${\mathcal{E}}:{\rm Rep}(V)\to{\rm Rep}(A(V))$ and also a natural linear inverse equivalence ${\mathcal{S}}:{\rm Rep}(A(V))\to{\rm Rep}(V)$ in the same paper if the vertex operator algebra $V$ satisfies suitable semisimplicity conditions.
-Section 15 abstracts Zhu linear equivalence and its inverse, by considering a pair of the Tannakian linear equivalence ${\mathcal{E}}:{\mathcal{C}}\to{\rm Rep}(A)$ associated to ${\mathcal{F}}:{\mathcal{C}}\to{\rm Vec}$ and a linear inverse equivalence ${\mathcal{S}}:{\rm Rep}(A)\to{\mathcal{C}}$ and describes how to make ${\rm Rep}(A)$ into a braided tensor category.
+Section 16 in [13] abstracts Zhu linear equivalence and its inverse, by considering a pair of the Tannakian linear equivalence ${\mathcal{E}}:{\mathcal{C}}\to{\rm Rep}(A)$ associated to ${\mathcal{F}}:{\mathcal{C}}\to{\rm Vec}$ and a linear inverse equivalence ${\mathcal{S}}:{\rm Rep}(A)\to{\mathcal{C}}$ and describes how to make ${\rm Rep}(A)$ into a braided tensor category.
 Thus, in the case of vertex operator algebras satisfying semisimplicity conditions, one may reduce the analysis of the category of infinite dimensional representations of the vertex operator algebra $V$ to a category of finite dimensional representations of the Zhu algebra $A(V)$ in a natural way.
@@
 7. A central method in the quantum group and VOA setting: the fundamental generating representation $V$ for all Lie types, following H.Wenzl.
-A central method of proof for both parts (b) and (c) is the use of the fundamental generating representation $V$ and its properties, following primarily the work of Wenzl.
-The paper [91] plays a main role especially for the proof of (b).
-For part (c) we shall later describe results and related works in the literature on which our proof is based.
-Part (b) of Theorem 2.1 gives coincidence of the two braided tensor structures for all Lie types (two pairs of associativity morphisms together with their braiding morphisms satisfying the two hexagonal diagrams) on a specific collection of their variables, in the two settings of quantum groups and vertex operator algebras.
+A central method of proof for parts (b), (c), and (d) is the use of the fundamental generating representation $V$ and its properties, following primarily the work of Wenzl.
+The paper [99] plays a main role especially for the proof of (b), (c).
+For part (d) we shall later describe results and related works in the literature on which our proof is based.
+Parts (b), (c) of Theorem 2.1 gives coincidence of the two braided tensor structures for all Lie types (two pairs of associativity morphisms together with their braiding morphisms satisfying the two hexagonal diagrams) on a specific collection of their variables, in the two settings of quantum groups and vertex operator algebras.
 These collections do not generate additively all the variables of the braiding morphisms or associativity morphisms, because they do not contain all the irreducible objects.
@@
 Thus the uniqueness theorem 9.19 is formulated with suitable assumptions, where again in the application the abstract notion of ${\mathcal{V}}$-pre-associator of CFT-type for the fundamental representation $V$ is used in a central way both on the side of quantum groups and that of vertex operator algebras to verify coincidence of these special braid and associativity morphisms, after application of our Drinfeld-Kohno theorem.
-But this brief description of Theorems 8.13 and Theorem 8.19 put together with coincidence of the two braided tensor structures on special collections of objects proved in part (b), is still not enough to describe our method of proof to complete the proof of part (c) and with it the proof of Theorem 2.1.
+But this brief description of Theorems 9.13 and Theorem 9.19 put together with coincidence of the two braided tensor structures on special collections of objects proved in part (b), (c), is still not enough to describe our method of proof to complete the proof of part (d) and with it the proof of Theorem 2.1.
 We need more that we describe in the following sections.
-The first next step is a somewhat complicated cohomological reduction work done in Theorems 8.13 and 8.19, where we use cohomology of the pentagon equation and its interplay with the hexagonal equations defining the braided symmetry for a general semisimple braided tensor category endowed with a generating object.
+The first next step is a somewhat complicated cohomological reduction work done in Theorems 9.13 and 9.19, where we use cohomology of the pentagon equation and its interplay with the hexagonal equations defining the braided symmetry for a general semisimple braided tensor category endowed with a generating object.
 The cohomological reduction is useful to reduce the problem to the last step, which uncovers a central duality property between the fundamental representation of the quantum group and corresponding representations of the braid group in the intertwining spaces of the fundamental representation, in the various Lie types.
 The first example of this is Schur-Weyl duality in the type $A$ case, and works in the literature that we shall recall greatly extend this first classical example to to other Lie types and to the quantum cases.
-We shall come back to this, but let us anticipate similarities and differences with the original Drinfeld-Kohno theorem, useful for the proof of (b), that will also useful to clarify how we were led to uncover the key role of duality of the fundamental representation $V$, for the proof of the last step (c).
-8. Main tools to prove (b): Analogies and differences between Drinfeld-Kohno Theorem and Theorem 28.4 in [11]
-We next outline general ideas that make our strategy possible, useful to prove part (b) of Theorem 2.1.
+We shall come back to this, but let us anticipate similarities and differences with the original Drinfeld-Kohno theorem, useful for the proof of (b), (c), that will also useful to clarify how we were led to uncover the key role of duality of the fundamental representation $V$, for the proof of the last step (d).
+8. Main tools to prove (b), (c): Analogies and differences between Drinfeld-Kohno Theorem and Theorem 30.4 in [13]
+We next outline general ideas that make our strategy possible, useful to prove part (b), (c) of Theorem 2.1.
 As already explained, the starting point of our work is the paper [99].
@@
 In our case we adopt the idea of Drinfeld coboundary that we describe for complex weak quasi-Hopf algebras in an analytic setting.
-This general theory is considered in Sections 26, 27, 28.
+This general theory is considered in Sections 27, 28, 29, 30.
 On one hand the coboundary matrix in an analytic setting induces a unitary structure which shows striking differences with unitary structures of usual Hopf $C^{*}$-algebras, such as for example the compact quantum groups, where the -involution commutes with the coproduct.
@@
 Similarities with the original Drinfeld-Kohno theorem also concern the identification method between the two associativity morphisms, in our case the twisted structure of that arising from quantum groups and that coming from Huang Lepowsky tensor structure both on the Zhu algebra.
-In Proposition 3.16 of [25], Drinfeld identifies the most difficult parts of the corresponding strictures for Drinfeld category using his uniqueness Theorem 3.15 based on cohomology of the same paper.
+In Proposition 3.16 of [27], Drinfeld identifies the most difficult parts of the corresponding structures for Drinfeld category using his uniqueness Theorem 3.15 based on cohomology of the same paper.
 In our case, we similarly develop a uniqueness theorem but for the associativity morphisms which we verify of our weak Hopf algebras $A_{W}({\mathfrak{g},q})$ and with very different methods, based on the generating representation $V$ for which we develop cohomological reduction Theorem 9.19 in [13] and extension to quantum case and different Lie types of Schur-Weyl duality.
 We discuss this more in the next two sections.
-9. Differences between the original Drinfeld-Kohno theorem and our analogue Theorem 28.4 in the treatment of the associativity morphisms, to prove (c).
+9. Differences between the original Drinfeld-Kohno theorem and our analogue Theorem 30.4 in the treatment of the associativity morphisms, to prove (d).
 Differences regard mainly the treatment of the associativity morphisms, that in our case as before explained resorts to determining the braiding morphisms via an analogue of Drinfeld square root construction of the coboundary matrix, and to braid group duality for $V$ and to cohomology of a braided tensor structure of a generating representation.
@@
 Differently from the original Drinfeld-Kohno theorem, we do not study the problem of equivalence associativity morphisms in our version version of Drinfeld-Kohno theorem.
-We do this in Section 8 in [11] with the mentioned different ideas that appeal to duality properties of the fundamental representation $V$ with respect to the representation of the braid group in the tensor powers of $V$.
+We do this in Sect.
+9 in [13] with the mentioned different ideas that appeal to duality properties of the fundamental representation $V$ with respect to the representation of the braid group in the tensor powers of $V$.
 More specifically, we first construct semisimple weak Hopf algebras $A_{W}({\mathfrak{g}},q)$ associated to fusion categories of quantum groups at roots of unity for all Lie types.
@@
 Drinfeld coboundary is closely related to the ribbon structure, and thus it is relevant for the ribbon preserving property of the KLF equivalence, similarly to the case of Drinfeld-Kohno theorem.
-10. Main tools to prove (c): cohomological reduction of a braided tensor structure of a category with a generating object $V$ to braid group duality of $V$
+10. Main tools to prove (d): cohomological reduction of a braided tensor structure of a category with a generating object $V$ to braid group duality of $V$
 Theorem 9.13 and Theorem 9.19 show uniqueness of the associativity morphisms (and also braiding morphisms in the latter case) under some assumptions.
@@
 For the associativity morphisms, methods are similar to those used to verify the assumption of Theorem 9.13.
-We apply Wenzl isomorphism and Drinfeld twist construction and then we use the structural properties of the Huang-Leposwky associativity morphisms in Theorem 31.9.
-For the special braiding morphisms we take into account the formulas given by Toledano-Laredo in Theorem 31.12.
+We apply Wenzl isomorphism and Drinfeld twist construction and then we use the structural properties of the Huang-Leposwky associativity morphisms in Theorem 34.7 in [13].
+For the special braiding morphisms we take into account the formulas given by Toledano-Laredo in Theorem 34.11 in [13].
 These two comparisons first motivated our method of proof for both braided symmetry and associativity morphisms.
-The proof of (c) of Theorem 2.1 is reduced to the proof of (b) and application of Theorem 8.19.
-More in detail, for part (c) we use the fact that the special variables of the associativity morphisms and braided symmetry generate all the triples multiplicatively, in the sense that all the components except one in the variables of the associativity and braiding morphisms are taken by the generating representation in each Lie type.
+The proof of (d) of Theorem 2.1 is reduced to the proof of (c) and application of Theorem 9.19.
+More in detail, for part (d) we use the fact that the special variables of the associativity morphisms and braided symmetry generate all the triples multiplicatively, in the sense that all the components except one in the variables of the associativity and braiding morphisms are taken by the generating representation in each Lie type.
 In the proof of Theorem 9.19, and the antecedent Theorems 9.13, we explain the generating property of these special values of the associativity and braiding morphisms with the pentagon equation of associativity morphisms and the hexagonal equations of a braided symmetry in a semisimple tensor category with a generating object but, as anticipated, another important property concerning the braid group representations in the category is needed, to determine the associativity morphisms uniquely.
-11. The role of braid group duality for the generating object in determining the full braided tensor structure for the proof of (c).
-The uniqueness Theorems 8.13 and 8.19 both apply under a central duality property of the fundamental representation $V$ with respect to the braid group representation in the quantum group tensor category, which is at the core of quantum group representation theory.
+11. The role of braid group duality for the generating object in determining the full braided tensor structure for the proof of (d).
+The uniqueness Theorems 9.13 and 9.19 both apply under a central duality property of the fundamental representation $V$ with respect to the braid group representation in the quantum group tensor category, which is at the core of quantum group representation theory.
 The first instance of this property is Schur-Weyl duality for the general linear group.
@@
 We need the property that the centralizer algebras $(V^{n},V^{n})$ with respect to the action of the Drinfeld-Jimbo quantum group $U_{q}({\mathfrak{g}})$ are generated by the natural representation of the braid group induced by the $R$-matrix.
-In our case $q$ is a root of unity as usually assumed corresponding to the unitary cases of the associated fusion category, specified in [91], [85], see also Definition 20.2 in [11].
+In our case $q$ is a root of unity as usually assumed corresponding to the unitary cases of the associated fusion category, specified in [99], [92], see also Def.
+20.1, 21.2 in [13].
 The tensor powers $V^{n}$ appearing in our setting are canonical choices of truncations of the full tensor powers of the quantum group $U_{q}({\mathfrak{g}})$ obtained by induction over $n$.
-The construction of the iterated truncations $V^{n}$ of $U_{q}({\mathfrak{g}})$ at the root of unity is explained in [91] and is the starting point to construct our weak Hopf algebras in Sections 29 and 30.
-The duality property between the truncated tensor powers $V^{n}$ of the fundamental representation and the representation of the braid group is needed as a further property to complete the proof of FKL theorem on coincidence between associativity morphisms and braiding morphisms on all pairs and triples and is used in our cohomological theorems 8.13 and 8.19.
+The construction of the iterated truncations $V^{n}$ of $U_{q}({\mathfrak{g}})$ at the root of unity is explained in [99] and is the starting point to construct our weak Hopf algebras in Sect.
+32.
+The duality property between the truncated tensor powers $V^{n}$ of the fundamental representation and the representation of the braid group is needed as a further property to complete the proof of FKL theorem on coincidence between associativity morphisms and braiding morphisms on all pairs and triples and is used in our cohomological theorems 9.13 and 9.19.
 These theorems complete with cohomogical arguments what was proved up to that point in our work, coincidence of these structure maps on pairs and triples where one or two coordinated are fixed to the fundamental representation $V$.
@@
 This is explained among other things in Prop.
-2.4 in [91], see also Prop 29.2 for the same result with a slightly more detailed proof, where an important role is played by the continuous curve described in that proposition, that passes through generic values of $q$.
+2.4 in [99], see also Prop 31.2 in [13] for the same result with a slightly more detailed proof, where an important role is played by the continuous curve described in that proposition, that passes through generic values of $q$.
 See also Section 3.5 in [99] for the special tensor products $V_{\lambda}\otimes V$ of representations of $U_{q}({\mathfrak{g}})$.
@@
 Central theorems are Theorem 5.2 in [102] for the even case, and the previous Theorem 3.3 in [101] for the odd case.
-A special case has been studied by Rowell and Wenzl in [84].
+A special case has been studied by Rowell and Wenzl in [91] in their study of a conjecture by Rowell and his coauthors regarding the characterization of finiteness of the image of the braid group in the morphism algebras of tensor powers of a simple object of a braided fusion category by integrality of the squared Frobenius-Perron dimension of the object.
 Type $G_{2}$ was shown by Lehrer and Zhang [70], Morrison [75], and Martirosyan and Wenzl [72].
-In the application of Theorems 8.13 and 8.19, we work on the side of the quantum group fusion category and by the results above, our assumptions on braid group duality for the fundamental representation of our weak Hopf algebras $A_{W}({\mathfrak{g}},q)$ constructed in Section 30 are verified for the defining structure arising from Tannakian construction on the quantum group fusion category ${\mathcal{C}}({\mathfrak{g}},q))$ and Wenzl functor.
+In the application of Theorems 9.13 and 9.19, we work on the side of the quantum group fusion category and by the results above, our assumptions on braid group duality for the fundamental representation of our weak Hopf algebras $A_{W}({\mathfrak{g}},q)$ constructed in Sect.
+32 in [13] are verified for the defining structure arising from Tannakian construction on the quantum group fusion category ${\mathcal{C}}({\mathfrak{g}},q))$ and Wenzl functor.
 12. More remarks from the viewpoint of Doplicher-Roberts reconstruction in AQFT on the relevance of duality for the fundamental representation in the construction of quantum gauge group
@@
 Yamashita, we learned that modularity is used in their recent work on classification of fusion categories with fusion rules of type BCD [42].
-Thus the fact that braid group duality for the fundamental representation $V$ was the relevant missing observation helpful to complete the proof of Theorem 2.1 in the previous version of [11], which led us to Theorems 8.13 and 8.19 as a completion of Lemma 7.10 in version 7 of [11], was for us a further important recognition of the relevance of the role of quantum groups.
+Thus the fact that braid group duality for the fundamental representation $V$ was the relevant missing observation helpful to complete the proof of Theorem 2.1 in the previous version of [13], which led us to Theorems 9.13 and 9.19 as a completion of Lemma 7.10 in version 7 of [13], was for us a further important recognition of the relevance of the role of quantum groups.
 We refer the reader to the beautiful historical paper by V.
@@
 26, 27, 29 in [13] extends the notion of symmetric functors that plays a key role in Doplicher-Roberts theory [23], [24].
+In Sect.
+28 we explain the relation with the more familiar class of hermitian ribbon structures.
 13. Outlook
@@
 Other interesting future developments may regard applications in noncommutative geometry in the sense of Connes [15] of our weak Hopf algebras $A_{W}({\mathfrak{g}},q)$ associated to the quantum group fusion category ${\mathcal{C}}({\mathfrak{g}},q)$ or their twisted weak quasi-Hopf algebra structures on the Zhu algebra of an affine vertex operator algebra at a positive integer level $V_{{\mathfrak{g}}_{k}}$ to the study of Dirac operators, by expanding an analog in conformal field theory of the groundbreaking work by Neshveyev and Tuset on Dirac operators for compact quantum groups in the setting of Kazhdan-Lusztig theorem for real values of the quantization parameter $q$ [79].
- [^1] Claims, arguments, and reasoning of the recent papers [6], [7], [8] are essentially dependent on the results and in-depth discussions between S.C.
-and C.P.
-that took place from 2018 to 2022 on the contents of an entire project including [11] and the topics of the above papers, during a period in which S.C.
-was a co-author, before breaking the collaboration in early 2023 by his decision, with a minor contribution in a draft of a single section.
-In addition, C.P.
-department digital devices were subject to a serious hacking attack for several months in 2021 during the previous collaboration, which was recognized by authorities, that severely damaged her intellectual property and health, and the hacked material has not been restituted.
-Acknowledgements The author is grateful to Yi-Zhi Huang for several comments on the proof of modularity for fusion categories of affine vertex operator algebras, and the history of Finkelberg-Kazhdan-Lusztig theorem.
+ [^1] The author department digital devices were subject to a severe hacking attack for several months in 2021 which was recognized by authorities.
+The hacking severely damaged her intellectual property and health, and the hacked material has not been restituted.
+Related results have appeared in the papers [8], [9], [10].
+These papers have been sued by the author to the Italian Procura della Repubblica between fall 2021 and spring 2023.
+Their claims, arguments, and reasoning essentially depend on [13], in-depth discussions between the author and S.
+Carpi that took place from 2018 to 2022 on the contents of an entire project including the topics of the above papers before his unilateral break of the collaboration in early 2023 and formal give up of his contribution, that consisted in a draft of a single section in 2018 based on [12], and yet unpublished joint work previously done by Sergio Ciamprone and the author (the first 60-page-long version of [13]), the original Finkelberg theorem, and work by Hans Wenzl in [99].
+The author shared proofs of these attacks to several colleagues from 2022 to 2024.
+Acknowledgements The author is grateful to Yi-Zhi Huang for several comments on the proof of modularity for fusion categories of vertex operator algebras, and the history of Finkelberg-Kazhdan-Lusztig theorem.
 She is also grateful to Marco V.

### v5

@@
 Abstract.
-We explain the structure of proof of our results on a problem that Sergio Doplicher posed in the 90s in the setting of conformal nets, and on a related problem posed by Yi-Zhi Huang in the setting of vertex operator algebras.
-The latter asks to find a direct construction of an equivalence due to Finkelberg between the quantum group modular fusion category at certain roots of unity and the fusion categories of affine vertex operator algebras at positive integer levels.
-The former asks for the construction of a quantum gauge group and field algebra for the category of localized endomorphisms of a conformal net, extending Doplicher-Roberts theory of the compact gauge group and field algebra in high dimensional algebraic QFT.
-In this summary we mostly focus on the affine VOA.
-We solve Huang’s problem for all the Lie types for which it is known that the centralizer algebras of the non-negligible tensor powers of a given generating object $V$ for the quantum group fusion category (a list of $V$ for which fusion tensor powers are well defined has been given by Hans Wenzl for all the Lie types), is generated by the representation of the braid group associated to the $R$-matrix, and for Huang-Lepowsky ribbon braided tensor structure for the vertex operator algebras.
-In the classical type $A$, this property becomes the classical Schur-Weyl duality for the vector representation, and was used by Doplicher and Roberts to construct ${\rm SU}(d)$ as a special case of their unique compact gauge group and field algebra in algebraic QFT.
-Presently, known results in the literature on this generating property allow to construct our equivalence for the Lie types $A$, $B$, $C$, $D$, $G_{2}$.
-Our approach may be regarded in the setting of Connes noncommutative geometry.
-We also solve a problem by Frenkel and Zhu on the construction of a (unitary) weak quasi-Hopf algebra structure on the Zhu algebra, associated to the minimum energy functor for all the Lie types, equivalent to that from quantum groups.
-Our main tool is the construction of a unitary coboundary weak Hopf $C^{*}$-algebra in a sense extending Drinfeld coboundary quasi-Hopf algebras on a formal variable, associated to Wenzl functor for the quantum group fusion category, and of a Drinfeld twist.
-We apply a de-quantization continuous curve of Wenzl and a Drinfeld twist method to reach the Zhu algebra.
-The two Drinfeld coboundary symmetries we find in both contexts extend Doplicher-Roberts symmetric functors for rigid symmetric tensor categories with simple unit which play a central role in their duality theory.
+This paper establishes connections between a problem posed by Doplicher and Roberts in the 1990s regarding construction of quantum gauge groups associated to the category of localized endomorphisms of conformal nets, and a problem by Huang concerning a direct proof of the Finkelberg equivalence theorem.
+We also unify the treatment of high dimensional algebraic QFT with the low dimensional case, resuming the study of Mack-Schomerus quantum symmetries, for which we previously proved a simplifying Hopf algebra property and studied the unitary structure of our global quantum gauge groups $A_{W}({\mathfrak{g}},q)$ of the WZW model.
+This program leads us to a solution of Huang’s problem in the setting of vertex operator algebras.
+Building upon our previous results, our approach utilizes a quantum gauge group framework to implement a unitary weak quasi-Hopf algebra structure on the Zhu algebra via a Drinfeld twist and Wenzl’s de-quantization continuous curve derived from the unitary coboundary weak Hopf algebra $A_{W}({\mathfrak{g}},q)$ previously associated to the quantum group fusion category.
+Our construction is of an analytic nature, and was originally inspired by Drinfeld-Kohno theorem for Drinfeld category.
+Using this framework, in our previous work we established an independent rigid braided tensor structure on the module category of the affine vertex operator algebra $V_{{\mathfrak{g}}_{k}}$ at positive integer level, providing a braided tensor equivalence with the modular fusion category of the corresponding quantum group at root of unity for all Lie types.
+In this paper, for the Lie types $A$, $B$, $C$, $D$, and $G_{2}$ we completely identify our structure with the Huang-Lepowsky braided tensor structure.
+The last step of the complete identification relies on the generating property of the braid group representation in the centralizer algebras of the tensor powers of the fundamental representation of the quantum group, which is currently known for these stated Lie types.
+This reduction builds on uniqueness results of braided tensor structures in presence of a generating object.
 Dedicated to Sergio Doplicher, with gratitude
 1. Introduction
-Moore and Seiberg [73], [74] discovered a connection between their equations and a certain structure of a tensor category arising from rational conformal field theories.
-The notion of ribbon braided tensor category has been settled in mathematics, and we refer to [23], [76], [30], [29], and, for their relations to compact quantum groups, to [80], and more generally to weak (quasi-) Hopf algebras to [13].
+1.1. Historical Background: Conformal Field Theories and Quantum Groups
+Primary fields, introduced under a different name by Mack and Salam in the context of higher-dimensional conformal field theories [68], were investigated in 2D by Belavin, Polyakov, and Zamolodchikov.
+They recognized that the local symmetry algebra is given by the infinite-dimensional Virasoro algebra [5], building upon previous work by Feigin and Fuchs [29] and Kac [62] on its representation theory.
+Witten introduced an interacting $2D$ conformal field theory, now known as the Wess-Zumino-Witten (WZW) model [103].
+He showed that this model possesses a symmetry satisfying the relations of an affine Kac-Moody algebra.
+In the same year, Knizhnik and Zamolodchikov showed that the correlation functions of a WZW model satisfy a system of partial differential equations [65].
+This Kac-Moody framework was related to the Virasoro algebras by Goddard, Kent and Olive [37], who introduced the coset construction to realize Virasoro minimal models directly from WZW models.
+The mathematical formulation of the KZ equations was made rigorous by Tsuchiya and Kanie [95], who studied the model for ${\hat{\mathfrak{s}l}}_{2}$ and showed that the associated monodromy yields representations of the braid group that factor through the Hecke algebra at a root of unity.
+The discovery of quantum groups first arose in physics literature in the work by Faddeev, introduced as noncommutative function algebras.
+In the mid 80s breakthrough papers by Drinfeld and Jimbo discovered them as quantization of universal enveloping algebras, and developed their properties.
+We refer to [8] for an in-depth presentation and historical information.
+Kohno proved that the monodromy representations of the KZ equations are equivalent to those derived from the universal $R$-matrix of a quantum group [66].
+Before recalling the subsequent deep developments of the theorem by Kohno, we need first review the emergence of tensor categories in constructive conformal field theory.
+Moore and Seiberg [73], [74] derived certain equations from rational conformal field theory and discovered a connection between their equations and the structure of a tensor category.
+The corresponding mathematical framework, specifically the notion of ribbon braided tensor category, has since been settled in mathematics, we refer to [26], [27], and references therein.
+An independent derivation of strict braided tensor $C^{*}$-categories arose from the axiomatic operator algebraic approach to QFT in the 1970s, see [20], [76], [45] and references therein.
+The connections between tensor $C^{*}$-categories and compact quantum groups are studied [80].
 Turaev gave a definition of modular tensor category, based on previous work with Reshetikhin on the construction of knot and 3-manifold invariants see [96].
@@
 We denote by ${\mathcal{C}}({\mathfrak{g}},q)$ the resulting modular fusion category, with $q=e^{i\pi/d\ell}$, and $d\in\{1,2,3\}$ a suitable integer structure constant from Lie theory.
-Unitary tensor structures have been constructed by Wenzl and Xu [99], [105], and the former paper, as well as more recent papers by Wenzl and his coauthors play a major role in our work, and we try to explain this in this paper.
-On the side of CFT, our main interest in this paper is the Wess-Zumino-Witten model.
+Unitary tensor structures have been constructed by Wenzl and Xu [99], [105].
+Wenzl’s work plays a major role in our work, and we try to explain this in this paper.
+In the constructive approach to CFT, the Wess-Zumino-Witten model is of central interest.
 Modular, or more generally ribbon rigid braided tensor category constructions involved many authors among physicists and mathematicians.
@@
 Another main difficulty is to show rigidity of the category.
+Non-strict examples motivated by conformal field theory arose from the work of Drinfeld on quasi-Hopf algebras [24].
 Let $\mathfrak{g}$} be a complex simple Lie algebra.
-These problems were first solved by Drinfeld in [25], [27] [28] for the Drinfeld category with the introduction of the notion and construction of non-trivial examples of quasi-Hopf algebras over $U({\mathfrak{g}})[[h]]$, via the Knizhnik-Zamolodchikov differential equation which is satisfied by the correlation functions of the WZW model.
+The mentioned problems were first solved by Drinfeld in [22], [24] [25] for the Drinfeld category with the introduction of the notion and construction of non-trivial examples of quasi-Hopf algebras over $U({\mathfrak{g}})[[h]]$, via the KZ differential equation.
 Drinfeld used this equation to construct the associativity morphisms compatible with naturally assigned braiding morphisms.
-Furthermore he established rigidity of this category by showing a remarkable equivalence with the module category arising from the quantized universal enveloping algebra $U_{h}({\mathfrak{g}})$, which is a Hopf algebra.
-To construct the equivalence, he used Drinfeld coboundary matrix ${\overline{R}}$, which is a modification of the $R$-matrix by a square root of the coboundary defined by the ribbon structure.
-This is his remarkable Drinfeld-Kohno theorem, expanding an earlier theorem by Kohno [69] stating that the representation of the braid group arising from monodromy of the KZ equation is equivalent to that arising from the action of the $R$-matrix of $U_{h}({\mathfrak{g}})$ with respect to the vector representation, for the classical Lie types.
-Drinfeld-Kohno theorem connects via a Drinfeld twist, the quasi-Hopf algebra $U({\mathfrak{g}})[[h]]$ with the quantized universal enveloping algebra $U_{h}({\mathfrak{g}})$ which is a Hopf algebra.
-As above mentioned, rigidity of Drinfeld category is derived from the equivalence constructed in Drinfeld-Kohno theorem and rigidity of the representation category of $U_{h}({\mathfrak{g}})$.
+Furthermore he established rigidity of this category by showing a remarkable ribbon, braided tensor equivalence with the module category arising from the quantized universal enveloping algebra $U_{h}({\mathfrak{g}})$, which is a Hopf algebra.
+To construct the ribbon, braided, tensor equivalence, he used Drinfeld coboundary matrix $\overline{R}$}, which is a modification of the $R$-matrix by a square root of the coboundary defined by the ribbon structure.
+This is his remarkable Drinfeld-Kohno theorem, expanding the earlier theorem by Kohno [66].
+Drinfeld-Kohno theorem connects via a Drinfeld twist, the quasi-Hopf algebra $U({\mathfrak{g}})[[h]]$ with the quantized universal enveloping algebra $U_{h}({\mathfrak{g}})$.
+In particular, the representations of the braid group arising from monodromy of the KZ equation are shown equivalent to those arising from the action of the $R$-matrix of $U_{h}({\mathfrak{g}})$ with respect to the vector representation, for the classical Lie types.
 Major work was carried out in a series of papers by Kazhdan and Lusztig in the mid 90s in the setting of affine Lie algebras.
-The authors construct rigid braided tensor structures for certain module categories of affine Lie algebras associated to ${\mathfrak{g}}$ at levels excluding the positive rationals.
-In their work they consider both generic and non generic values of the levels.
-In particular they consider the negative integer levels $-k$, $k\in{\mathbb{N}}$, and proved a braided tensor equivalence of their category ${\mathcal{O}}_{-k}$ with the non-semisimple category of certain modules of the Drinfeld-Jimbo quantum group $U_{q}({\mathfrak{g}})$ at the same type of roots of unity $q=e^{i\pi/d\ell}$, with the exception of some exceptional cases at low levels [66].
-The order $\ell$ appearing in the root of unity and the positive integer $k$ are related by a well known equality, see e.g.
-Def.
-20.1 and 21.2 in [13] and references therein.
-Shortly after, Finkelberg considered a certain semisimple subquotient $\tilde{\mathcal{O}}_{-k}$ of the representation category constructed by Kazhdan and Lusztig, paralleling the subquotient construction in the setting of quantum groups, and described a passage $\tilde{\mathcal{O}}_{k}\to\tilde{\mathcal{O}}_{-k}$ claimed as a braided tensor equivalence from the category $\tilde{\mathcal{O}}_{k}$ of certain integral modules of the affine Lie algebra at the positive integer level $k$ to the subquotient rigid braided tensor category $\tilde{\mathcal{O}}_{-k}$.
-Combining with the equivalence by Kazhdan and Lusztig, Finkelberg derived a connection between $\tilde{\mathcal{O}}_{k}\to{\mathcal{C}}({\mathfrak{g}},q)$.
+The authors construct rigid braided tensor structures for certain module categories of affine Lie algebras associated to $\mathfrak{g}$} at shifted levels excluding the non negative rationals.
+For convenience we denote by -$\ell$ the shifted levels allowed by Kazhdan and Lusztig.
+In their work they proved a braided tensor equivalence of their category $\mathcal{O}$$_{-\ell}$ with the non-semisimple category of certain modules of the Drinfeld-Jimbo quantum group $U_{q}({\mathfrak{g}})$ at the root of unity $q=e^{i\pi/d\ell}$, with $d$ the ratio of the square lenghts of the long to the short root, with the exception of some cases [63].
+With $k$ a positive integer, defined as the physical level, and $\ell=k+\check{h}$, with $\check{h}$ the dual Coxeter number of $\mathfrak{g}$}, Finkelberg considered a certain semisimple subquotient $\tilde{\mathcal{O}}_{-\ell}$ of the representation category constructed by Kazhdan and Lusztig, paralleling the subquotient construction of the quantum group fusion category ${\mathcal{C}}({\mathfrak{g}},q)$ in the setting of quantum groups at roots of unity.
+He defined an equivalence $\tilde{\mathcal{O}}_{\ell}\to\tilde{\mathcal{O}}_{-\ell}$ from the category $\tilde{\mathcal{O}}_{\ell}$ of certain integral modules of the affine Lie algebra at level $k$ introduced by Beilinson, Feigin, and Mazur to the subquotient rigid braided tensor category $\tilde{\mathcal{O}}_{-\ell}$.
+Combining with the equivalence by Kazhdan and Lusztig, Finkelberg derived an equivalence between $\tilde{\mathcal{O}}_{\ell}\to{\mathcal{C}}({\mathfrak{g}},q)$.
 See the introduction and Theorem 4.3 in [30].
-The original work by Finkelberg refers to the cited work in physics by Moore and Seiberg and in algebraic geometry by Beilinson, Feigin, and Mazur [4] for the rigid braided tensor structure on $\tilde{\mathcal{O}}_{k}$.
-As I learned from discussions with Huang, rigidity was not proved in [4], and there was no proof of rigidity using this approach.
-Remarkably, Huang and Lepowsky introduced and developed a rigorous direct construction of a modular (vertex) tensor category structure in a very general setting for vertex operator algebras, including as a particular case the affine vertex operator algebras at positive integer levels $k$ that we denote by $V_{{\mathfrak{g}}_{k}}$.
-The construction has been completed by Huang between 2005 and 2008 [55], [59] following a general construction of tensor product theory in the setting of vertex operator algebras by Huang and Lepowsky [50], [51], [52] and by Huang [53], [58].
-In particular [53] gives the construction of associativity morphisms which is the first difficulty in the construction of the tensor structure.
-An historical account may be found in [60].
-We also refer to the introductions of [54] and [61].
-These works give in particular the first mathematical proof of modularity, that applies moreover to a large class of representation categories of vertex operator algebras.
-In particular rigidity and modularity property in the setting of vertex operator algebras was first shown in [55].
-To show modularity, and in particular rigidity, Huang first discovered the need of the Verlinde formula.
-The paper [58] gives a proof of the Verlinde formula in the general case.
-See also [56] for the proof of convergence and extension property assumed in [53] and [57] for the proof of modular invariance conjecture of Moore and Seiberg needed in the proof of the Verlinde formula.
-By a private communication with Huang on the history of this problem, I learned that it finally turned out that the construction by Finkelberg described in [33] connecting $\tilde{\mathcal{O}}_{k}$ with Kazhdan and Lusztig structure inherited by $\tilde{\mathcal{O}}_{-k}$ could not suffice to derive rigidity of $\tilde{\mathcal{O}}_{k}$ from that of $\tilde{\mathcal{O}}_{-k}$.
-This implied that the proof of the equivalence $\tilde{\mathcal{O}}_{k}\to\tilde{\mathcal{O}}_{-k}$ was far from complete.
-I have been informed by Huang that these doubts were motivated by two facts.
-On one hand, his earlier direct communications in the 90s had clarified that rigidity was not proved in [4].
-On the other, the mentioned property that his proof of rigidity in the setting of vertex operator algebras involved the Verlinde formula indicated that more work was likely needed to complete the proof in [33].
-These facts together with early claims of rigidity in [33], motivated Huang to a comparison of the results in [33] with their results on rigidity and this led to note a gap in a proposition of [33] with implications on the main result.
-Huang’s study had led him to identify the gap in the fact that a suitable constant needed to be shown to be nonzero in a proposition useful to obtain the equivalence between the two local systems considered by Finkelberg and Kazhdan and Lusztig, and this proposition would imply derivation of rigidity in his case from previous work.
-On the other hand, Huang proof of rigidity in the setting of vertex operator algebras had similarly reduced to show nontriviality of a constant, and to prove this he had used the Verlinde formula as above alluded to, which was a main result in their work.
-This had led Huang to inform Finkelberg about utility of this formula to show rigidity in his case.
-As later reported by Finkelberg in [34], it turned out that the completion of the proof of that proposition leading to rigidity of $\tilde{\mathcal{O}}_{k}$ and the equivalence $\tilde{\mathcal{O}}_{k}\to\tilde{\mathcal{O}}_{-k}$, involved the Verlinde formula as well (which became another main result in Finkelberg’s paper).
-This reflects also on Finkelberg composed equivalence $\tilde{\mathcal{O}}_{k}\to\tilde{\mathcal{O}}_{-k}\to{\mathcal{C}}({\mathfrak{g}},q)$, with Kazhdan-Lusztig equivalence that moreover still passes through the negative integer levels.
-This implies in particular that the theorem does not include a few exceptional cases, including ${\mathfrak{g}}=E_{8}$ and $k=2$, which correspond to cases not treated in Kazhdan-Lusztig theory.
-Following the introductions in [54], [61], Finkelberg work [33], [34] may be reinterpreted as giving the construction of a braided tensor category structure on $\tilde{\mathcal{O}}_{k}$.
-One may also conclude that these papers give a second proof of rigidity and modularity of this category.
-An approach to modularity based on algebraic geometry may be found in Theorem 7.0.1 [1] but the proof of rigidity is not given and, as already mentioned it was known to experts that there was no proof based on their approach at the time.
-The author thanks Y.-Z.
-Huang for this information.
-We point out the very recent paper by Etingof and Penneys where the authors show among other things that weak rigidity in the sense of [1] is equivalent to rigidity for certain semisimple braided tensor categories of moderate growth with simple unit, see [31] giving a third proof of rigidity and give applications to a simplification of Huang’s proof of rigidity.
-They also answer positively to a question posed in [1] that for every finite semisimple category over an algebraically closed field the data of a modular functor is essentially equivalent to the structure of a modular fusion category, completing the proof of rigidity in [1].
-On the other hand, Finkelberg equivalence theorem is stated as Theorem 7.0.2 in [1] without proof.
-The braided tensor equivalence between $\tilde{\mathcal{O}}_{k}\to{\mathcal{C}}({\mathfrak{g}},q)$ obtained from the combined works by Finkelberg, Kazhdan and Lusztig is considered indirect by several authors.
-This leads to the desire of constructing the equivalence in a natural, direct, and explicit way, and this problem is discussed by Y.-Z.
-Huang in [61] in great detail, where we select only a small part.
+While the equivalence between quantum group fusion categories and categories of modules of affine Lie algebras at positive integer level was proposed in the mid 1990s, the complete proof of rigidity of $\tilde{\mathcal{O}}_{\ell}$ and derivation of the tensor equivalence was given in 2013 using the Verlinde formula [31] building on work by Faltings and Teleman.
+We refer the reader to the appendix for more historical remarks concerning the proof of rigidity for affine vertex operator algebras and affine Lie algebras at positive integer levels.
+The ribbon braided tensor equivalence established in [30], [31] relies on a passage through negative levels and then applies the equivalence with the quantum group fusion category by Kazhdan and Lusztig.
+As a consequence, Finkelberg’s proof excludes certain exceptional cases like $\mathfrak{g}=E_{8}$ at level $k=2$ since this case is excluded in [63].
+1.2. Huang’s Problem
+This led us to consider the following problem posed by Y.-Z. Huang in [58].
 Problem 4.4 in [58].
 Find a direct construction of this equivalence without using the equivalence given by Kazhdan-Lusztig so that this equivalence covers all the cases, including the important ${\mathfrak{g}}=E_{8}$ and $k=2$ case.
-We also mention Gannon’s observation in Sect. 6.2.3 of [38] regarding the desire of finding a direct connection between quantum groups and CFT.
-From the side of mathematical axiomatizations of CFT, we work with respect to Huang and Lepowsky braided tensor structure for the representation category of vertex operator algebras.
-A direct construction of the equivalence may be helpful to see how rigidity or modularity in the setting of CFT may be transported from corresponding results in the setting of quantum groups at roots of unity, which seem simpler, or vice versa, by a given explicit equivalence.
-Or to simplify or give new construction of other structures, e.g.
-unitary structures, with a quantum group viewpoint, or to see more connections with quantum groups of models derived from or generalizing these fundamental models.
-The study of how unitarization from the quantum group fusion categories [99] was useful to derive unitarization in representation categories of vertex operator algebras via an explicit equivalence that passes through weak quasi-Hopf algebras was important for us since the early stages of applications of the theory of weak quasi-Hopf algebras [13], and further motivated us to study the above problem.
-Indeed one of our first results on unitarization was that weak quasi-Hopf algebras associated to fusion categories known to be equivalent allow to transport the unitary structure of one category to unitarize the other.
-In the process, knowledge of the explicit equivalence gives an understanding of the transported unitary structure.
-We have results on the structure of the unitary structure of the affine vertex operator algebra at positive integer levels transported by our equivalence, which asserts that it must correspond to the classical compact real form, and necessarily equals the unitary structure previously directly constructed by Gui, see [13] for more details.
-Our starting point to approach this problem arose from a related problem generously posed by Sergio Doplicher in the 90s to the author in a different setting, that of algebraic quantum field theory.
-We try to explain this problem.
-Our approach to Huang problem is based on interpreting Finkelberg-Kazhdan-Lusztig equivalence as an extension to quantum groups of Doplicher-Roberts duality theory for the important class of WZW models that are described by affine Lie algebras at positive integer levels.
-We shall briefly come back to the previous history of these important models in physics.
-Algebraic quantum field theory had been initiated by Haag and Kastler in [47] in the 60s, and developed by founding breakthrough works by Doplicher, Haag, and Roberts [17], [18], [19], [20] in the 70s.
-We refer to the book by Haag [48], for an in-depth introduction to this theory, based on functional analysis, and more specifically on operator algebras, and also for more references.
-In this context, I learned from private discussions that for high dimensional quantum field theories categorical structures in their study first emerged from the discovery by Doplicher of the notion of localized endomorphism of the local observable algebra, allowing composition of two of them which was understood as a strict tensor product structure.
-The mentioned works by Doplicher, Haag, and Roberts led to a deep construction of the structure of a rigid symmetric strict tensor $C^{*}$-category with simple unit, starting from first principles in physics.
+Moreover, in light of Drinfeld’s construction of the ribbon braided tensor equivalence in Drinfeld-Kohno theorem, and his consequent derivation of rigidity of the Drinfeld category, is natural to ask whether the Verlinde formula is necessary to prove rigidity, and whether rigidity is necessary to prove the ribbon braided tensor equivalence $\tilde{\mathcal{O}}_{\ell}\to{\mathcal{C}}({\mathfrak{g}},q)$.
+We also mention Gannon’s observation in Sect.
+6.2.3 of [35] regarding the desire of finding a direct connection between quantum groups and CFT.
+Our work provides a direct constructive method–addressing Huang’s Problem 4.4 in [58]–by utilizing a global quantum gauge group method originating in algebraic quantum field theory, and the Zhu algebra to establish the equivalence.
+From the side of mathematical axiomatizations of CFT, Huang and Lepowsky have done extensive work on the construction of tensor structures of modules and showing all the properties of modular tensor category [47], [48], [49], [50], [51], [52], [53], [54], [55], [56].
+An explicit construction of the equivalence is helpful to see how rigidity, modularity or unitarity in the setting of CFT may be transported from corresponding results in the setting of quantum groups at roots of unity, which seem simpler, by our explicit equivalence.
+1.3. Inspiration from Algebraic Quantum Field Theory
+Our starting point to approach this problem arose from a related problem posed by Doplicher and Roberts in the axiomatic setting of the operator algebraic approach to quantum field theory (AQFT).
+Problem posed in Section 7 in [21].
+Can we extend the duality theory for compact groups [61] Theorem 6.1 to more general objects, by replacing strict symmetric monoidal C*-categories by strict braided monoidal C*-categories?
+What is the full class of compact group-like objects which arise in this way?
+We next explain this problem and connect it to Huang’s problem.
+Algebraic quantum field theory was initiated by Haag and Kastler in [44] in the 1960s, and developed by founding breakthrough works by Doplicher, Haag, and Roberts [14], [15], [16], [17] in the 1970s.
+The operator algebraic approach to QFT joins high rigour provided by the operator algebra methods with first principles in physics on which the theory is formulated.
+We refer to the book by Haag [45], for an in-depth introduction to this theory.
+In the context of high dimensional quantum field theories, categorical structures first emerged from the discovery of the notion of localized endomorphism of the local observable algebra, allowing composition of two of them which was understood as a strict tensor product structure.
+The mentioned works by Doplicher, Haag, and Roberts led to a deep construction of the structure of a rigid symmetric strict tensor $C^{*}$-category with simple unit.
 Rigidity, called charge conjugation in this setting, describes particle-antiparticle duality.
-In their papers rigidity is derived as a weak limit of inner automorphisms induced by transfer operators moving the support of the given localized endomorphism $\rho$ to (space-like) infinity.
+In their work rigidity is derived as a weak limit of inner automorphisms induced by transfer operators moving the support of the given localized endomorphism $\rho$ to (space-like) infinity.
 This gives a left inverse $\Phi_{\rho}$ of $\rho$, which is a completely positive map.
@@
 It was shown that the values of the statistics parameter (which is the constant $\Phi_{\rho}(\varepsilon)$, with $\varepsilon$ the action of the generator of the permutation group in the algebra of intertwining operators of the tensor powers of $\rho$), can be completely classified, and it can only take the values given by the reciprocal of a positive integer $d$ (the order of the parastatistics) up to a sign which classifies charges of para-Bose or para-Fermi type.
-This analysis also needed to prove that the statistics parameter can not take the value 00 and was excluded by physically motivated conditions [20], [48].
+This analysis also needed to prove that the statistics parameter can not take the value 0 and was excluded by physically motivated conditions [17], [45].
 Because of the representation of the permutation group in the category, together with their results on the classification of the statistics parameter, and the $C^{*}$-structure, in the 70s it arose the question of whether that was the category of representations of a compact group, because those data arising from a single localised endomorphism were also satisfied by the representation category of a compact Lie subgroup of ${\rm SU}(d)$, that would be a gauge group of global symmetries.
@@
 [26].
-In a series of important papers dated the end of the 80s, Doplicher and Roberts were able to solve the problem in full generality.
+In a series of papers dated the end of the 80s, Doplicher and Roberts were able to solve the problem in full generality.
 Their construction gives a unique compact gauge group with representation category equivalent to that of a given symmetric rigid strict tensor $C^{*}$-category with simple unit, yielding the formulation of a new abstract duality theory for compact groups [18], [20],[19], and led to desired applications to high dimensional Algebraic QFT [21].
@@
 This property allows to find their unique compact gauge group $G$.
-In two or three dimensional quantum field theory, it was noted that the corresponding categories were braided rather than symmetric.
-This implies that Bose-Fermi alternative does not hold any longer.
-Models were first described by Streater and Wilde.
-Doplicher, Haag, Roberts theory of superselection sectors was generalized to braid groups statistics by Fredenhagen, Rheren, Schroer [35], [36].
-The representation of the braid group in the tensor $C^{*}$-category is unitary.
-The discovery of quantum groups first arose in physics literature in the work by Faddeev, introduced as noncommutative function algebras.
-In the mid 80s breakthrough papers by Drinfeld and Jimbo discovered them as quantization of universal enveloping algebras, and developed their properties.
-We refer to the introduction in [11] and more historical information at the end of each chapter in their book.
-Woronowicz developed a theory of compact quantum groups as noncommutative function algebras in the setting of $C^{*}$-algebras and constructed the first important examples as deformation of ${\rm SU}(d)$ by a real parameter, see [80] and reference to the original papers, via a concrete Tannaka-Krein duality for $C^{*}$-algebras extending the classical case.
-After the discovery of the quantum groups in the mid eighties, Doplicher asked the author to try to extend their theory and consequent applications on the construction of a quantum gauge group and a related field algebra.
-Of course, several studies were dedicated to extend Doplicher-Roberts duality theory to quantum groups, see Ch.
-IV.5 in [48] for more references on an early stage of this problem, see also the introduction in [13].
-Our approach may be regarded in the setting of Connes noncommutative geometry [15], and follows naturally our previous work in the setting of compact quantum groups, see also the introduction in [14], [16].
-In [83], [84], [88], [85], [86], [87], [89] we have developed at the level of rigid tensor $C^{*}$-categories with simple unit an analogue of Doplicher-Roberts construction of a compact gauge group and field algebra, or more precisely the analogue of [23] in the setting of compact quantum groups.
-The analogy cannot be performed in a strict way due to new difficulties in the quantum case, see the mentioned papers and [13].
-One important difference was that for rigid braided tensor categories we do not have braided tensor functors to the categories of vector spaces, beyond the symmetric ones, because the category of vector spaces does not admit a braided symmetry which is not a permutation symmetry.
-Giving up a compatibility property in the symmetric case implies lack of uniqueness of the functor, and thus of the gauge group.
-This was noticed early as an easy consequence from their construction.
-To describe our approach, we recall that Drinfeld original motivation for introducing quasi-Hopf algebras was to interpret them as symmetries to describe braided tensor categories from important models of rational conformal field theories motivated by the work by Moore and Seiberg.
+1.4. Quantum symmetries and primary field reconstruction in low-dimensional AQFT
+In low-dimensional quantum field theories (specifically in two and three spacetime dimensions), the categories of superselection sectors naturally admit a braided tensor structure rather than a symmetric one.
+Consequently, the standard DHR Bose-Fermi alternative does not hold any longer, allowing for the existence of non-integer statistics.
+Early models exhibiting such behavior were first constructed by Streater and Wilde.
+The DHR theory of superselection sectors was subsequently generalized to accommodate braid group statistics by Fredenhagen, Rehren, and Schroer [32, 33], who showed that the representations of the braid group within these tensor $C^{*}$-categories are unitary.
+Motivated by the Drinfeld-Kohno theorem and the representation theory of quantum groups at roots of unity–where Hermitian representations of $U_{q}(\mathfrak{sl}_{2})$ become non-semisimple and exhibit null vectors–Mack and Schomerus sought to recover the truncated fusion tensor products of conformal field theory (CFT) from a semisimple quotient algebra.
+To this end, they extended Drinfeld’s framework to define weak quasi-Hopf algebras.
+They proposed that the true global quantum symmetries of minimal conformal models are weak quasi-Hopf algebras canonically associated with $U_{q}(\mathfrak{sl}_{2})$ at roots of unity.
+As a first main example, they rigorously showed that the primary field operators of the chiral Ising model satisfy weak quasi-Hopf-covariant operator product expansions and obey local braid relations [69].
+Their program culminated in Schomerus’s generalized construction of a weak quasi-Hopf algebra and a corresponding field algebra for any rational conformal field theory, extending the Doplicher-Roberts reconstruction theorem to the braided case [70], [93].
+Mack-Schomerus quantum symmetry program saw further but limited development.
+A primary obstacle was that the emerging quantum gauge groups in the general theory were highly non-unique and possessed a complicated algebraic structure, largely due to the non-triviality of the associator.
+These ambiguities remained unclarified, leaving an incomplete reconciliation with the standard Doplicher-Roberts theory.
+Furthermore, the pioneering examples provided by Mack and Schomerus were derived from quantum group representation theory, rather than reconstructed intrinsically from the local observables.
+This approach raises doubts as to whether their symmetries can be understood as weak versions of Hopf algebras–a notion and a connection they did not address.
+The algebraic results of Tsuchiya and Kanie were developed into the operator algebraic approach to conformal field theory framework by A.
+Wassermann [98].
+Utilizing von Neumann algebras, Wassermann constructed the algebras of observables for chiral model of the CFT describing the WZW model, known as the loop group conformal net.
+Subsequently, the theory developed toward the direct computation and classification of the tensor categories of localized endomorphisms.
+By largely leaving the quantum gauge group viewpoint aside, this structural shift proved successful for 2D classifications.
+However, it also resulted in a lack of a unifying mathematical language between low-dimensional braided theories and higher-dimensional symmetric theories governed by standard compact gauge groups.
+1.5. Our Approach to the two Problems: Quantum Gauge Groups and Drinfeld Twists
+Our approach to Huang’s problem and to Doplicher-Roberts problems is derived from synthesizing the developments in affine Lie algebras by Wassermann and Toledano-Laredo on the loop group conformal net [94] with the unitary properties of quantum group fusion categories previously highlighted by Wenzl [99], from the view point of quantum symmetries in algebraic conformal field theory.
+In our approach primary fields play a prominent role to establish the equivalence, as originally envisaged by Mack and Schomerus.
+In our approach, a weak quasi-Hopf algebra is naturally associated to the conformal field theory, while a weak Hopf algebra, in a suitable new sense, is naturally associated to the quantum group fusion category, and we connect them by a Drinfeld twist.
+Specifically, we solve Huang’s problem and advance Mack-Schomerus quantum symmetry program, by introducing a global quantum gauge group approach.
+We construct finite-dimensional weak Hopf $C^{*}$-algebras $A_{W}({\mathfrak{g}},q)$ with unitary coboundary structures associated to the fusion category of quantum groups at roots of unity.
+To achieve this, we utilize a specific linear fiber functor $W$ originating from the work of Wenzl on ${\mathcal{C}}({\mathfrak{g}},q)$ to the category of Hilbert spaces [99].
+By applying an explicit Drinfeld twist–analogous to Drinfeld’s original proof of the Drinfeld-Kohno theorem–we implement a unitary weak quasi-tensor structure directly onto the Zhu algebra.
+This explicit construction allows us to naturally and directly transport the rigid, modular, and unitary tensor structures from the quantum group category to the representation category of the affine vertex operator algebra (VOA) at positive integer levels, thereby providing the desired direct link in Huang’s problem.
+Our work has roots in our previous works in the setting of compact quantum groups of Woronowicz, [11], [13], [83][89]; however the different nature of categories involved here necessitates a very different framework.
+Woronowicz developed compact quantum groups as noncommutative function algebras in the setting of $C^{*}$-algebras via a concrete Tannaka-Krein duality for $C^{*}$-algebras extending the classical case.
+In his case, categories are naturally endowed with a fibre functor to the category of f.d.
+Hilbert spaces.
+Our approach extends Woronowicz Tannaka-Krein duality on several aspects.
+The first aspect is that our categories do not have a fibre functor a priori.
+Similarly to Doplicher-Roberts setting, we must construct one naturally, compatible with the braided tensor structure of the category.
+We need to do this first for the quantum group fusion category ${\mathcal{C}}({\mathfrak{g}},q)$, and then for the VOA fusion category ${\rm Rep}(V_{{\mathfrak{g}}_{k}})$.
+To describe our approach, we go back to the constructive approach of rigid braided tensor categories motivated by CFT, mentioned in Subsect.
+1.1.
+We recall that Drinfeld original motivation for introducing quasi-Hopf algebras was to interpret them as symmetries to describe braided tensor categories from important models of rational conformal field theories motivated by the work by Moore and Seiberg.
 Although as noted in [23], they could not describe precisely simple models such as the Ising, Drinfeld evaluated quasi-Hopf algebras a good description of quantum symmetries for the WZW model in an approximate way, i.e.
 in the limit to infinity of the central charge.
-From this perspective, we regard the equivalence described by Drinfeld-Kohno theorem as a Tannakian counterpart at the level of quasi-Hopf algebras of Kazhdan-Lusztig-Finkelberg equivalence $\tilde{\mathcal{O}}_{k}\to{\mathcal{C}}({\mathfrak{g}},q)$.
-This point of view is the author understanding of the simplified proof given by Neshveyev and Tuset of Drinfeld-Kohno theorem following the proof by Kazhdan and Lusztig of the same theorem over the complex numbers, and for real values of the deformation parameter $q$, in the setting of compact quantum groups.
-Their approach inspired us to a similar strategy in our case of an abstract duality theory for both the fusion categories of quantum groups at roots of unity and of affine vertex operator algebras at positive integer levels [78].
-There have been studies that construct analogues of Doplicher-Roberts field algebra using generalizations of quasi-Hopf algebras (the weak quasi-Hopf algebras) as quantum gauge groups, and they showed complications reflecting non-triviality of the associator, see [93].
-It would be desirable to be able to use quantum symmetry groups that may be regarded in the family of Hopf algebras.
-To treat the difficulty of finding a replacement of Doplicher-Roberts symmetric functors, in [12] and [13] we have taken into account Drinfeld notion of coboundary symmetry $\overline{R}$, which plays a role in the construction of Drinfeld twist in Drinfeld-Kohno theorem [27].
-This is a symmetry of the category obtained modifying the braided symmetry defined by the $R$-matrix by a square root of the $2$-coboundary defined by the ribbon structure, and by construction the coboundary symmetry is an involutive symmetry, similarly to the permutation symmetry.
-This symmetry finally allows for a generalization of the notion of symmetric functor.
-Moreover we have followed both the original choice by Mack and Schomerus [71] of a given dimension function corresponding to non-negligible tensor powers of the vector representation of quantum ${\mathfrak{s}l}_{2}$ at roots of unity and also the study of classification of compact quantum groups corresponding to a given representation category of deformation compact quantum groups by Neshveyev and Yamashita [81], [82].
-They both considered a specific positive integer valued dimension function associated to the given category of the quantum group, which was the classical dimension function.
-This suggested to us to consider in [12] and [13] a specific dimension function on a functor related to classical data, and this was first Wenzl functor [99] for the quantum group fusion category now at roots of unity, and then the minimum energy functor for the WZW model [106] in the setting of vertex operator algebras.
-It was particularly important for us that in [99] Wenzl describes an explicit quantization curve associated to his functor that leads to an algebra that identifies with the Zhu algebra for that model in the setting of vertex operator algebras.
-Since in [12] for the type $A$ case we had constructed a natural quantum group structure on a semisimple associative algebra from the quantum group fusion category endowed with Wenzl functor, compatible with the tensor structure, this arose the question of constructing another natural quantum group structure on the Zhu algebra at the time of [12], following the quantization curve on the unitary structure of tensor products with the aim of making it classical.
-This involved in [13] the construction of a Drinfeld twist in our setting, the above mentioned notion that Drinfeld had introduced to study equivalence of the Drinfeld category with the quantum group category in a formal setting.
-We have done this for all the Lie types via our version of Drinfeld-Kohno Theorem 30.4 in [13] and applications to affine vertex operator algebras at positive integer level, via Zhu linear equivalence, in [13].
+Our approach to Huang problem is based on the program of understanding Finkelberg-Kazhdan-Lusztig combined equivalence as an extension to quantum groups of Doplicher-Roberts duality theory for the class of WZW models that are described by affine Lie algebras at positive integer levels.
+From this perspective, we regard the equivalence described by Drinfeld-Kohno theorem as a Tannakian counterpart that needs to be implemented on the categories appearing on the sides of Kazhdan-Lusztig-Finkelberg equivalence $\tilde{\mathcal{O}}_{\ell}\to{\mathcal{C}}({\mathfrak{g}},q)$.
+In building this abstract duality theory, we were inspired by Neshveyev and Tuset’s simplified proof of the Drinfeld-Kohno theorem [78] in the setting of compact quantum groups, viewing it as a Tannakian counterpart implemented at the level of Hopf and quasi-Hopf algebras.
+The construction of the weak Hopf algebras, weak quasi-Hopf algebras starting from the category and a linear fibre functors are not present in their work.
+Drinfeld original approach was the opposite, to start from the construction of Hopf and quasi-Hopf algebras.
+Their methods for constructing the twist are inspired by Kazhdan-Lusztig work, our methods are inspired by Drinfeld original twist.
+Differences between the two situations arise from the different unitary structures at roots of unity or at real values of $q$.
+The weak Hopf algebras $A_{W}({\mathfrak{g}},q)$ that we construct in [10] from the quantum group fusion category are generalizations of duals of the compact quantum groups of Wornowicz: They have algebraic and unitarity properties that extend properly that of the the duals of the compact quantum groups.
+The main algebraic difference is that the coproduct $\Delta$ can not be unital in our case due to the fusion rules of the quantum grpoup fusion category.
+It is also not strictly associative either, but this lack of strict associativity is rather mild, because we construct a fibre functor satisfying two tensorial diagrams.
+Moreover the antipode satisfies the usual Hopf algebra axioms and commutes with the -involution.
+The -involution of $A_{W}({\mathfrak{g}},q)$ is anticomultiplicative, rather than multiplicative as in the case of compact quantum groups.
+The hermitian structure of quantum group fusion categories at roots of unity was analysed in [99].
+This structure implies in a non-trivial way that the adjoint bialgebra and the opposite bialgebra of $A_{W}({\mathfrak{g}},q)$ coincide, reflecting a similar structure of the hermitian structure of $U_{q}({\mathfrak{g}},q)$–although $A_{W}({\mathfrak{g}},q)$ is not obtained by a co-ideal.
+To form tensor product of -representations, we need a twist relating the adjoint bialgebra of $A_{W}({\mathfrak{g}},q)$ with the given bialgebra.
+The existence of the twist allows to unify and extend the theory with that of the compact quantum groups.
+Our twist is inspired by the twist that Drinfeld used for a different purpose: the proof of the Drinfeld-Kohno theorem.
+To construct this twist, Drinfeld made use of Drinfeld coblundary $\overline{R}$-matrix.
+In the case of symmetric tensor categories, the coboundary matrix is the identity.
+Otherwise it induces an intertwiner $\rho\otimes\sigma\to\sigma\otimes\tau$ in the fusion category, similarly to how an $R$-matrix induces the braiding as an intertwiner.
+However for the coboundary, the intertwiner does not satisfy the braid group relation, but has always of period $2$, extending the property of the permutation symmetry in the classical case.
+We summarize the structure of $A_{W}({\mathfrak{g}},q)$ by saying that this algebra is a unitary coboundary weak Hopf algebra.
+1.6. Main theorem
+Theorem 1.1.
+Let $\mathfrak{g}$} be a complex simple Lie algebra and $k$ a positive integer, the level.
+Let $q=e^{\frac{i\pi}{d(k+h^{\vee})}}$, with $h^{\vee}$ the dual Coxeter number and $d$ the ratio of the square lengths of the long to the short root.
+Then the unitary coboundary weak Hopf algebra $A_{W}({\mathfrak{g}},q)$ induces a rigid braided tensor category structure on the module category of the affine vertex operator algebra $V_{{\mathfrak{g}}_{k}}$.
+Furthermore we construct, via the Drinfeld twist method, a braided tensor equivalence between this category and the modular quantum group fusion category ${\mathcal{C}}({\mathfrak{g}},q)$ for all Lie types.
+Regarding the comparison with the Huang-Lepowsky ribbon braided tensor structure:
+  • (1)
+    For the Lie types $A$, $B$, $C$, $D$, $G_{2}$ we obtain a complete identification.
+  • (2)
+    For the remaining exceptional Lie types we obtain the identification for the tensor product structure, the ribbon structure, and those braiding and associativity axioms for morphisms where one or two variables are the fundamental representation $V$, respectively.
+The previous result is stated in greater detail as Theorem 3.1.
+Our starting point for this theorem was in the context of constructive rigorous conformal field theory, in the setting of operator algebras.
+The main references important for us to start to set up the connection between quantum groups and affine vertex operator algebras, are the work of Wenzl [99], and the work by Wassermann [98] in the setting of affine Lie algebras.
+A key concept is played by the notion of primary field, and its initial term, which determines the primary field uniquely, as shown by Tsuchiya and Kanie [95].
+We use this fact in a fundamental way to connect quantum group fusion category to representations of the Zhu algebra and their special tensor products with Wenzl fundamental representation, see Sect.
+35 in [10].
+Most importantly, we describe a connection between Wenzl and Wassermann work via Wenzl quantisation curve, and building on their work in [10] we obtain Theorems 2.1, 2.2, Cor.
+2.3, and Theorem 2.4 which compares with the tensor product structure by Huang and Lepowsky, to which we refer the reader also for important historical references.
+The connection from affine Lie algebra setting to Huang-Lepowsky tensor product theory may be found in Sections 33–38 in [10].
+A main idea that we follow is to consider the fundamental representation $V$ of $\mathfrak{g}$} for each Lie type as in [99] where the author explains its generating property and good fusion structure with respect to tensor products with an arbitrary irreducible representation allowing a well-defined truncation procedure to obtain non-negligible objects, with the aim of studying unitarizability of the fusion category.
+As anticipated in the abstract, the restriction to the mentioned Lie types is due to the fact that we need to use a key property of the theory of quantum groups, namely the extension to the quantum case of generalized Schur-Weyl duality.
+Currently, this property is known to hold for these Lie types.
+This is part (d) of the same theorem.
+The final stage of the proof, detailed in Sect.
+9–11, relies on a process of cohomological reduction and heavily utilizes extensions of quantum Schur-Weyl duality.
+By explicitly exploiting the generating property of the braid group representation within the centralizer algebras of the truncated tensor powers of the fundamental representation, we achieve the direct equivalence with Huang-Lepowsky braided tensor structure of ${\rm Rep}(V_{{\mathfrak{g}}_{k}})$ necessary to solve Huang’s problem for the classical and $G_{2}$ Lie types.
+The paper is organized as follows.
+In Sect.
+2 we give the main definition of unitary coboundary weak Hopf algebra and recall the construction of $A_{W}({\mathfrak{g}},q)$.
+In Sect.
+3 we state the main result, Theorem 3.1.
+Sects 4–7 explain the main ideas to prove parts (a)–(d) of this result.
+We give details for the completion of the proof of part (d) in Sects 9–11.
+Sect.
+12 concerns connection with Doplicher-Roberts theorem regarding the role of generalized Schur-Weyl duality.
+Sect.
+13 concerns future developments.
+We put in the Appendix historical remarks on the proof of rigidity in module categories of affine Lie algebras and affine vertex operator algebras at positive integer level.
+2. Unitary coboundary weak Hopf algebras $A_{W}({\mathfrak{g}},q)$
+The notion of weak quasi-Hopf algebra was first introduced by Mack and Schomerus motivated by the work of Drinfeld on quasi-Hopf algebras, specifically the Drinfeld-Kohno theorem [70].
+The weakness refers to a non-unital coproduct, and is needed to construct a coalgebra structure on the quotient $A$ of $U_{q}({\mathfrak{g}})$ with $q$ a root of unity by the ideal of elements annihilated by representations with zero quantum dimension.
+Since the ideal is not a coideal, the quotient algebra $A$ is not a bialgebra in the natural way.
+For ${\mathfrak{s}l}_{2}$, Mack and Schomerus defined a weak coproduct on $A$ and an associator derived from the fusion rules of the semisimple fusion category ${\mathcal{C}}({\mathfrak{s}l}_{2},q)$, a counit and a natural antipode derived from $U_{q}({\mathfrak{s}l}_{2})$, and have verified the axioms of a weak quasi-Hopf algebra on their example.
+Their construction is based on the special multiplicity free fusion rules of the category.
+In [10] we have extended Mack and Schomerus construction to all Lie types and to all roots of unity of the unitary type–i.e.
+the roots of unity appearing in the statement of our main Theorem 1.6–and we have derived algebraic and analytic properties that were overlooked in the literature.
+The main algebraic property that our quotient algebra $A_{W}({\mathfrak{g}},q)$ of $U_{q}({\mathfrak{g}})$ satisfies is that of being the Hopf counterpart of a weak quasi-Hopf algebra, in a position completely analogous to Hopf algebras among quasi-Hopf algebras.
+For this reason, we shall refer to $A_{W}({\mathfrak{g}},q)$ as a weak Hopf algebra.
+We have established the weak Hopf algebra structure of $A_{W}({\mathfrak{g}},q)$ building upon the quantum group fusion category ${\mathcal{C}}({\mathfrak{g}},q)$ and Wenzl’s linear fibre functor $W$ associated to it [99].
+Here we limit ourselves to recall the axioms of the definition of weak Hopf algebra which mostly differ from those of Hopf algebras.
+For the complete properties of antipode, counit and $R$-matrix of weak Hopf algebras we refer the reader to Sects.
+9–14 of [10].
+Definition 2.1.
+Let $A$ be a semisimple unital complex algebra endowed with a coproduct $\Delta:A\to A\otimes A$ given by a not necessarily unital homomorphism and a counit satisfying the usual counit axioms for bialgebras.
+The tensor product $\rho\otimes\sigma$ of two f.d.
+representations of $A$ is the representation defined by $\Delta$ restricted to the range of $\Delta(I)$ in the tensor product space.
+We call $A$ a weak bialgebra if the category ${\rm Rep}(A)$ of representations of $A$ becomes a tensor category with associativity morphisms
+$\alpha_{\rho,\sigma,\tau}:(\rho\otimes\sigma)\otimes\tau\to\rho\otimes(\sigma\otimes\tau),\quad\rho,\sigma,\tau\in{\rm Rep}(A)$
+defined by the action of $1\otimes\Delta(\Delta(I))\Delta\otimes 1(\Delta(I))$ between the representation spaces, with $I$ the unit of the algebra, and $1$ the identity map.
+In particular, if the coproduct is unital, $A$ is a coassociative bialgebra.
+A weak Hopf algebra is a weak bialgebra endowed with an antipode $S:A\to A$ satisfying the axioms of a Hopf algebra antipode.
+Among weak quasi-bialgebras, weak bialgebras admit a characterization in terms of weak tensoriality of their fibre functor.
+Proposition 2.2.
+A semisimple weak quasi-bialgebra $A$ is a weak bialgebra if and only if the fibre functor ${\mathcal{F}}:{\rm Rep}(A)\to{\rm Vec}$ is a weak tensor functor.
+By this we mean that the associated natural transformations given by projection and inclusion, respectively:
+$F_{\rho,\sigma}:{\mathcal{F}}(\rho)\otimes{\mathcal{F}}(\sigma)\to{\mathcal{F}}({\rho\otimes\sigma})$
+and
+$G_{\rho,\sigma}:{\mathcal{F}}({\rho\otimes\sigma})\to{\mathcal{F}}(\rho)\otimes{\mathcal{F}}(\sigma)$
+satisfy for all objects $\rho$, $\sigma$, $\tau$,
+(2.1) $F_{\rho,\sigma}\circ G_{\rho,\sigma}=1_{{\mathcal{F}}({\rho\otimes\sigma})}$
+(2.2) $F_{\rho^{\prime},\sigma^{\prime}}\circ{\mathcal{F}}(S)\otimes{\mathcal{F}}(T)={\mathcal{F}}(S\otimes T)\circ F_{\rho,\sigma},\quad{\mathcal{F}}(S)\otimes{\mathcal{F}}(T)\circ G_{\rho,\sigma}=G_{\rho^{\prime},\sigma^{\prime}}\circ{\mathcal{F}}(S\otimes T)$
+for objects $\rho$, $\sigma$, $\rho^{\prime}$, $\sigma^{\prime}\in{\mathcal{C}}$ and morphisms $S:\rho\to\rho^{\prime}$, $T:\sigma\to\sigma^{\prime}$.
+(2.3) ${\mathcal{F}}(\alpha_{\rho,\sigma,\tau})=F_{\rho,\sigma\otimes\tau}\circ 1_{{\mathcal{F}}(\rho)}\otimes F_{\sigma,\tau}\circ\alpha^{\prime}_{{\mathcal{F}}(\rho),{\mathcal{F}}(\sigma),{\mathcal{F}}(\tau)}\circ G_{\rho,\sigma}\otimes 1_{{\mathcal{F}}(\tau)}\circ G_{\rho\otimes\sigma,\tau}$
+(2.4) ${\mathcal{F}}(\alpha_{\rho,\sigma,\tau}^{-1})=F_{\rho\otimes\sigma,\tau}\circ F_{\rho,\sigma}\otimes 1_{{\mathcal{F}}(\tau)}\circ{\alpha^{\prime}}^{-1}_{{\mathcal{F}}(\rho),{\mathcal{F}}(\sigma),{\mathcal{F}}(\tau)}\circ 1_{{\mathcal{F}}(\rho)}\otimes G_{\sigma,\tau}\circ G_{\rho,\sigma\otimes\tau},$
+where $\alpha$ are the associativity morphisms of ${\rm Rep}(A)$ defined by the associator of $A$, and $\alpha^{\prime}$ are the (trivial) associativity morphisms of $\rm Vec$}.
 We refer to Sect.
-25 in [13] for an extension of certain results by Neshveyev and Yamashita in the setting of quantum group fusion categories of type $A$ based on Kazhdan-Wenzl theory at roots of unity in our case, and on the weak Hopf algebras constructed in [12], in place of the compact quantum groups.
-This section is important for us as it suggested that in general the associativity morphisms in a category arising from CFT may be of a simple nature from the viewpoint of the associated quantum gauge groups that need to be constructed, and this finally proved to be the case, we shall come back to this with more precision.
-On the other hand, in the process of extending Doplicher-Roberts theory in the setting of compact quantum groups, we had started by embedding in a tensorial way part of the morphisms of the category into the Hilbert spaces.
-This was the tensor subcategory generated by solutions of the conjugate equations [83], [85], [89].
-Via Tannaka-Krein-Woronowicz duality this subcategory corresponds to Wang and Van Daele orthogonal or universal compact quantum groups $A_{o}(F)$, or $A_{u}(F)$ [97], their representation category has been studied by Banica [2], [3].
-The choice of the compact quantum group arising in this way depends on whether a generating object $\rho$ is selfconjugate as in the case of $A_{o}(F)$ or not as for $A_{u}(F)$.
-In doing this embedding we had already made a choice of a dimension function, the rank of the associated invertible matrix $F$, and from that choice we had constructed a tensor functor from that subcategory.
-To consider the full category, the analogue Doplicher-Roberts compact gauge group is described by an ergodic action of the corresponding quantum group on a quantum homogeneous space rather than a quantum group itself, due to lack of commutativity of this $C^{*}$-algebra [85].
-The homogeneous space is described by a $C^{*}$-algebra, naturally associated to the category and to the embedding of the subcategory.
-This property is likely to hold also in other models arising from conformal field theory starting from the minimum energy functor, unless one can show that certain known obstructions already met in [84] for compact quantum groups, vanish, we shall come back to this in the outlook section 13.
-However, for the WZW models arising from affine Lie algebras at positive integer levels in [12] for the type $A$ case, and in [13] for all the Lie types, we obtain genuine quantum groups as quantum gauge groups, described as semisimple weak Hopf algebras with $R$-matrix, $C^{*}$-structure explicitly associated to Drinfeld coboundary matrix, and with unitary representations of the braid group, derived from the quantum group fusion category.
-We prove that these weak Hopf algebras have representation category equivalent to both the starting fusion category for all the Lie types and also to that of the corresponding affine vertex operator algebra with ribbon braided Huang-Lepowsky tensor structure, for the classical Lie types and $G_{2}$.
-Constructing the equivalence from the former to the latter in the setting of semisimple weak quasi-Hopf algebras as we do in [13] for these Lie types is our proof of Kazhdan-Lusztig-Finkelberg theorem, see Theorem 2.1.
-Among our first references in the context of conformal field theory important for us to start to set up the connection with the work of Wenzl in quantum groups, is the work by Wassermann [98] in the setting of affine Lie algebras, where the key concept is played by the notion of primary field, and its initial term, which we use in a fundamental way to connect quantum group fusion category to representations of the Zhu algebra and their special tensor products with the fundamental representation, see Sect.
+6, 7, 10 in [10] for the general definition of weak quasi-Hopf algebra, their $R$-matrix how to obtain a rigid braided tensor category from it.
+This structure of weak quasi-Hopf algebra is stable under a twist operation–first introduced by Drinfeld for quasi-Hopf algebras–.
+The twist of a weak quasi-Hopf algebra is automatically a weak quasi-Hopf algebra, and provides an equivalent rigid tensor category.
+We next describe the properties concerning the unitary structure of weak quasi-Hopf algebras that are most relevant for this paper.
+Definition 2.3.
+A (compatible) unitary coboundary weak quasi-Hopf algebra $A$ is defined by the following data:
+  • a)
+    A unital weak quasi-Hopf algebra $A$ endowed with a $C^{*}$-algebra involution with an antipode $(S,\alpha,\beta)$
+  • b)
+    a ribbon structure $(R,v)$ for $A$ associated to $(S,\alpha,\beta)$ such that the ribbon element $v\in A$ is unitary,
+  • c)
+    a unitary central square root $w\in A$ of $v$ such that $\varepsilon(w)=1$, $S(w)=w$,
+  • d)
+    $\tilde{A}=A^{{\rm op}}$ as quasitriangular weak quasi-bialgebras,
+  • e)
+    $\overline{R}=Rw^{-1}\otimes w^{-1}\Delta(w)$ is positive.
+The element $R$ is the $R$-matrix; $v\in A$ is a central element defining the ribbon structure; the operator $\overline{R}$ is Drinfeld coboundary matrix and gives the $\Omega$-involution $\Omega=\overline{R}$ in the sense of Sect.
+11, 12 in [10].
+The $\Omega$-involution induces a -involution on ${\rm Rep}(A)$ making it into a tensor -category.
+In the case where the -involution is induced by a coboundary matrix, the braided symmetry of ${\rm Rep}(A)$ induced by the $\overline{R}$-matrix is unitary by Theorem 27.9 in [10].
+The triple $(S,\alpha,\beta)$ describes the antipode of a weak quasi-Hopf algebra, and makes ${\rm Rep}(A)$ into a rigid category.
+The presence of invertible elements $\alpha$, $\beta$ is necessary in the quasi-cases, it arises from twist deformations of the usual antipode of a Hopf algebra.
+If $A$ is weak Hopf algebra, its antipode is required to satisfy the same axioms as a Hopf algebra antipode: is defined with $\alpha=\beta=I$; $\varepsilon$ denotes the counit and satisfies the same axioms as in the Hopf algebra case; $A^{{\rm op}}$ and $\tilde{A}$, denote the opposite and adjoint weak quasi-Hopf algebras.
+The opposite structure $A^{{\rm op}}$ is widely used in the literature, while the adjoint structure $\tilde{A}$ is defined by coproduct $\tilde{\Delta}(a)=\Delta(a^{*})^{*}$, and similarly for the remaining structure, see Sect.
 10 in [10].
-Most importantly, we describe a connection between Wenzl and Wassermann work via Wenzl quantisation curve, and building on their work in [13] we obtain Theorems 3.1, 3.2, Cor.
-3.3, and Theorem 3.4 which compares with the tensor product structure by Huang and Lepowsky, to which we refer the reader also for important historical references.
-The connection from affine Lie algebra setting to Huang-Lepowsky tensor product theory may be found in Sections 34–39 in [13].
-Our abstract duality methods need new further tools as we shall see from the theory of quantum groups, mainly from the work of Drinfeld and Wenzl, that are not available in the theory of groups.
-However, we shall find that in both contexts, quantum groups at roots of unity and affine vertex operator algebras at positive integer levels, both representation categories are described by functors extending the symmetric tensor functors in Doplicher-Roberts abstract duality theory, see Sect.
+Except for the $C^{*}$-structure of the algebra and the positivity requirement of the coboundary operator in e), the axioms are satisfied by the -algebra of $U_{q}({\mathfrak{g}})$, which for this reason represents the motivating example of our definition.
+These properties of $U_{q}({\mathfrak{g}})$ and positivity of a truncated coboundary operator operator of $U_{q}({\mathfrak{g}})$ on certain special tensor products are due to Wenzl [99], see also Sect.
 30 in [10].
-We call these structures hermitian coboundary weak quasi-Hopf algebras, as the hermitian structure is given by Drinfeld coboundary matrix, which is a modification of the $R$-matrix with a square root of the ribbon structure.
-We notice on the other hand that a stronger property is needed to establish the connection, than the purely categorical characterisation of functors in Sect.
-29 in [13].
-This property is expressed in terms of quantum groups, and the first example, in a non-semisimple setting, is the hermitian structure of $U_{q}({\mathfrak{g}})$ at a root of unity by Theorem 31.9 in [13].
-This property says that the -involution enjoys a precise anticommutation relation (rather than commutation relation as is familiar in Hopf $C^{*}$-algebras) with the coproduct and was first noticed in [99].
-We axiomatize this property in Sect.
-30 in [13], as a compatibility property of hermitian coboundary weak quasi-Hopf algebras.
-Further examples are our weak bi-algebras $A_{W}({\mathfrak{g}},q)$ associated to Wenzl functor $W:{\mathcal{C}}({\mathfrak{g}},q)\to{\rm Vec}$ and to the quasi-bialgebras structures on the Zhu algebra $A({V_{{\mathfrak{g}}_{k}}})$ associated to Zhu functor ${\rm Rep}({V_{{\mathfrak{g}}_{k}}})\to{\rm Vec}$ via our Drinfeld twist construction, see Sections 32, 34 in [13].
-The anticommutation property of the -involution and coproduct of $A_{W}({\mathfrak{g}},q)$ is not a direct consequence of the corresponding property of $U_{q}({\mathfrak{g}})$, as its coproduct is not compatible with the fusion tensor product.
-We need some delicate analysis of the fusion tensor product described in [99].
-For $A({V_{{\mathfrak{g}}_{k}}})$, we show that the twist makes the coproduct of $A({V_{{\mathfrak{g}}_{k}}})$ symmetric, and this is a result analogous to the symmetry property of the usual coproduct of Drinfeld quasi-Hopf algebra $U({\mathfrak{g}})[[h]]$.
-For details we refer to the cited sections.
-On the other hand, as an algebra $A_{W}({\mathfrak{g}},q)$ is a natural quotient of $U_{q}({\mathfrak{g}})$, and the antipode of $A_{W}({\mathfrak{g}},q)$ is directly derived from that of $U_{q}({\mathfrak{g}})$.
-In this way, $A_{W}({\mathfrak{g}},q)$ becomes a weak Hopf algebra.
-It follows that ${\rm Rep}(A_{W}({\mathfrak{g}},q))$ becomes a rigid ribbon braided tensor category naturally equivalent to ${\rm Rep}(A({V_{{\mathfrak{g}}_{k}}}))$ via our twist, and therefore so does ${\rm Rep}({V_{{\mathfrak{g}}_{k}}})$ via Zhu equivalence.
-As above motivated, we shall derive more structure beyond rigidity and modularity on ${\rm Rep}({V_{{\mathfrak{g}}_{k}}})$ from the quantum group fusion category, a further example is unitarization of the vertex operator algebra representation category from the corresponding quantum group fusion category, via the explicit equivalence, see Sects.
-22, 23, 24, 26 in [13].
-To this aim it is important that our equivalence passes through the weak quasi-Hopf algebras, and the fact that these are naturally associated to the categories.
-Indeed it was known that a mere tensor equivalence between the categories would not suffice to transport the unitary structures see [49].
-Sect.
-16 in [13] treats this for general tensor categories that admit submultiplicative dimension functions (called weak dimension functions).
-Our equivalence, since it is given by a Drinfeld twist analogous to [25], [27], and is also motivated by the requirement of making the unitary structure classical, at the level of the associated weak Hopf or quasi-Hopf algebras gives an understanding of the unitary structures of representation categories of affine vertex operator algebras as classical compact real forms, or untwisted.
-We refer to the outlook section for some more information on possible extension of our viewpoint to models beyond the affine vertex operator algebras at positive integer levels, see also [13].
-Our main result is Theorem 2.1.
-In the subsequent sections we outline the plan of our proof.
-A main idea that we follow is to consider the fundamental representation $V$ of ${\mathfrak{g}}$ for each Lie type as in [99] where the author explains its generating property and good fusion structure with respect to tensor products with an arbitrary irreducible representation allowing a well-defined truncation procedure to obtain non-negligible objects, with the aim of studying unitarizability of the fusion category.
-It should also be noted that a main application of that paper is construction of inclusion of subfactors of finite Jones index [63] for all Lie types, extending his previous work.
-For all the Lie types we obtain complete identification for the tensor product structure, all the ribbon structure, and those braiding and associativity morphisms where one or two variables are the fundamental representation $V$, respectively, with Huang-Lepowsky ribbon braided tensor structure.
-These are parts (a), (b), (c) of Theorem 2.1.
-The term weak in the setting of quasi-Hopf algebras was first introduced by Mack and Schomerus motivated by the work of Drinfeld for quasi-Hopf algebras and Drinfeld-Kohno theorem [71].
-Weakness refers to a non-unital coproduct, needed to exclude the non-physical representations with zero quantum dimension, and with the need of compatibility with the fusion rules of the category arising from a suitable model in CFT, described by a quantized universal enveloping algebra of ${\mathfrak{s}l}_{2}$ at a root of unity.
-Mack and Schomerus did not consider the Hopf counterpart in their theory.
-Weak Hopf algebras were later introduced by Böhm and Szlachanyi motivated by the work of Mack and Schomerus, but with a strictly coassociative coproduct [7].
-The definition of weak Hopf algebra has become standard and understood in the sense by Böhm and Szlachanyi.
-In [12], [13], and also in this paper, weak Hopf algebras are understood with a different meaning.
-While a different name would be helpful to avoid confusion, so far we have not been able to find a good name for our algebras to clarify the situation, and we use the term weak Hopf algebras in the sense defined in Sect.
-10 in [13].
-In [13] we develop a general theory of weak Hopf algebras, review and expand the theory of weak quasi-Hopf algebras, mainly Tannaka-Krein aspects, both in the algebraic and in the $C^{*}$-algebraic setting, with non-trivial unitary structure, discuss a suitable Drinfeld-Kohno theorem, and its application to the construction of semisimple weak Hopf and quasi-Hopf algebras with structure naturally compatible with that of the involved ribbon braided tensor categories, for quantum groups at roots of unity and affine vertex operator algebras at positive integer levels, including unitary structure as heredited from the unitary structure of the category due to Wenzl [99].
-We follow the strategy of Drinfeld for Drinfeld-Kohno theorem which includes the comparison of braiding morphisms of the two settings, but we do not include the comparison of the associativity morphisms in our analogue, Theorem 30.4 in [13] (more details are given later).
-We construct a Drinfeld twist for our weak Hopf algebras useful for the braided symmetry in a way similar to Drinfeld method, by taking a square root of the coboundary matrix, in our case this is slightly more delicate because of the truncation procedure and analytic aspects.
-Similarly to Drinfeld-Kohno theorem, we establish equality of the associativity morphisms arising from quantum group fusion category after the twist applied to our weak Hopf algebras with Huang-Lepowsky associativity morphisms via a uniqueness theorem of a cohomological nature, but in our case this theorem seems different, in that we use cohomology of the pentagon equation on the generating representation $V$ and extensions to quantum case and Lie types of Schur-Weyl duality.
-We shall come back to this in the last part of the introduction and in the following sections of this paper.
-We apply our concrete twist to transport unitary structures to representation categories of affine vertex operator algebras at positive integer levels with the tensor structure due to Huang and Lepowsky, for all the mentioned Lie types, by transferring the unitary structure from quantum group fusion category as constructed in [99].
-On the other hand, the remaining Lie types may be either derived from Finkelberg equivalence, making now the ribbon braided fusion category $\tilde{\mathcal{O}}_{k}$ unitary, or treated with our partial results in the setting of Huang-Lepowsky structure for representation categories of vertex operator algebras.
-Direct constructions of unitary structures in this setting have been obtained by Gui as already mentioned [43], [44], [45], [46].
-Our weak Hopf algebras while having a non-unital coproduct, are not in the class of Bohm and Szlachanyi.
-For example their coproduct is not strictly coassociative.
-They are not extension of Bohm and Szlachanyi weak Hopf algebras, and we know from results in [13] that the intersection of the two classes are the genuine Hopf algebras.
-On the other hand, in our opinion the weak Hopf algebras considered in [12] and [13] still deserve to be regarded as Hopf algebras in a weak sense, for various reasons.
-On one hand they seem to play the role of the quantized universal enveloping algebras in Drinfeld-Kohno theorem, following the original motivation of Mack and Schomerus.
-Moreover, their representation category is characterized by a tensor functor to the category of Hilbert spaces with non-invertible natural transformations (projections endowed with natural inverse inclusions, in our case) satisfying tensoriality diagrams, which reduces to the notion of tensor functor when the natural transformations are invertible.
-Their counit is antimultiplicative and the antipode satisfy the same relations as in the case of Hopf algebras.
-As already mentioned, we prove the complete identification of braiding and associativity morphisms of the two structures for all the classical Lie types $A$, $B$, $C$, $D$, and the case $G_{2}$ in part (d) of Theorem 2.1, by further refining methods of the previous part but we use more structure in Sect.
-9 of [13] that has not emerged in other parts.
-As above mentioned, in the type $A$ case, the extra structure that we need to use is known as quantum Schur-Weyl duality between the vector representation and the natural representation of the braid group between the tensor powers of the vector representations.
-In the classical case, Schur-Weyl duality is a key property for the general linear groups starting with the work of Schur and Weyl.
-We note that Schur-Weyl duality also takes an important role in Doplicher-Roberts abstract duality theory for compact groups and applications to construction of the field algebra.
-This classical duality emergences as the starting point for the construction of global gauge symmetry groups in tensor categories arising in [24].
-Duality with the braid group representation is a concept at the basis of the origin of quantum groups following V. F. R. Jones [64].
-More generally, for Lie types beyond $A$, we need a corresponding duality between the centralizer algebras of the truncated tensor powers of the fundamental representation $V$ of the Drinfeld-Jimbo quantum group $U_{q}({\mathfrak{g}})$ at the root of unity and the natural representation of the braid group in these algebras, and this leads to the restriction stated in (d).
-This allows us to apply our abstract uniqueness Theorem 9.19 in [13] for braided tensor structures in a pre-tensor category.
-In our work, the need of quantum duality between the braid group representation and the fundamental representation $V$ emerged on the occasion of the proof of our earlier Lemma 7.10, which was not complete in version 7 of [13], and we worked on completion with the same methods.
-It was for us a sudden realization of an even greater recognition than we had realized up to that point of the contribution of the authors on extensions of Schur-Weyl duality which has historical interest on his own in representation theory, and more work by Wenzl and his collaborators, and several other authors in the classical and quantum cases for the various Lie types.
-We shall come back to this in more detail.
-In our setting, quantum extension of Schur-Weyl duality to the Lie types has a role to see uniqueness of the associativity morphisms in a tensor category associated to the given braided symmetry with a generating representation $V$, see Theorem 9.19 in [13].
-This applies to both fusion categories associated to quantum groups at roots of unity and of affine VOAs.
-Presently it is not known to the author whether the fundamental representation for the exceptional Lie types $E$ and $F$ satisfies braid group duality.
-The $E_{8}$ case is more delicate, because fusion of tensor powers of $U_{q}({\mathfrak{g}})$ at the root of unity is more complicated, as explained in [99], but more work by Wenzl is available for the study of the centralizer algebra [100].
-We are grateful to M.
-Yamashita for pointing out this reference to us.
-Among possible applications of our proof are the study of unitary structures of representation categories of vertex operator algebras, due to the clarity offered by a quantum group viewpoint of the same problem as explained in [99]; the study of equivalences between representation categories of conformal nets and vertex operator algebras via the construction of a common quantum gauge group; to extension of Doplicher-Roberts field algebra construction from the local observable algebra to the conformal field theory setting.
-Some more discussion may be found in the outlook section 13.
+Compatible unitary weak quasi-Hopf algebras in the sense of Definition 2.3 have been introduced and studied in Sect.
+29 of [10].
+They satisfy the stronger property d) than their weaker variant considered in Sect.
+27 of the same paper, where d) is relaxed up to a so called trivial twist.
+If the coproduct is unital, the two definitions coincide.
+We do not have examples in the weak case that distinguishes among the two definitions.
+The weaker variant is interesting for the fact that it admits a categorical characterization extending the notion of symmetric fibre functor –a core notion in Doplicher-Roberts theory.– to the coboundary symmetry induced by $\overline{R}$ in ${\rm Rep}(A)$.
+The stronger notion on the other hand allows a formulation of an abstract Drinfeld-Kohno theorem, see Theorem 29.4 in [10], that we use in our main theorem.
+In this paper we shall not need the weaker version, therefore we shall always assume the compatibility condition given by property d) when not clearly stated.
+In [99] Wenzl has constructed a linear -functor
+$W:{\mathcal{C}}({\mathfrak{g}},q)\to{\rm Hilb}.$
+The following result has been shown in Theorem 31.16 in [10].
+Theorem 2.4.
+The semisimple -algebra $A_{W}({\mathfrak{g}},q):={\rm Nat}(W)$ of natural transformations of $W$ admits a natural structure of unitary compatible coboundary weak Hopf algebra derived from ${\mathcal{C}}({\mathfrak{g}},q)$ and $W$.
+Thanks to Theorem 29.4 in [10]–our version of Drinfeld-Kohno theorem–the Zhu algebras $A(V_{{\mathfrak{g}}_{k}})$ become unitary coboundary weak quasi-Hopf algebras with a Drinfeld twist from $A_{W}({\mathfrak{g}},q)$, by our version of Drinfeld-Kohno theorem.
+This is summarized in the following section, and detailed in Sect.
+33 of [10].
 3. The main result
@@
     $Z:{\rm Rep}_{\rm HL}(V_{{\mathfrak{g}}_{k}})\to{\rm Rep}_{\rm QG}(A(V_{{\mathfrak{g}}_{k}}))$
-    endowed with the same structure maps as in (a) but w.r.t.
+    endowed with the same natural transformation as in (a) but w.r.t.
 the new ribbon braided tensor structure of ${\rm Rep}_{\rm QG}(A(V_{{\mathfrak{g}}_{k}}))$ described in (b), satisfies the ribbon braided tensor equivalence equations for the ribbon structure for all objects, the braid morphisms for pairs $(V_{\lambda},V)$ and $(V_{\lambda},V)$, and the associativity morphisms for triples
@@
     If $\mathfrak{g}$} is of one of the Lie types $A$, $B$, $C$, $D$, $G_{2}$ then Zhu’s equivalence in (c) is a ribbon braided tensor equivalence. Specifically, the composition
-[…]
+    (3.1) ${\rm Rep}_{\rm HL}(V_{{\mathfrak{g}}_{k}})\xrightarrow{Z}{\rm Rep}_{\rm QG}(A(V_{{\mathfrak{g}}_{k}}))\xrightarrow{}{\rm Rep}(A_{W}({\mathfrak{g}},q))\xrightarrow{{\rm TK\ ribbon\ equiv}}{\mathcal{C}}({\mathfrak{g}},q)$
     is a ribbon tensor equivalence by application of the indicated theorems.
-The definition of unitary coboundary weak Hopf algebra, its restricted version named unitary compatible coboundary weak Hopf algebra, are given in Sects.
-31, 32 in [13], by abstracting features of the hermitian structure of $U_{q}({\mathfrak{g}})$ for $|q|=1$, discovered by Wenzl in [99] in his proof of unitarity of Kirillov inner product, for the associated fusion category as certain roots of unity, based on a complex Drinfeld coboundary matrix.
-The more general notion admits a categorical characterization generalizing Doplicher-Roberts symmetric functors.
-For part (b), we refer the reader to Theorem 3.2 (a) in [13] and further indications therein where the statement is explained.
-We would like to notice that the passages described in (2.1) are explicit.
-The first and the last equivalences are due respectively, to the work by Zhu combined with that of Huang and Lepowsky.
-The last is the functor due to Wenzl enriched with our weak tensor structure, preserving Drinfeld coboundary structure derived from the ribbon and $R$-matrix, see the above cited sections in [13].
-The middle equivalence in (2.1) is application of the explicit symmetry provided by our analogue of Drinfeld-Kohno theorem 30.4 in [13].
+We would like to notice that the equivalences described in (3.1) are explicit.
+The equivalence on the left is Zhu’s linear equivalence endowed with the natural transformations given by the restriction to the zero modes the braided tensor structure by Huang and Lepowsky.
+The fact that the braided tensor structure of the range of $Z$ induced by Huang-Lepowsky structure in (a) coincides with the structure arising from quantum groups on the Zhu algebra, follows from our work in (c), done in [10], and the first statement of part (d), established in this paper.
+The equivalence on the right is the functor due to Wenzl enriched with our weak tensor structure by Theorem 2.4.
+The middle equivalence is an application of the explicit symmetry provided by our analogue of Drinfeld-Kohno theorem 29.4 in [10], and stated in parts (a) and (b) of Theorem 2.2 in [10].
 This symmetry is given by a Drinfeld twist explicitly defined by the action of the $R$-matrix, more precisely by the braided and ribbon structure in the two settings, that allow to see in the vertex operator algebra setting same structure as that in the quantum group setting.
-3. Motivation and framework of our work. Constructing braided tensor equivalences via a quantum gauge group given by a discrete weak Hopf algebra and a Drinfeld twist.
-Before going into the explanation of the statement and outline the main tools of the proof in more detail, we summarize the general framework that motivated and led to our work.
-In the following, where not indicated the cited sections and results refer to [13], and also for unexplained notation.
-We were originally motivated by Doplicher-Roberts abstract duality theory for compact groups and their application to the construction of a global compact gauge group and reconstruction of the net of field algebras from the observable algebras in high dimensional quantum field theories [23], [24].
-In the setting of low dimensional quantum field theories, starting with the fusion categories of quantum groups at all the roots of unity that are known to correspond to unitary structures by the work of Wenzl and Xu [99], [105], we in particular construct discrete quantum group-like structures with non-trivial twisted unitary structures for all the Lie types as in [99].
-These discrete quantum groups are thought of as duals of compact quantum groups, but their preduals admit no analogy to the case of the compact quantum groups of Woronowicz, because they are non-associative algebras.
-However, Woronowicz work was strongly inspiring for us.
-Indeed our first examples in the type $A$ case were first constructed in our previous work [12] via an abstract Tannaka-Krein method in the search of a noncommutative function algebra, using the fusion rules of the corresponding fusion category.
-On the other hand, their duals are quite well-behaved discrete $C^{*}$-algebras with almost associative coproducts, antimultiplicative antipode and multiplicative counit satisfying the same relation as for Hopf algebras.
-In this sense they resemble the duals of the compact quantum groups, the main differences being non-unitality of the coproduct and the unitary structure of the tensor product of their representations, which is twisted via a positive operator associated to the $R$-matrix (Drinfeld coboundary matrix), and makes the representation of the braid group unitary.
-The need of a twist for the tensor product representation arises from the fact that coproduct and -involution satisfy an anticommutativity condition, rather than commutativity as it usually is the case for Hopf algebras in the operator algebra setting.
-This anticommutativity is derived from the same property of the -involution of the non-semisimple $U_{q}({\mathfrak{g}})$, first pointed out by Wenzl [99].
-The weak Hopf algebras constructed in [12] have been extended to all Lie types in Sections 31, 32 of [13] with uniform methods for all the Lie types.
-The construction of associativity morphisms satisfying the two hexagonal equations with respect to natural braiding morphisms associated to representations of the universal enveloping algebra $U({\mathfrak{g}})[[h]]$ of a simple Lie algebra (in a formal setting) with classical coproduct, were first solved by Drinfeld with the introduction and construction of quasi-Hopf algebras, via the Knizhnik-Zamolodchikov differential equation.
-This category is known as Drinfeld category, motivated by previous work by Kohno on monodromy of a connection associated to the vector representation for all ${\mathfrak{g}}$ of classical type [69].
-Drinfeld developed the remarkable and beautiful Drinfeld-Kohno theorem connecting $U({\mathfrak{g}})[[h]]$ with its natural $R$-matrix and non-trivial co-associator with the quantized universal enveloping algebra $U_{h}({\mathfrak{g}})$ via a Drinfeld twist, describing a ribbon braided tensor equivalence.
-In this case, rigidity of Drinfeld category is derived from the equivalence constructed by Drinfeld in Drinfeld-Kohno theorem [25], [27], [28].
-Our algebras are examples of Mack-Schomerus weak quasi-Hopf algebras introduced in [71] as generalization of Drinfeld quasi-Hopf algebras with non-unital coproduct.
-For ${\mathfrak{g}}={\mathfrak{s}l}_{2}$ our examples reduce to Mack-Schomerus examples for the algebraic structure.
-We understand their lack of coassociativity as quite mild, because their representation category is endowed with a weak unitary tensor functor in the category of Hilbert spaces.
-Thus we regard them as the Hopf counterpart of weak quasi-Hopf algebras, in a position similar to that of Hopf algebras among quasi-Hopf algebras, that in our knowledge has been overlooked.
-Here the term weak is understood in the sense first introduced by Mack and Schomerus, the coproduct is non-unital, a feature needed to exclude the non-physical representations with zero quantum dimension, and with the need of compatibility with the tensor structure of the category.
-The weakness of tensoriality of the embedding functor of their representation category to the category of Hilbert spaces means that the structure natural transformations are surjections onto subspaces of the tensor product spaces endowed with left inverses, rather than isomorphisms, and satisfy two tensoriality diagrams.
-Much of Woronowicz theory can be extended to our case.
-From this via a quantization isomorphism described by Wenzl in [99] and the construction of a Drinfeld twist we construct and describe in some more detail later a quantum group-like structure on the Zhu algebra associated to the affine models described as in the statement of the main result.
-These constructions are in some quantum analogy with Doplicher-Roberts construction in the setting of high dimensional algebraic quantum field theory, in that the categories that we study are not given as Tannakian categories.
-In our case, differently from their case, we have naturally defined linear functors emerging as the minimum energy functor due to Zhu in the setting of vertex operator algebras.
-Also differently form their case, where the linear functor needed to be constructed but the tensor structure was trivial, in our case we need to construct tensor structures associated to these naturally occurring linear functors to the category of vector spaces.
-Our starting point is what we shall refer to as Wenzl functor associated to the quantum group fusion category described in [99].
-This paper plays a central role, both on the side of quantum groups and on the side of affine vertex operator algebras, as we next try to explain.
-As said, on the the side of quantum groups, our approach has been influenced by results in quantum group theory by Woronowicz but also by the dual viewpoint by Majid for quasi-Hopf algebras, extended to weak quasi-Hopf algebras by Haring-Oldenburg.
-We start reviewing these Tannakian reconstruction theorems from the viewpoint of Majid and Haring-Oldenburg in Sect.
-6, 7, in [13], where the reader may find indication of the original references.
-Moreover Kazhdan and Wenzl theory is also important to us, and the reader may consult Sect.
-25 in [13] and references therein.
-It is by no means an exaggeration to assert that our paper [13] would have not existed without the paper by Neshveyev, Tuset on Kazhdan-Lusztig theorem for compact quantum groups [78].
-This paper was strongly inspiring for our strategy, although methods are different.
-For similar reasons, we were also guided by the presentation of Drinfeld-Kohno theorem in the book by Bakalov and Kirillov [1].
-These last two references first inspired our work in the search of a unified treatment following a Tannakian approach in the idea of Majid, in our case.
-An important aspect in our case as far as Majid Tannakian duality for quasi-Hopf algebras, is the search of natural tensor structures for the involved embedding functors into the category of Hilbert spaces (Wenzl functor and Zhu functor in the two settings).
-Primary fields first emerged in the physics literature, [5], using previous work by Feigin and Fuks [32] and Kac [65] on the representation theory of the Virasoro algebra.
-In 1984 Witten introduced what are now known the WZW models, due to Wess, Zumino, Novikov and Witten [103].
-In the same year, Knizhnik and Zamolodchikov showed that their correlation functions satisfy a system of partial differential equations [68].
-Tsuchiya and Kanie started a rigorous study in [95]. Goddard, Kent and Olive realized the chiral unitary Virasoro minimal models via a coset construction [40].
-Moore and Seiberg conjectured some analytic properties of the quantum fields and described the construction of a modular category associated to a general conformal field theory.
-In the same paper [69] Kohno had also considered monodromy of $n$-point functions of a conformal field theory on ${\mathbb{P}}^{1}$ associated to the affine Lie algebra $\hat{\mathfrak{s}l}_{2}$.
-The mathematical contribution by Tsuchiya and Kanie [95], and the subsequent works by A.
-Wassermann [98] and Toledano-Laredo [94] on affine Lie algebras, compared with the structure that had previously emphasized by Wenzl in the study ofquantum group fusion categories and their unitary structures [99] that led to our approach.
-The reader may find reference to more of the several original papers in mathematics and physics in Sections 34–39 in [13].
-4. The general plan of our proof.
-We briefly explain here the plan of our proof of the main theorem for parts (a), (b), (c), and (d). In the following sections we describe the main ideas to prove these parts in greater detail.
-Part (a) is an application of a general construction of a natural structure of a (ribbon, braided) tensor equivalence on a given equivalence ${\mathcal{C}}\to{\rm Rep}(A)$ of semisimple linear categories endowed with an inverse linear equivalence ${\rm Rep}(A)\to{\mathcal{C}}$, with $A$ a discrete algebra.
-We outline the main ideas of this tensor structure construction in the next section.
-The reader may refer to Section 16 in [13] for the detailed construction.
-For part (b), we follow a suitable Drinfeld twist method originally used by Drinfeld for Drinfeld-Kohno theorem [25], [27], [28], and in our case for different categories, defined in the statement of the main result.
-This is the main part of the result.
-The first main step is the construction of a class of finite dimensional weak Hopf $C^{*}$-algebras $A_{W}({\mathfrak{g}},q)$ for all Lie types and positive integer levels, extending the former examples of [12] in the type $A$ case.
-They have unitary coboundary structures associated to the fusion category of quantum groups at roots of unity, associated to a specific linear fiber functor $W:{\mathcal{C}}({\mathfrak{g}},q)\to{\rm Vec}$, that emerges from the work of Wenzl [99].
-Complete definitions of unitary coboundary structures on weak quasi-Hopf algebras are given in [13], and are motivated by the work of several authors, among them Drinfeld, Wenzl, Mack and Schomerus.
-In Sect.
-9 of [13] we prove two main abstract uniqueness results for associativity morphisms and braided symmetries in a semisimple pre-tensor category with a generating object.
-These are Theorem 9.13 and its extension Theorem 9.19.
-We explain here briefly how Theorem 9.19 applies to imply part (d) of Theorem 2.1.
-The proof of parts (b), (c), and (d) is based on the general theory of semisimple weak quasi-Hopf algebras, weak Hopf algebras, and the theory of Drinfeld twist treated in Sections 5–8, and 10–18 in [13].
-This abstract theory splits in two parts.
-Theory on which part (b), (c) of Theorem 2.1 is based is developed in Sections 5–8, and 10–17.
-The applicative part of the proof of part (b), (c) is developed in Sections 20–39.
-The abstract aspects that lead to the proof of part (d) of Theorem 2.1 are the main subject of Section 9.
-For parts (b), (c) we develop an abstract analogue of Drinfeld-Kohno theorem 30.4 in [13], and apply it to the setting of fusion categories of quantum groups at suitable roots of unity and of module categories of affine vertex operator algebras at positive integer levels.
-We construct weak Hopf algebras $A_{W}({\mathfrak{g}},q)$ associated to the quantum group fusion category for all the Lie types in Section 32 in [13], by constructing a weak tensor structure on a linear functor from this category to the category of vector spaces, that emerges in [99] for the special tensor products described in (b), (c).
-The weak tensor structure is constituted by fusion intertwiners with respect to the action of $U_{q}({\mathfrak{g}})$.
-By application of our analogue of Drinfeld-Kohno theorem we obtain the structure of a weak quasi-Hopf algebra on the Zhu algebra $A(V_{{\mathfrak{g}}_{k}})$ associated to a corresponding affine vertex operator algebra $V_{{\mathfrak{g}}_{k}}$.
+4. The general plan of our proof
+Part (a) of Theorem 3.1 follows from the application to Zhu’s functor of the general construction.
+Let us consider a (ribbon, braided) semsimple tensor category $\mathcal{C}$} together with a discrete complex algebra $A$ (a direct sum of full matrix algebras), a linear equivalence ${\mathcal{C}}\to{\rm Rep}(A)$ and an inverse linear equivalence ${\rm Rep}(A)\to{\mathcal{C}}$.
+We wish to use these equivalences to transfer all the structure of $\mathcal{C}$} to ${\rm Rep}(A)$, and make it into a (ribbon, braided) tensor category in such a way that the given equivalences become (ribbon, braided) tensor equivalences.
+The definition of the structure on ${\rm Rep}(A)$ is very natural and verification of the axioms of (ribbon, braided) tensor category and the natural transformations on the equivalences and the needed properties is straightforward.
+The special case of making ${\rm Rep}(A(V_{{\mathfrak{g}}_{k}}))$ into a ribbon braided tensor category from Huang-Lepowsky structure on ${\rm Rep}(V_{{\mathfrak{g}}_{k}})$ corresponds to pass from the structure on a category of infinite dimensional modules to the category of their zero modes.
+We have explicitly described this construction in Sect.
+37 in [10].
+This construction was also considered in [72].
+We have used a similar construction, in the reverse direction, and for a different purpose in Theorems 15.6 and 22.10 of [10].
+For part (b), we follow a suitable Drinfeld twist method originally used by Drinfeld for Drinfeld-Kohno theorem [22], [24], [25], and in our case for the stated categories.
+This step splits in various parts.
+The first main step is the construction of the class of finite dimensional unitary coboundary weak Hopf $C^{*}$-algebras $A_{W}({\mathfrak{g}},q)$ recalled in Theorem 2.4 for all the Lie types and positive integer levels.
+This construction extends with different methods the previos examples of [9] in the type $A$ case.
+The unitary coboundary structure of $A_{W}({\mathfrak{g}},q)$ naturally arises from the unitary structure of the fusion category of quantum groups at roots of unity, associated to the specific linear fiber functor $W:{\mathcal{C}}({\mathfrak{g}},q)\to{\rm Vec}$, that emerges from the work of Wenzl [99].
+The second main step to prove (b) is the construction of the Drinfeld twist $F$ defined as a suitable square root of $\overline{R}$.
+A third step consists in the development of an abstract analogue of Drinfeld-Kohno theorem 29.4 in [10], that we apply to our twist composed with a quantization isomorphism described in [99] for the unitary structure of irreducible respresentations of the fusion category, between $A_{W}({\mathfrak{g}},q)$ and a corresponding quotient of $U({\mathfrak{g}})$ that identifies with the Zhu algebra $A(V_{{\mathfrak{g}}_{k}})$.
+By application of our analogue of Drinfeld-Kohno theorem we obtain the structure of a unitary coboundary weak quasi-Hopf algebra on the Zhu algebra $A(V_{{\mathfrak{g}}_{k}})$ associated to a corresponding affine vertex operator algebra $V_{{\mathfrak{g}}_{k}}$.
 Thus a natural weak quasitensor structure defined by fusion intertwiners with respect to the action of $\mathfrak{g}$} compatible with the fusion rules of modules of the vertex operator algebra.
 This induces the structure of a ribbon braided tensor category on the module category of $V_{{\mathfrak{g}}_{k}}$, via transport from the the Zhu algebra $A(V_{{\mathfrak{g}}_{k}})$.
+In particular, we get an independent construction of the associativity morphisms and proof of rigidity of ${\rm Rep}(V_{{\mathfrak{g}}_{k}})$, directly arising from the quantum group fusion category via the equivalence induced by the twist.
 By construction, this structure is ribbon braided tensor equivalent to the structure arising from the quantum group fusion category.
@@
 In our case, the unitary structure of the fusion category of quantum groups at roots of unity in [99] is used in the construction of the Drinfeld twist of $A_{W}({\mathfrak{g}},q)$.
-On the module category of $V_{{\mathfrak{g}}_{k}}$ we also have the ribbon braided tensor structure constructed by Huang and Lepowsky, and we need to compare the two structures.
+Part (c) initiates the comparison between Huang-Lepowsky braided tensor structure on ${\rm Rep}_{\rm HL}(V_{{\mathfrak{g}}_{k}})$ and our structure on ${\rm Rep}_{\rm QG}(A(V_{{\mathfrak{g}}_{k}}))$ and the axioms of braided tensor equivalence is done in [10] by a direct comparison for the special braiding morphisms and the special associativity morphisms listed in (c) on the Zhu algebra, passing through (a).
+Up to this point we do not need any knowledge on the space of intertwiners in the two categories, the linear equivalence is an isomorphism between any two corresponding morphism spaces by the Drinfeld twist construction.
+The necessary general theory of semisimple weak quasi-Hopf algebras, weak Hopf algebras, and the theory of Drinfeld twist treated in Sections 4–7, and 9–17 in [10].
+This abstract theory splits in two parts.
+Theory on which part (b), (c) of Theorem 3.1 is based is developed in Sections 4–7, and 9–16.
+The applicative part of the proof of part (b), (c) is developed in Sections 19–38.
+The abstract aspects that lead to the proof of part (d) of Theorem 3.1 are the main subject of Section 9.
+In Sect.
+8 of [10] we explain two main abstract uniqueness results for associativity morphisms and braided symmetries in a semisimple pre-tensor category with a generating object.
+These are Theorem 8.13 and its extension Theorem 8.19 in [10].
+In the subsequent sections we explain how Theorem 8.19 applies to establish a first step to part (d) of Theorem 3.1.
+These are the main new abstract results to obtain the complete braided tensor equivalence, and rely on key aspects of the discovery of quantum groups.
+In Sect.
+11 we apply these theorems to complete to proof of part (d) of Theorem 3.1.
+In this case we work on the comparison between the two structures transported to the side of the quantum groups, $A_{W}({\mathfrak{g}},q)$, and we need knowledge on the generating property by the braid group between tensor powers of the fundamental representation, to apply our uniqueness theorem.
 5. Outline of the construction of a weak quasi-tensor structure on the Zhu algebra $A(V_{{\mathfrak{g}}_{k}})$ via Wenzl quantisation isomorphism and Drinfeld twist method.
@@
 Direct verification of coincidence of the two braided tensor structures on all the irreducible objects is more complicated due to multiplicity of fusion and non-semisimplicity of $U_{q}({\mathfrak{g}})$.
-To this aim rather than pursuing a closer analogy with the original Drinfeld-Kohno theorem for all objects, we develop two uniqueness theorems, Theorem 9.13 and its generalization Theorem 9.19.
-This generalization is made necessary by the actual conclusions reached in our work in the applicative sections on the comparison between special cases of the two associativity morphisms and their corresponding braided symmetries.
-Thus the uniqueness theorem 9.19 is formulated with suitable assumptions, where again in the application the abstract notion of ${\mathcal{V}}$-pre-associator of CFT-type for the fundamental representation $V$ is used in a central way both on the side of quantum groups and that of vertex operator algebras to verify coincidence of these special braid and associativity morphisms, after application of our Drinfeld-Kohno theorem.
+To this aim rather than pursuing a closer analogy with the original Drinfeld-Kohno theorem for all objects, we develop two uniqueness theorems, Theorem 8.13 and its generalization Theorem 8.19 in [10].
+This generalization is made necessary by the actual conclusions reached in our work in the applicative sections on the direct comparison between special cases of the two associativity morphisms and their corresponding braided symmetries.
+Thus the uniqueness theorem 8.19 is formulated with suitable assumptions, where in the application the abstract notion of $\mathcal{V}$}-pre-associator of CFT-type for the fundamental representation $V$ is used in a central way both on the side of quantum groups and that of vertex operator algebras to verify coincidence of these special braid and associativity morphisms, after application of our Drinfeld-Kohno theorem.
 But this brief description of Theorems 8.13 and Theorem 8.19 put together with coincidence of the two braided tensor structures on special collections of objects proved in part (b), (c), is still not enough to describe our method of proof to complete the proof of part (d) and with it the proof of Theorem 3.1.
 We need more that we describe in the following sections.
-The first next step is a somewhat complicated cohomological reduction work done in Theorems 9.13 and 9.19, where we use cohomology of the pentagon equation and its interplay with the hexagonal equations defining the braided symmetry for a general semisimple braided tensor category endowed with a generating object.
+The first next step is a somewhat complicated cohomological reduction work done in Theorems 8.13 and 8.19, where we use cohomology of the pentagon equation and its interplay with the hexagonal equations defining the braided symmetry for a general semisimple braided tensor category endowed with a generating object.
 The cohomological reduction is useful to reduce the problem to the last step, which uncovers a central duality property between the fundamental representation of the quantum group and corresponding representations of the braid group in the intertwining spaces of the fundamental representation, in the various Lie types.
@@
 This approach is particularly useful for braiding on tensor products of the form $V_{\lambda}\otimes V$ or $V\otimes V_{\lambda}$, where multiplicity-free and structural fusion properties of $V$ allow for a direct verification of the equivalence between the two braided symmetries on these objects.
-Moreover in our case on special triples (or $n$-tuples) where two (or consecutive $n-1$-tuples) coordinates are fixed to $V$, verification of the equivalence of the two associativity morphisms similarly put in a nutshell reduces to the same properties of $V$.
+Moreover in our case on special triples (or $n$-tuples) where two (or consecutive n1n-1-tuples) coordinates are fixed to $V$, verification of the equivalence of the two associativity morphisms similarly put in a nutshell reduces to the same properties of $V$.
 We do not use the KZ equation to this aim, or in any part of our work, other than the fact that the KZ equation was used by Drinfeld, Kazhdan and Lusztig, and by Huang and Lepowsky in the setting of affine vertex operator algebras, to construct the structure of a braided tensor category.
@@
 11. The role of braid group duality for the generating object in determining the full braided tensor structure for the proof of (d).
-The uniqueness Theorems 9.13 and 9.19 both apply under a central duality property of the fundamental representation $V$ with respect to the braid group representation in the quantum group tensor category, which is at the core of quantum group representation theory.
+The uniqueness Theorems 8.13 and 8.19 both apply under a central duality property of the fundamental representation $V$ with respect to the braid group representation in the quantum group tensor category, which is at the core of quantum group representation theory.
 The first instance of this property is Schur-Weyl duality for the general linear group.
@@
 In our case $q$ is a root of unity as usually assumed corresponding to the unitary cases of the associated fusion category, specified in [99], [92], see also Def.
-20.1, 21.2 in [13].
+19.1, 20.2 in [10].
 The tensor powers $V^{n}$ appearing in our setting are canonical choices of truncations of the full tensor powers of the quantum group $U_{q}({\mathfrak{g}})$ obtained by induction over $n$.
@@
 32.
-The duality property between the truncated tensor powers $V^{n}$ of the fundamental representation and the representation of the braid group is needed as a further property to complete the proof of FKL theorem on coincidence between associativity morphisms and braiding morphisms on all pairs and triples and is used in our cohomological theorems 9.13 and 9.19.
+The duality property between the truncated tensor powers $V^{n}$ of the fundamental representation and the representation of the braid group is needed as a further property to complete the proof of FKL theorem on coincidence between associativity morphisms and braiding morphisms on all pairs and triples and is used in our cohomological theorems 8.13 and 8.19.
 These theorems complete with cohomogical arguments what was proved up to that point in our work, coincidence of these structure maps on pairs and triples where one or two coordinated are fixed to the fundamental representation $V$.
@@
 This is explained among other things in Prop.
-2.4 in [99], see also Prop 31.2 in [13] for the same result with a slightly more detailed proof, where an important role is played by the continuous curve described in that proposition, that passes through generic values of $q$.
+2.4 in [99], see also Prop 30.2 in [10] for the same result with a slightly more detailed proof, where an important role is played by the continuous curve described in that proposition, that passes through generic values of $q$.
 See also Section 3.5 in [99] for the special tensor products $V_{\lambda}\otimes V$ of representations of $U_{q}({\mathfrak{g}})$.
@@
 Type $G_{2}$ was shown by Lehrer and Zhang [67], Morrison [75], and Martirosyan and Wenzl [71].
-In the application of Theorems 9.13 and 9.19, we work on the side of the quantum group fusion category and by the results above, our assumptions on braid group duality for the fundamental representation of our weak Hopf algebras $A_{W}({\mathfrak{g}},q)$ constructed in Sect.
+In the application of Theorems 8.13 and 8.19, we work on the side of the quantum group fusion category and by the results above, our assumptions on braid group duality for the fundamental representation of our weak Hopf algebras $A_{W}({\mathfrak{g}},q)$ constructed in Sect.
 31 in [10] are verified for the defining structure arising from Tannakian construction on the quantum group fusion category ${\mathcal{C}}({\mathfrak{g}},q))$ and Wenzl functor.
@@
 Yamashita, we learned that modularity is used in their recent work on classification of fusion categories with fusion rules of type BCD [39].
-Thus the fact that braid group duality for the fundamental representation $V$ was the relevant missing observation helpful to complete the proof of Theorem 2.1 in the previous version of [13], which led us to Theorems 9.13 and 9.19 as a completion of Lemma 7.10 in version 7 of [13], was for us a further important recognition of the relevance of the role of quantum groups.
-We refer the reader to the beautiful historical paper by V.
+Braid group duality for the fundamental representation $V$ was the relevant observation helpful to prove part (d) of Theorem 3.1 with respect to the previous parts of the same theorem, which could reach the equivalence only partially as stated in the previous parts (a)–(c).
+This property was indeed used to apply Theorems 8.13 and 8.19 in [10] and reach the full braided tensor equivalence for the stated Lie types.
+We refer the reader to the historical paper by V.
 F.
@@
 We also would like to point out that our results of the categorical characterization of the class of unitary coboundary weak Hopf algebras that we obtain in Sect.
-26, 27, 29 in [13] extends the notion of symmetric functors that plays a key role in Doplicher-Roberts theory [23], [24].
+25, 26, 28 in [10] extends the notion of symmetric functors that plays a key role in Doplicher-Roberts theory [20], [21].
 In Sect.
@@
 Other interesting future developments may regard applications in noncommutative geometry in the sense of Connes [12] of our weak Hopf algebras $A_{W}({\mathfrak{g}},q)$ associated to the quantum group fusion category ${\mathcal{C}}({\mathfrak{g}},q)$ or their twisted weak quasi-Hopf algebra structures on the Zhu algebra of an affine vertex operator algebra at a positive integer level $V_{{\mathfrak{g}}_{k}}$ to the study of Dirac operators, by expanding an analog in conformal field theory of the groundbreaking work by Neshveyev and Tuset on Dirac operators for compact quantum groups in the setting of Kazhdan-Lusztig theorem for real values of the quantization parameter $q$ [79].
- [^1] The author department digital devices were subject to a severe hacking attack for several months in 2021 which was recognized by authorities.
-The hacking severely damaged her intellectual property and health, and the hacked material has not been restituted.
-Related results have appeared in the papers [8], [9], [10].
-These papers have been sued by the author to the Italian Procura della Repubblica between fall 2021 and spring 2023.
-Their claims, arguments, and reasoning essentially depend on [13], in-depth discussions between the author and S.
-Carpi that took place from 2018 to 2022 on the contents of an entire project including the topics of the above papers before his unilateral break of the collaboration in early 2023 and formal give up of his contribution, that consisted in a draft of a single section in 2018 based on [12], and yet unpublished joint work previously done by Sergio Ciamprone and the author (the first 60-page-long version of [13]), the original Finkelberg theorem, and work by Hans Wenzl in [99].
-The author shared proofs of these attacks to several colleagues from 2022 to 2024.
-Acknowledgements The author is grateful to Yi-Zhi Huang for several comments on the proof of modularity for fusion categories of vertex operator algebras, and the history of Finkelberg-Kazhdan-Lusztig theorem.
+Appendix A Historical remarks on rigidity in affine Lie algebra and vertex operator algebra categories
+The original work by Finkelberg refers to the cited work in physics by Moore and Seiberg and in algebraic geometry by Beilinson, Feigin, and Mazur [4] for the rigid braided tensor structure on $\tilde{\mathcal{O}}_{\ell}$.
+As I learned from discussions with Huang, rigidity was not proved in [4], and there was no proof of rigidity using this approach.
+Remarkably, Huang and Lepowsky introduced and developed a rigorous direct construction of a modular (vertex) tensor category structure in a very general setting for vertex operator algebras, including as a particular case the affine vertex operator algebras at positive integer levels $k$ that we denote by $V_{{\mathfrak{g}}_{k}}$.
+The construction has been completed by Huang between 2005 and 2008 [52], [56] following a general construction of tensor product theory in the setting of vertex operator algebras by Huang and Lepowsky [47], [48], [49] and by Huang [50], [55].
+In particular [50] gives the construction of associativity morphisms which is the first difficulty in the construction of the tensor structure.
+An historical account may be found in [57].
+We also refer to the introductions of [51] and [58].
+These works give in particular the first mathematical proof of modularity, that applies moreover to a large class of representation categories of vertex operator algebras.
+In particular rigidity and modularity property in the setting of vertex operator algebras was first shown in [52].
+To show modularity, and in particular rigidity, Huang first discovered the need of the Verlinde formula.
+The paper [55] gives a proof of the Verlinde formula in the general case.
+See also [53] for the proof of convergence and extension property assumed in [50] and [54] for the proof of modular invariance conjecture of Moore and Seiberg needed in the proof of the Verlinde formula.
+By a private communication with Huang on the history of this problem, I learned that it finally turned out that the construction by Finkelberg described in [30] connecting $\tilde{\mathcal{O}}_{\ell}$ with Kazhdan and Lusztig structure inherited by $\tilde{\mathcal{O}}_{-\ell}$ could not suffice to derive rigidity of $\tilde{\mathcal{O}}_{\ell}$ from that of $\tilde{\mathcal{O}}_{-\ell}$.
+This implied that the proof of the equivalence $\tilde{\mathcal{O}}_{\ell}\to\tilde{\mathcal{O}}_{-\ell}$ was far from complete.
+I have been informed by Huang that these doubts were motivated by two facts.
+On one hand, his earlier direct communications in the 90s had clarified that rigidity was not proved in [4].
+On the other, the mentioned property that his proof of rigidity in the setting of vertex operator algebras involved the Verlinde formula indicated that more work was likely needed to complete the proof in [30].
+These facts together with early claims of rigidity in [30], motivated Huang to a comparison of the results in [30] with their results on rigidity and this led to note a gap in a proposition of [30] with implications on the main result.
+Huang’s study had led him to identify the gap in the fact that a suitable constant needed to be shown to be nonzero in a proposition useful to obtain the equivalence between the two local systems considered by Finkelberg and Kazhdan and Lusztig, and this proposition would imply derivation of rigidity in his case from previous work.
+On the other hand, Huang proof of rigidity in the setting of vertex operator algebras had similarly reduced to show nontriviality of a constant, and to prove this he had used the Verlinde formula as above alluded to, which was a main result in their work.
+This had led Huang to inform Finkelberg about utility of this formula to show rigidity in his case.
+As later reported by Finkelberg in [31], it turned out that the completion of the proof of that proposition leading to rigidity of $\tilde{\mathcal{O}}_{\ell}$ and the equivalence $\tilde{\mathcal{O}}_{\ell}\to\tilde{\mathcal{O}}_{-\ell}$, involved the Verlinde formula as well (which became another main result in Finkelberg’s paper).
+This reflects also on Finkelberg composed equivalence $\tilde{\mathcal{O}}_{\ell}\to\tilde{\mathcal{O}}_{-\ell}\to{\mathcal{C}}({\mathfrak{g}},q)$, with Kazhdan-Lusztig equivalence that moreover still passes through the negative integer levels.
+This implies in particular that the theorem does not include a few exceptional cases, including ${\mathfrak{g}}=E_{8}$ and $k=2$, which correspond to cases not treated in Kazhdan-Lusztig theory.
+Following the introductions in [51], [58], Finkelberg work [30], [31] may be reinterpreted as giving the construction of a braided tensor category structure on $\tilde{\mathcal{O}}_{\ell}$.
+One may also conclude that these papers give a second proof of rigidity and modularity of this category.
+An approach to modularity based on algebraic geometry may be found in Theorem 7.0.1 [1] but the proof of rigidity is not given and, as already mentioned it was known to experts that there was no proof based on their approach at the time.
+The author thanks Y.-Z.
+Huang for this information.
+We point out the very recent paper by Etingof and Penneys where the authors show among other things that weak rigidity in the sense of [1] is equivalent to rigidity for certain semisimple braided tensor categories of moderate growth with simple unit, see [28] giving a third proof of rigidity and give applications to a simplification of Huang’s proof of rigidity.
+They also answer positively to a question posed in [1] that for every finite semisimple category over an algebraically closed field the data of a modular functor is essentially equivalent to the structure of a modular fusion category, completing the proof of rigidity in [1].
+On the other hand, Finkelberg equivalence theorem is stated as Theorem 7.0.2 in [1] without proof.
+The braided tensor equivalence between $\tilde{\mathcal{O}}_{\ell}\to{\mathcal{C}}({\mathfrak{g}},q)$ obtained from the combined works by Finkelberg, Kazhdan and Lusztig is considered indirect by several authors.
+Acknowledgements.
+The methodology and results presented here are the outcome of a long-term project initiated in 2016.
+The author wishes to clarify that S.
+Carpi joined the collaboration in January 2018 and participated in the joint development of a draft section based on the ongoing project, an early 60-page version of [10] v1.
+He formally withdrew his signature on December 19, 2022 with a letter to C.P.
+stating wave of intellectual property rights.
+The authors of the papers resulting from this project are the solely main contributors and responsible for the constructions and the entirety of the proofs presented.
+The methods in the project build upon the author’s previous joint work with S.
+Ciamprone (2016–2018) and M.
+V.
+Giannone (2020–2022).
+The author is grateful to Yi-Zhi Huang for several comments on the proof of modularity for fusion categories of vertex operator algebras, and the history of Finkelberg-Kazhdan-Lusztig theorem.
 She is also grateful to Marco V.
-Giannone for pointing out to her incompleteness of the proof of Lemma 7.10 in version 7 of [13] and to Hans Wenzl for kind email correspondence on the centraliser algebras for the spinor representations, and for references on duality for tensor powers of the fundamental representation of Drinfeld-Jimbo quantum groups for the Lie types $B$, $D$, $G_{2}$, leading to the present paper, and to Makoto Yamashita for interesting discussions on [42] and [100] during his visit at the math department of Rome, Sapienza, and interesting comments on our work.
+Giannone for pointing out to her incompleteness of the proof of Lemma 7.10 in version 7 of [10] and to Hans Wenzl for kind email correspondence on the centraliser algebras for the spinor representations, and for references on duality for tensor powers of the fundamental representation of Drinfeld-Jimbo quantum groups for the Lie types $B$, $D$, $G_{2}$, leading to the present paper.
+She also thanks Makoto Yamashita for interesting discussions on [39] and [100] during his visit at the math department of Rome, Sapienza, and interesting comments on our work.
 Finally she thanks Pavel Etingof for informing her of [28].
+The author thanks Sapienza University of Rome for financial support Finanziamento di Ateneo per la Ricerca Scientifica, years 2018; 2019; 2020; 2021, 2023.
+Author’s Note.
+Our work was the subject of a severe cyber-attack in 2021, recognized by the judicial authority in Italy.
+The legal actions currently undergone at the Italian Procura della Repubblica (2021–2023) concern the unauthorized use of materials from that project in other publications.
+The papers Carpi et al, Commun.
+Math.
+Phys., 390 (2022), 169; Carpi et al, Commun.
+Math.
+Phys.
+402 (2023), 169; Adamo et al: Commun.
+Math.
+Phys.
+404, (2023), 1231, are legally involved in the related penal procedure.